### v2

# Codes Correcting Two Bursts of Exactly $b$ Deletions

Abstract
In this paper, we explore constructions for codes that correct two bursts of deletions, with each burst having length exactly $b$. Previously, the best known construction, derived using the syndrome compression technique, achieved a redundancy of at most $7\log n+O\left(\log n/\log\log n\right)$ bits. In this work, we present new constructions for all $q\geq 2$ that achieve redundancy at most $7\log n+O(\log\log n)$ when $b>1$. Additionally, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
Index Terms:
deletion, burst-deletion, error-correcting codes, DNA-based storage
I Introduction
A subset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], where by proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lowere bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that codes in [3, 4] have redundancy optimal up to a constant.
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log(n)+O(\log\log(n))$ [21]. And a following work [22] gave a $q$-ary $2$-deletion correcting code with redundancy $5\log(n)+O(\log\log(n))$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 24], wireless sensor networks and satellite communication devices [25]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein gave a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [26]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [27], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [28]. Schoeny’s result was generalized to non-binary alphabets in [29, 30]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[31]. It was proved in [28, 30] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [26] and [31, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [26, 28, 32, 33, 21, 34, 35, 31].
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. To the best of our knowledge, there is no explicit results about such codes. A related result can be found in a work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. In this paper, we construct codes which outperform the one derived by the syndrome compression technique. Our contributions include:
  • A binary code of length $n$ with redundancy at most $7\log n+(15b-5)\log\log n+O(1)$;
  • A $q$-ary code of length $n$ with redundancy at most $7\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $7\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
Here, it is assumed that $q$ and $b$ are constants with respect to $n$.
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV and Section V deal with binary and non-binary codes, respectively. In Section VI, we show that our idea in Section V can result in a new construction of non-binary two-deletion correcting code. Finally Section VII concludes this paper.
II Preliminaries
In this section, we introduce some necessary terminologies and notations.
For an integer $q\geq 2$ and a positive integer $n$, denote $\Sigma_{q}=\left\{0,1,\ldots,q-1\right\}$ and $\Sigma_{q}^{n}$ the set of all $q$-ary sequences with $n$ symbols. Let $\mathbf{x}\in\Sigma_{q}^{n}$ be a sequence. Unless otherwise stated, the $i$th coordinate of $\mathbf{x}$ is denoted by $x_{i}$, i.e., $\mathbf{x}=x_{1}\cdots x_{n}$. We call $n$ the length of $\mathbf{x}$ and denote $\left|\mathbf{x}\right|=n$. For a finite set $A$, we denote by $\left|A\right|$ the cardinality of $A$.
For two integers $m$ and $n$ such that $m\leq n$, let $[m,n]$ denote the set $\left\{m,m+1,\ldots,n\right\}$. If $m=1$, denote $[n]=[1,n]$. For a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ and a subset $I=\left\{i_{1},i_{2},\ldots,i_{t}\right\}\subseteq[n]$ where $i_{1}<i_{2}<\cdots<i_{t}$, we define $\mathbf{x}_{I}\triangleq x_{i_{1}}x_{i_{2}}\cdots x_{i_{t}}$. For each $I\subseteq[n]$, we say $\mathbf{x}_{I}$ is a subsequence of $\mathbf{x}$. In particular, if $I$ is an interval of $[n]$ (i.e., $I=[i,j]$ for some $1\leq i\leq j\leq n$), we say $\mathbf{x}_{I}$ is a substring of $\mathbf{x}$. A run in $\mathbf{x}$ is a maximal substring consisting of the same symbols. The number of runs of $\mathbf{x}$ is denoted by $r(\mathbf{x})$. For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, $1$, 00 and $1$. So $r\left(\mathbf{x}\right)=5$.
The concatenation of two sequences $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\mathbf{y}$. For example, let $\mathbf{x}=102$ and $\mathbf{y}=121$ be two sequences in $\Sigma_{3}^{3}$, then $\mathbf{x}\mathbf{y}=102121\in\Sigma_{3}^{6}$. Let $b$ and $n$ be two positive integers satisfying $b<n$. When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $11\leq i\leq n-b+$ 1.
In this paper, we focus on codes correcting two $b$-burst-deletions. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$, where $n>2b$. There are two ways to define two $b$-burst-deletions:
  1. (D1)
    the two bursts happen sequentially: from $\mathbf{x}$, we get $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then we get $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$ from $\mathbf{z}$;
  2. (D2)
    the two bursts happen simultaneously: from $\mathbf{x}$, we get $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n]. This definition is motivated by the nanopore sequencing technology, in which a DNA strand pass a nanoscale hole nucleotide by nucleotide [36, 37]. So we can assume that the two bursts do not overlap.
In fact, (D1) and (D2) are equivalent. Firstly, it is clear that see that (D1) covers (D2). Next, we show that (D2) also covers (D1).
Observation II.1
Let $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}$ is obtained from $\mathbf{x}$ by process (D1). Then there exist two intervals $I_{1}=[j_{1},j_{1}+b-1]$, $I_{2}=[j_{2},j_{2}+b-1]\subseteq[n]$, where $j_{2}-j_{1}\geq b$, such that $\mathbf{y}=\mathbf{x}_{[n]\setminus(I_{1}\cup I_{2})}$. In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlap substrings of length $b$ from $\mathbf{x}$.
Proof:
By assumption, there is $11\leq i_{1}\leq n-b+$ 1 and $11\leq i_{2}\leq n-2b+$ 1 such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$. If $i_{2}\geq i_{1}$, let $j_{1}=i_{1}$ and $j_{2}=i_{2}+b$. Then the conclusion follows.
Now suppose $11\leq i_{2}\leq i_{1}-$ 1. If $i_{2}\leq i_{1}-b$, let $j_{1}=i_{2}$ and $j_{2}=i_{1}$. Then the conclusion follows. If $i_{1}-b<i_{2}<i_{1}$, it is clear that $\mathbf{y}=\mathbf{x}_{[n]\setminus[i_{2},i_{2}+2b-1]}$. Let $j_{1}=i_{2}$ and $j_{2}=i_{2}+b$. Then the conclusion follows. ∎
Remark II.1
There is another possibility: the two bursts might overlap and result in a single burst that is shorter than $2b$. We do not take this situation into account, since it is covered by a more comprehensive problem: correcting two bursts of deletions, where each burst has length at most $b$. Our idea in this paper fails in this situation. We left this problem for future research.
For $t\in\{1,2\}$ and $n>tb$, define
[…]
When $b=1$, we use notation $\mathcal{B}_{t}(\mathbf{x})$ instead of $\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)$.
Definition II.1
Let $\mathcal{C}$ be a subset of $\Sigma_{q}^{n}$ with $\left|\mathcal{C}\right|\geq 2$. Suppose $t\in\{1,2\}$. We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$. In particular, if $b=1$, we call $\mathcal{C}$ a $t$-deletion correcting code.
Clearly, if any $\mathbf{x}\in\mathcal{C}$ can be uniquely and efficiently recovered from any given $\mathbf{x}^{\prime}\in\mathcal{B}_{t}^{b}(\mathbf{x})$, then $\mathcal{C}$ is a $t$-$b$-burst-deletion correcting code. Here, “efficiently” means that the time complexity of decoding $\mathbf{x}$ from $\mathbf{y}$ is polynomial in $n$. In this paper, we construct $t$-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
The redundancy of a code $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is defined to be $\rho(\mathcal{C})=\log\left(q^{n}/\left|\mathcal{C}\right|\right)$. All logarithms in this paper are to the base $2$. In addition, we always assume that $q$ and $b$ are fixed with respect to the code-length $n$.
Let $n^{\prime}$ and $n$ be two positive integers satisfying $n^{\prime}<n$. For each sequence $\mathbf{x}\in\Sigma_{q}^{n}$, let $\widetilde{\mathbf{x}}$ be the zero padding of $\mathbf{x}$ to the shortest length that is greater than $n$ and is divisible by $n^{\prime}$, that is, $\widetilde{\mathbf{x}}=\mathbf{x}0^{\left\lceil n/n^{\prime}\right\rceil n^{\prime}-n}$, and then $n^{\prime}\mid\left|\widetilde{\mathbf{x}}\right|$. We can represent $\mathbf{x}$ as an $n^{\prime}\times\left\lceil n/n^{\prime}\right\rceil$ array $A\left(\mathbf{x},n,n^{\prime}\right)=\left[a_{i,j}\right]$, where $a_{i,j}=\tilde{x}_{i+n^{\prime}j}$ for all $1\leq i\leq n^{\prime}$ and $10\leq j\leq\left\lceil n/n^{\prime}\right\rceil-$ 1. In other words, the $i$-th row of $A\left(\mathbf{x},n,n^{\prime}\right)$ is
$A\left(\mathbf{x},n,n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix representation of $\mathbf{x}$. If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
For example, let $n=7$, $\mathbf{x}=1011010\in\Sigma_{2}^{7}$ and $n^{\prime}=2$. Then $\widetilde{\mathbf{x}}=10110100\in\Sigma_{2}^{8}$ and
$A\left(\mathbf{x},7,2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}.$
If $n^{\prime}=3$, then $\widetilde{\mathbf{x}}=\mathbf{x}00$ and so
$A\left(\mathbf{x},8,3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}.$
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as an array $A(\mathbf{x},n,b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$.
The next observation is a straightforward corollary of II.1.
Observation II.2
For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{2}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$. Any two $b$-burst-deletions in $\mathbf{x}$ will induce two deletions in each row of $A(\mathbf{x})$. Furthermore, if the positions of the two deletions in $A(\mathbf{x})_{1}$ are $j_{1}$ and $j_{2}$, then for each $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred at coordinate $j_{1}-1$ or $j_{1}$, and the other deletion occurred at coordinate $j_{2}-1$ or $j_{2}$.
II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for all even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$. In Section VI, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
For a subset $\mathcal{E}\subseteq\Sigma_{q}^{n}$ and a sequence $\mathbf{x}\in\mathcal{E}$, define
$\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\mathcal{E}~{}:~{}\mathbf{y}\neq\mathbf{x}\text{ and }\mathcal{B}_{2}^{b}\left(\mathbf{y}\right)\cap\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)\neq\emptyset\right\}.$
Lemma II.1
[13, 20, 38] Let $\mathcal{E}\subseteq\Sigma_{q}^{n}$ be a code and $N>\max\left\{\left|\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)\right|~{}:~{}\mathbf{x}\in\mathcal{E}\right\}$. Suppose that the function $f:\Sigma_{q}^{n}\rightarrow\{0,1\}^{R(n)}$ (where $R(n)$ is a function of $n$ and $R(n)\geq 2$) satisfies the following property:
  1. $(\textup{{P}}1)$
    if $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}\in\mathcal{N}_{\Sigma_{q}^{n}}\left(\mathbf{x}\right)$, then $f\left(\mathbf{x}\right)\neq f\left(\mathbf{y}\right)$.
Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time, such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
Let $\mathcal{E}$ be a $1$-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1. Then Lemma II.1 asserts that if $\mathbf{x},\mathbf{y}\in\mathcal{E}$ are distinct codewords and $\bar{f}(\mathbf{x})=\bar{f}(\mathbf{y})$, we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$. Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code =$\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$ [^2] We select $\mathcal{E}$ to be a $1$-$b$-burst correcting code to obtain a better redundancy. If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$..
For the choice of $\mathcal{E}$, we have the following result.
Lemma II.2
[31, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a $1$-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$. So we let $N=q^{b}n^{3}$. Now the redundancy of $\mathcal{E}^{\prime}$ is at most $7\log n+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. To conclude our discussion, it remains to find an $f$ satisfying (P1) in Lemma II.1 such that $R(n)=O(\log n)$, which is given in Lemma II.5. The proof of Lemma II.5 is based on Lemmas II.3 and II.4.
The following lemma is a corollary of [17, Theorem 2].
Lemma II.3
[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}^{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
Lemma II.4
Suppose $q\geq 2$. There is a function $\xi_{1}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7\left\lceil\log q\right\rceil\log n+O_{q}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q}(1)$ is a constant dependent only on $q$.
Proof:
Any $\mathbf{x}\in\Sigma_{q}^{n}$ can be uniquely represented as an array
$M(\mathbf{x})\triangleq\begin{pmatrix}x_{1,1}&\cdots&x_{1,n}\\ \vdots&\cdots&\vdots\\ x_{\left\lceil\log q\right\rceil,1}&\cdots&x_{\left\lceil\log q\right\rceil,n}\end{pmatrix},$
where $x_{k,i}\in\{0,1\}$ be such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
Denote the $k$-th row of $M(\mathbf{x})$ by $M(\mathbf{x})_{k}$. Suppose $\mathbf{y}\in\mathcal{B}_{2}(\mathbf{x})$. It is clear that $M(\mathbf{y})_{k}\in\mathcal{B}_{2}\left(M(\mathbf{x})_{k}\right)$ for all $1\leq k\leq\left\lceil\log q\right\rceil$. Let $\xi(\cdot)$ be the function defined in Lemma II.3. For $\mathbf{x}\in\Sigma_{q}^{n}$, define $\xi_{1}(\mathbf{x})\triangleq\left(\xi\left(M(\mathbf{x})_{1}\right),\ldots,\xi\left(M(\mathbf{x})_{\left\lceil\log q\right\rceil}\right)\right)$. Then by Lemma II.3, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi\left(M(\mathbf{x})_{k}\right)$ is a binary vector of length $7\log n+O(1)$, we can see that $\xi_{1}(\mathbf{x})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log n+O_{q}(1)$. ∎
Lemma II.5
Suppose $b>1$, $n>2b$ and $q\geq 2$. There is a function $\psi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Proof:
Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{y})=A(\mathbf{y},n-b,b)$. Since $\mathbf{y}\in\mathcal{B}_{2}^{b}(\mathbf{x})$, we have $A(\mathbf{y})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. Define $\psi(\mathbf{x})\triangleq\left(\xi_{1}\left(A(\mathbf{x})_{1}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b}\right)\right)$. Then by Lemma II.4, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi_{1}(A(\mathbf{x})_{i})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log(n/b)+O_{q}(1)$, we can see that $\psi(\mathbf{x})$ is a binary vector of length $7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)$. ∎
Taking $f=\psi$ gives a function satisfying (P1) as needed.
Before proceeding to subsequent sections, we introduce a well-known result from number theory, which will be frequently used.
Lemma II.6 (Bertrand’s Postulate/Bertrand–Chebyshev theorem)
For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
III Bounds
We could not find any existing upper or lower bounds on the maximum size of a $2$-$b$-burst-deletion code. In this section, we will derive these bounds.
Let $M_{q,n,b}$ be the maximum size of a $2$-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
Theorem III.1
$M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
Proof:
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. It is clear that a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [39, page 100, Theorem 1] that
$M_{q,n,b}\geq\sum_{\mathbf{x}\in\Sigma_{q}^{n}}\frac{1}{d(\mathbf{x})+1}.$ (1)
For $\mathbf{x}\in\Sigma_{q}^{n}$, by II.1, we conclude that each $\mathbf{y}$ (including $\mathbf{x}$ itself) with $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$ can be obtained as follows: 1) deleting two non-overlap substrings of length $b$ from $\mathbf{x}$, resulting in a sequence $\mathbf{z}\in\Sigma_{q}^{n-2b}$; 2) inserting two sequences of length $b$ into $\mathbf{z}$. Therefore, we have $d(\mathbf{x})+1\leq\binom{n}{2}^{2}q^{2b}$. ∎
In [40, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
Theorem III.2
For $q\geq 2$, let $f(q)=\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. When $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
$M_{q,n,b}\leq\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}.$
Proof:
Define $m=n/b-1$. Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a $2$-$b$-busrt-deletion correcting code. Set $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and $t=\left(1-\frac{1}{q}-\epsilon\right)m$. We partition $\mathcal{C}$ into two disjoint subsets: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}_{1}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\geq t+2\text{ for some }1\leq i\leq b\right\}$ and $\mathcal{C}_{2}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\leq t+1\text{ for all }1\leq i\leq b\right\}$ (recall that $r(\cdot)$ denotes the number of runs). To derive an upper bound of $\left|\mathcal{C}\right|$, it is sufficient to upper bound $\left|\mathcal{C}_{1}\right|$ and $\left|\mathcal{C}_{2}\right|$.
For any $\mathbf{x}\in\mathcal{C}$, define $\mathcal{A}_{2}^{b}(\mathbf{x})=\left\{A(\mathbf{x}^{\prime}):\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})\right\}$. Let $A(\ell,k)$ be the array obtained by deleting the $\ell$-th and the $k$-th columns of $A(\mathbf{x})$ for $1\leq\ell\neq k\leq n/b$. It is clear that $A(\ell,k)\in\mathcal{A}_{2}^{b}(\mathbf{x})$ and
$\cup_{1\leq\ell\neq k\leq\frac{n}{b}}\left\{A(\ell,k)_{i}\right\}=\mathcal{B}_{2}(A(\mathbf{x})_{i}),$
for all $1\leq i\leq b$. By [41, eq. (11)], we know that
$\binom{r-1}{2}\leq\left|\mathcal{B}_{2}(\mathbf{v})\right|\leq\binom{r+1}{2}$
for any sequence $\mathbf{v}\in\Sigma_{q}^{n}$ with exactly $r$ runs. Since $\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ and $\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|=\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|$ for all $\mathbf{x}$, it follows that
$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2},$ (2)
for all $\mathbf{x}\in\mathcal{C}_{1}$. Since $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a $2$-$b$-burst-deletion correcting code. So we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for all distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}_{1}$. Then it follows from Equation 2 that $\left|\mathcal{C}_{1}\right|\binom{t+1}{2}\leq\sum_{\mathbf{x}\in\mathcal{C}_{1}}\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\leq q^{n-2b}$ and hence
$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ $=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$ (3)
$=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2b^{2}}{(1-1/q-\epsilon)^{2}(1-b/n)^{2}}$
$=\frac{2b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}\cdot\frac{1}{\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}}$
$\leq\frac{3b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}$
as long as $\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, which is possible when $n$ is sufficiently large.
Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define a set
$\mathcal{C}^{\prime}\triangleq\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$
Since $r\left(A(\mathbf{x})_{i}\right)\leq t+1$ for any $\mathbf{x}\in\mathcal{C}_{2}$ and $1\leq i\leq b$, we conclude that $\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|$. So it suffices to estimate an upper bound of $\left|\mathcal{C}^{\prime}\right|$.
Since $m=n/b-1$, by definition of $\mathcal{C}^{\prime}$, we have
$\left|\mathcal{C}^{\prime}\right|$ =|$=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (4)
$=\left|\cup_{j=1}^{t+1}\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|^{b}$
$=\left(\sum_{j=1}^{m+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
=$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
where (a) follows from the well-known result (see the proof of [42, Theorem 3.1]):
|$\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|=\binom{m}{j-1}q(q-1)^{j-1}.$
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [43, Proposition 3.3.3] that
$\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\leq q^{mH_{q}(t/m)}=q^{mH_{q}(1-1/q-\epsilon)},$ (5)
where $H_{q}(x)\triangleq x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function. Here, $\log_{q}x=\frac{\log x}{\log q}$ for any positive real number $x$.
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. So by Lemma .1, we have.
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Now it follows from Equations 4 and 5 that
$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}\overset{(b)}{<}(1.121)^{3b}\frac{q^{n}}{n^{3}},$ (6)
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 3 and 6. ∎
Then next corollary is a direct consequence of Theorems III.1 and III.2.
Corollary III.1
Suppose that $q$ and $b$ are fixed. Let $f(q)$ be defined in Theorem III.2. When $n\geq\max\left\{30,\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}\right\}$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
$2\log n+(2b-2)\log q+2\log(q-1)-2\log b-2\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code with maximum size.
IV Binary $2$-$b$-Burst-Deletion Correcting Codes
In this section, we will construct a class of binary $2$-$b$-burst-deletion correcting codes with redundancy at most $7\log n+O(\log\log n)$. First, we introduce some $1$-deletion correcting codes and $2$-deletion correcting codes, which will be useful later.
Lemma IV.1
[21, Lemma 1] For any integer $n\geq 3$, there exists a function $\mathsf{VT}:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{\log(n+1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\mathsf{VT}\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
Definition IV.1 (regularity)
A sequence $\mathbf{x}\in\Sigma_{2}^{n}$ is said to be $d$-regular if each substring of $\mathbf{x}$ of length at least $d\log n$ contains both $00$ and $11$.
For the number of regular sequences, we have the following lemma.
Lemma IV.2
[19, Lemma 11] The number of $d$-regular sequences of length $n$ is at least $2^{n-1}$, as long as $d\geq 7$ and $n$ be such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
Lemma IV.3
[19, Theorem 7] Suppose $d\geq 7$. There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{d}(1)$ denotes a constant depending only on $d$.
If the first row $A(\mathbf{x})_{1}$ is regular, then by correcting the two deletions in $A(\mathbf{x})_{1}$, we can approximate the location of the two deletions, as ensured by the following lemma.
Lemma IV.4
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. Then we can
  1. $(1)$
    either find distinct runs $\mathbf{x}_{J_{1}}$ and $\mathbf{x}_{J_{2}}$ of $\mathbf{x}$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, such that $i_{1}\in J_{1}$ and $i_{2}\in J_{2}$.
  2. $(2)$
    or find an interval $J\subseteq[n]$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, of length at most $3d\log n$ such that $i_{1},i_{2}\in J$.
From now on, we assume that $A(\mathbf{x})_{1}$ is $7$-regular. Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$. Then by II.2, we have $A\left(\mathbf{x}^{\prime}\right)_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Lemma IV.3 ensures that when given $\eta\left(A(\mathbf{x})_{1}\right)$, one can uniquely recover $A(\mathbf{x})_{1}$ from $A\left(\mathbf{x}^{\prime}\right)_{1}$. Then by Lemma IV.4, we can know the approximate positions of deletions in $A(\mathbf{x})_{1}$. This further enables us to approximately determine the positions of deletions in $A(\mathbf{x})_{i}$ for all $2\leq i\leq b$. Based on the two cases in Lemma IV.4, we split our discussion into Section IV-A and Section IV-B. Our codes for correcting two $b$-burst-deletions is presented in Theorem IV.1, whose proof is divided in into Lemmas IV.5, IV.6 and IV.7.
IV-A When Case (1) in Lemma IV.4 happens
Let $A(\mathbf{x})_{1,I_{j}}$ be the $j$th run of $A(\mathbf{x})_{1}$, where $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$ (recall that $r\left(A(\mathbf{x})_{1}\right)$ is the number of runs in $A(\mathbf{x})_{1}$) and $I_{j}\subseteq[n/b]$ is the index set of $A(\mathbf{x})_{1,I_{j}}$. For simplicity, denote $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$ and $p_{r\left(A(\mathbf{x})_{1}\right)}=n/b$. Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $p_{j}-p_{j-1}=\left|I_{j}\right|\leq 7\log(n/b)$. Note that $I_{j}$ depends on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $j_{1}<j_{2}$. Then for all $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (if $j_{1}=1$, then this deletion occurred in the interval $[1,p_{1}]$) and the other occurred in the interval $\left[p_{j_{2}-1},p_{j_{2}}\right]$. So $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin\left[p_{j_{1}-1},p_{j_{1}}\right]\cup\left[p_{j_{2}-1},p_{j_{2}}\right]$. Specifically, we have
$A(\mathbf{x})_{i,k}=\begin{cases}A(\mathbf{x}^{\prime})_{i,k},\text{ if }k<p_{j_{1}-1},\\ A(\mathbf{x}^{\prime})_{i,k-1},\text{ if }p_{j_{1}}<k<p_{j_{2}-1},\\ A(\mathbf{x}^{\prime})_{i,k-2},\text{ if }k>p_{j_{2}}.\end{cases}$ (7)
for all $2\leq i\leq b$.
Therefore, we need to decode $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$, for all $2\leq i\leq b$. Notice that when $j\geq 2$, the interval $[p_{j-1},p_{j}]$ intersects with both $I_{j-1}$ and $I_{j}$. Our process of decoding $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$ is divided in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
For the first step, for all $1\leq j<r\left(A(\mathbf{x})_{1}\right)$, define
$f_{j}(\mathbf{x})=\left(A(\mathbf{x})_{2,p_{j}},\ldots,A(\mathbf{x})_{b,p_{j}}\right),$
which can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{b-1}$. For $s=0,1$, let
$K_{s}=\left\{2\leq j\leq r\left(A(\mathbf{x})_{1}\right):j\equiv s\pmod{2}\right\}.$
Let $Q_{1}\geq\max\{n/b,N_{1}\}$ be the smallest prime. For $s=0,1$, Define
$f^{(s,0)}(\mathbf{x})=\sum_{j\in K_{s}}f_{j-1}(\mathbf{x})\pmod{2N_{1}},$ (8)
$f^{(s,1)}(\mathbf{x})=\sum_{j\in K_{s}}(j-1)f_{j-1}(\mathbf{x})\pmod{Q_{1}}.$
In the following, we may assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4.
Lemma IV.5
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. Then given $f^{(s,0)}(\mathbf{x})$ and $f^{(s,1)}(\mathbf{x})$ ($s=0,1$), we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$. By Equation 7, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
When $j_{1}=1$, we have $p_{j_{1}-1}=0$ and hence $A(\mathbf{x})_{p_{j_{1}-1}}$ is not defined. So we only need to find $A(\mathbf{x})_{p_{j_{2}-1}}$. Let $s\in\{0,1\}$ be such that $j_{2}\equiv s\pmod{2}$. If $s=1$, then $f_{j-1}(\mathbf{x})$ is known for all $j\in K_{s}\setminus\{j_{2}\}$. Let $\delta=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$. Then it follows from the first line of Equation 8 that
$f_{j_{2}-1}(\mathbf{x})\equiv\delta\pmod{2N_{1}}.$
Since $0\leq f_{j_{2}-1}(\mathbf{x})<N_{1}$ and $0\leq\delta<2N_{1}$, we conclude that $f_{j_{2}-1}(\mathbf{x})=\delta$.
If $s=0$, then $j_{1}+1\equiv s\pmod{2}$. Since $j_{2}-j_{1}\geq 2$, we have $j_{1}+1\neq j_{2}$. Then $f_{j-1}(\mathbf{x})$ is known for all $j\in K_{s}\setminus\{j_{1}+1,j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 8 that
$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$ (9)
$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (10)
Since $0\leq f_{j_{1}}(\mathbf{x}),f_{j_{2}-1}(\mathbf{x})<N_{1}$ and $0\leq\delta^{(0)}<2N_{1}$, we conclude from Equation 9 that $f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})=\delta^{(0)}$. Combining this with Equation 10, we get
$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (11)
$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Since $0\leq j_{1}<j_{2}-1<r\left(A(\mathbf{x})_{1}\right)\leq n/b\leq Q_{1}$, we have $j_{1}\not\equiv j_{2}-1\pmod{Q_{1}}$, which implies that system (11) can be uniquely solved and the unique solution is $\left(f_{j_{1}}(\mathbf{x})\pmod{Q_{1}},f_{j_{2}-1}(\mathbf{x})\pmod{Q_{1}}\right)$. Since $0\leq f_{j_{1}}(\mathbf{x}),f_{j_{2}-1}(\mathbf{x})<N_{1}\leq Q_{1}$, we have $f_{j_{1}}(\mathbf{x})=f_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $f_{j_{2}-1}(\mathbf{x})=f_{j_{2}-1}(\mathbf{x})\pmod{Q_{1}}$.
Now assume $j_{1}>1$. First, suppose $j_{1}\equiv j_{2}\pmod{2}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$, where $s\in\{0,1\}$ be such that $j_{1}\equiv j_{2}\equiv s\pmod{2}$. Then it follows from Equation 8 that
$f_{j_{1}-1}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
$(j_{1}-1)f_{j_{1}-1}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Similar to previous case, we can uniquely solve $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$ from this equation system.
Now suppose $j_{1}\equiv s_{1}\pmod{2}$ and $j_{2}\equiv s_{2}\pmod{2}$, where $s_{1},s_{2}\in\{0,1\}$ and $s_{1}\neq s_{2}$. We first show how to calculate $f_{j_{2}-1}(\mathbf{x})$. Since $j_{1}\not\equiv s_{2}\pmod{2}$, it holds that $j_{1}+1\equiv s_{2}\pmod{2}$. Note that $j_{1}+1\neq j_{2}$ since $j_{2}-j_{1}\geq 2$ by assumption. It is clear that $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s_{2},0)}(\mathbf{x})-\sum_{j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s_{2},1)}(\mathbf{x})-\sum_{j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 8 that
$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Similarly, we can obtain $f_{j_{2}-1}(\mathbf{x})$ by solving this equation system.
It remains to show how to calculate $f_{j_{1}-1}(\mathbf{x})$. If $j_{2}=r\left(A(\mathbf{x})_{1}\right)$, then $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{1}}\setminus\{j_{1}\}$. Therefore,
$f_{j_{1}-1}(\mathbf{x})=\left(f^{(s_{1},0)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}.$
If $j_{2}<r\left(A(\mathbf{x})_{1}\right)$, then $j_{2}+1\equiv j_{1}\pmod{2}$. It is clear that $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}$. Let $\delta^{(0)}=\left(f^{(s_{1},0)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s_{1},1)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 8 that
$f_{j_{1}-1}(\mathbf{x})+f_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
$(j_{1}-1)f_{j_{1}-1}(\mathbf{x})+j_{2}f_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
We can obtain $f_{j_{1}-1}(\mathbf{x})$ by solving this equation system. ∎
Upon knowing the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$, it remains to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$. Recall that $j_{2}-j_{1}\geq 2$. This implies $p_{j_{2}-1}>p_{j_{1}}$ and thus for all $2\leq i\leq b$, the deletion in $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ does not affect $A(\mathbf{x})_{i,I_{j_{1}}}$. Therefore, we can conclude that $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$. So we can apply Lemma IV.1 to decode $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}$, respectively.
For $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
$g_{j}(\mathbf{x})=\left(\mathsf{VT}\left(A(\mathbf{x})_{2,I_{j}}\right),\cdots,\mathsf{VT}\left(A(\mathbf{x})_{b,I_{j}}\right)\right).$
Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $\left|I_{j}\right|\leq 7\log(n/b)$. Then it follows from Lemma IV.1 that
$\mathsf{VT}\left(A(\mathbf{x})_{i,I_{j}}\right)\in\{0,1\}^{\log\left(7\log(n/b)+1\right)}.$
So $g_{j}(\mathbf{x})$ can be viewed as an integer in $[0,N_{2}-1]$, where $N_{2}=2^{(b-1)\log\log(n)+O_{b}(1)}$. Here, $O_{b}(1)$ is a constant dependent only on $b$.
Let $Q_{2}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
$g^{(0)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}g_{j}(\mathbf{x})\pmod{2N_{2}},$ (12)
$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{2}}$
Lemma IV.6
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. If the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are known, then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
As discussed above, we have $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$. So by Lemma IV.1, to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, it is sufficient to obtain the values of $g_{j_{1}}(\mathbf{x})$ and $g_{j_{2}}(\mathbf{x})$.
Since the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are given, the value of $g_{j}(\mathbf{x})$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{2}}$. Then it follows from Equation 12 that
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{2}},$ (13)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{2}}.$ (14)
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 13 that $g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})=\delta^{(0)}$. Combining this with Equation 14, we get
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{2}},$ (15)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{2}}.$
Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{2}$, system (15) can be uniquely solved and the unique solution is
$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{2}},g_{j_{2}}(\mathbf{x})\pmod{Q_{2}}\right).$
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{2}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{2}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{2}}$. ∎
IV-B When Case (2) in Lemma IV.4 happens
In this case, we can find an interval $J\subseteq[n/b]$ of length at most $21\log(n/b)$ such that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$. Suppose $J=[c,d]$. Then for all $2\leq i\leq b$, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c-1,d]$, which is of length at most $121\log n+$ 1.
Let $P=21\log\left(n/b\right)+1$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/bP\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/bP\right\rceil-1.\end{cases}$
Let $\xi(\cdot)$ be the function defined in Lemma II.3. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
$h_{j}\left(\mathbf{x}\right)=\left(\xi\left(A(\mathbf{x})_{2,J_{j}}\right),\xi\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi\left(A(\mathbf{x})_{b,J_{j}}\right)\right).$
Since $\left|J_{j}\right|\leq 2P=42\log\left(n/b\right)+2$ for all $j$, each $h_{j}(\mathbf{x})$ is a binary vector of length at most $7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)$ and so we can view $h_{j}(\mathbf{x})$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)}$. Let
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ even}}h_{j}(\mathbf{x})\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ odd}}h_{j}(\mathbf{x})\pmod{N_{3}}.\end{array}$ (16)
Lemma IV.7
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$, where $J\subseteq[n/b]$ is an interval of length $21\log(n/b)$ and is known to us. Given $h_{e}\left(\mathbf{x}\right)$ and $h_{o}\left(\mathbf{x}\right)$, one can uniquely recover $A(\mathbf{x})_{i}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$. Without loss of generality, we assume that $j_{0}$ is even. Then $A(\mathbf{x})_{i,J_{j}}$ and thus $h_{j}(\mathbf{x})$ is known to us for each even $j\neq j_{0}$. So it remains to recover $A(\mathbf{x})_{J_{j_{0}}}$. If $J_{j_{0}}=[c,d]$, it is clear that $A(\mathbf{x}^{\prime})_{i,[c,d-2]}=\mathcal{B}_{2}\left(A(\mathbf{x})_{i,[c,d]}\right)$, for all $2\leq i\leq b$. By Lemma II.3, it is suffices to obtain $h_{j_{0}}(\mathbf{x})$. By Equation 16, we have
$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, it follows that
$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
Now the proof is completed. ∎
Now we are ready to present our main result in this section.
Theorem IV.1
Suppose $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1},Q_{2}$, $f^{(s,\ell)}(\mathbf{x})$, $g^{(s)}(\mathbf{x})$ (where $s,\ell\in\{0,1\}$), $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{s,0}<2N_{1}$ ($s=0,1$), $0\leq b_{s,1}<Q_{1}$ ($s=0,1$), $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{2}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f^{(s,\ell)}(\mathbf{x})=b_{s,\ell}\text{ for }s,\ell\in\{0,1\},\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $7\log n+(15b-5)\log\log n+O_{b}(1)$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n,b)$. Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Since $A(\mathbf{x})_{1}$ is $7$-regular, by Lemma IV.3, we can recover $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ based on $\eta\left(A(\mathbf{x})_{1}\right)$. Now according to Lemmas IV.4, IV.5, IV.6 and IV.7, it is clear that $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code.
Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. By Lemma IV.2, there are at least $2^{n-1}$ $7$-regular sequences when $n\geq 9$. It then follows from the pigeonhole principle that there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot(2N_{1})^{2}\cdot(Q_{1})^{2}\cdot(2N_{2})\cdot Q_{2}\cdot(N_{3})^{2}},$
which implies by definition that
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+2\log Q_{1}+\log Q_{2}+2\log(2N_{1})+\log(2N_{2})+2\log N_{3}+1$
$=4\log n+2\log Q_{1}+\log Q_{2}+(15b-5)\log\log n+O_{b}(1)$
$\leq 7\log n+(15b-5)\log\log n+O_{b}(1).$
The last inequality follows from the fact $n/b\leq Q_{1},Q_{2}<2n/b$, which is guaranteed by Lemma II.6. ∎
V Non-binary $2$-$b$-Burst-Deletion Correcting Codes
Let $q\geq 2$. In this section, we will extend our idea in Section IV to construct $q$-ary codes capable of correcting two $b$-burst-deletions. Surprisingly, our construction in this section gives rise to a binary code which is better than that in Theorem IV.1 when $b\geq 17$.
The first step is to establish a result similar to Lemma IV.4 to approximately determine the error positions. To that end, we associate with any $q$-ary sequence a binary sequence.
For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
Remark V.1
Although any $x\in\Sigma_{q}$ can be decomposed as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, it might exist some $u\in\{0,1\}$ and $0\leq v<\left\lceil q/2\right\rceil$ such that $\left\lceil q/2\right\rceil u+v\notin\Sigma_{q}$. In fact, if $q$ is even, then for any $u\in\{0,1\}$ and any $0\leq v<\left\lceil q/2\right\rceil$, we have $\left\lceil q/2\right\rceil u+v\in\Sigma_{q}$. But when $q$ is odd, we have $\left\lceil q/2\right\rceil+\left\lceil q/2\right\rceil-1=q\notin\Sigma_{q}$.
We first generalize the concept of $d$-regular sequence to sequences over general alphabets.
Definition V.1
A sequence $\mathbf{x}\in\Sigma_{q}^{n}$ is said to be $d$-regular if $u(\mathbf{x})$ is $d$-regular. In other words, $\mathbf{x}$ is $d$-regular if and only if any substring of $\mathbf{x}$ of length at least $d\log n$ contains two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$.
[…]
Lemma V.1
Let $q\geq 2$.
  • If $q$ is even, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n}/2\geq q^{n-1}$ for all $d\geq 7$ and $n$ such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
  • If $q$ is odd, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n-1}$ for all $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. Note that when $d\geq 10$, we have $-1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large. In particular, when $d=10$, it suffices to require $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
Proof:
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. Then it follows from Lemma IV.2 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
Now suppose that $q\geq 3$ is odd. In this case, previous argument does not hold. This is because when $q$ is odd, $u_{x}=1$ and $v_{x}=(q-1)/2$, we have $\left\lceil q/2\right\rceil u_{x}+(q-1)/2=q\notin\Sigma_{q}$. Fortunately, we can apply similar ideas in proofs of [19, Lemma 11] and [22, Lemma 5] to derive our lower bound on $\left|\mathcal{R}_{q,n,d}\right|$. For $m\geq 2$, define
$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }0\leq v_{i},v_{i+1}\leq(q-1)/2\right\},$
$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }(q+1)/2\leq v_{i},v_{i+1}<q\right\}.$
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}^{3}\right)=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
For $\mathbf{v}\in\Sigma_{q}^{m}$, let $\overline{\mathbf{v}}\triangleq(q-v_{1})\cdots(q-v_{m})$. It is easy to verify that if $\mathbf{v}\in S_{m}^{U}$, then $\overline{\mathbf{v}}\in S_{m}^{L}$. Then it follows that $\left|S_{m}^{U}\right|\leq\left|S_{m}^{L}\right|<(0.87q)^{m}$.
Let $m=\left\lfloor\frac{d}{2}\log n\right\rfloor$ and
$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\text{ and }v_{j}.v_{j+1}\geq(q+1)/2\right\}.$
Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$. Now let $k=\left\lfloor n/m\right\rfloor$ and define
=$\mathcal{R}^{\prime}=\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-k\left\lfloor n/m\right\rfloor}.\end{array}\right\}$
It is easy to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus
$\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}^{\prime}\right|$ $=\left|Q\right|^{k}q^{n-km}$
$>\left(q^{m}-2(0.87q)^{m}\right)^{k}q^{n-km}$
$=q^{n}\left(1-2(0.87)^{m}\right)^{k}$
$\overset{(a)}{\geq}q^{n}\left(1-2k(0.87)^{m}\right)$
$=q^{n}\left(1-2\left\lfloor\frac{n}{m}\right\rfloor 2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\geq q^{n}\left(1-2\frac{n}{m}2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\overset{(b)}{\geq}q^{n}\left(1-2\frac{n}{m}2^{\frac{d}{2}\log n\cdot\log(0.87)-\log(0.87)}\right)$
$=q^{n}\left(1-\frac{2}{0.87\left\lfloor\frac{d}{2}\log n\right\rfloor}n^{1+\frac{d}{2}\log(0.87)}\right)$
$\overset{(c)}{\geq}q^{n-1}.$
Here, (a) follows from the fact that $(1+x)^{r}\geq 1+rx$ for any integer $r\geq 1$ and any real number $x\geq-1$; (b) follows from the fact that $\log(0.87)<0$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor\geq\frac{d}{2}\log n-1$; (c) follow from the fact that $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. ∎
In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$. It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\}.$
Then $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. Like in Section IV-A, let $U(\mathbf{x})_{1,I_{j}}$ ($1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$) be all runs in $U(\mathbf{x})_{1}$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$.
Recall that $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$ We associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most $b-1$ in $[n]$ to the left of $I_{j}^{\prime}$. See Figure 1 for an illustration. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
[…]
(a) $U(\mathbf{x})$: $I_{j}$ in $\left[n/b\right]$
[…]
(b) $\mathbf{x}$: $I_{j}^{\prime}$ and $I_{j}^{L}$ in $\left[n\right]$
Figure 1: Illustration of relationship between $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$.
Lemma V.2
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.3. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can
  1. $(1)$
    either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$;
  2. $(2)$
    or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
Proof:
Since $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, we have $u(\mathbf{x}^{\prime})=u(\mathbf{x})_{[n]\setminus(D_{1}\cup D_{2})}$. It follows that $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Suppose that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $U(\mathbf{x})_{1}$ by deleting $U(\mathbf{x})_{1,k_{1}}$ and $U(\mathbf{x})_{1,k_{2}}$. Recall that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. So $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.3, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.4, we can
  • either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $k_{1}\in I_{j_{1}}$ and $k_{2}\in I_{j_{2}}$;
  • or find an interval $[c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Note that for the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
In the following, we always assume $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Furthermore, suppose that $A(\mathbf{x})_{1}$ is $d$-regular. We need to find conditions on $\mathbf{x}$ which enable us to recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$. As in Section IV, we split our discussion into Sections V-A and V-B. Our main result in this section is presented in Theorem V.1.
V-A When Case (1) in Lemma V.2 happens
In this case, we already found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Note that $I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]$ and $I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}=\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]$. Then it follows that
$x_{k}=\begin{cases}x^{\prime}_{k},\text{ if }k<p_{j_{1}-1}b-b+2,\\ x^{\prime}_{k-b},\text{ if }p_{j_{1}}b<k<p_{j_{2}-1}b-b+2,\\ x^{\prime}_{k-2b},\text{ if }p_{j_{2}}b<k\leq n.\end{cases}$ (17)
We can assume $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma V.2.
Our first goal is to recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$. Note that if $j=1$, then $p_{j-1}=0$. So we only need to focus on $\mathbf{x}_{I_{j}^{L}}$ for $j\geq 2$. For $s=0,1$, let
$K_{s}=\left\{2\leq j\leq r\left(U(\mathbf{x})_{1}\right):j\equiv s\pmod{2}\right\}.$
Recall that $\left|I_{j}^{L}\right|\leq b-1$. So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N_{1}-1]$, where $N_{1}=q^{b-1}$. Let $Q_{1}\geq\max\{n/b,N_{1}\}$ be the smallest prime. For $s=0,1$, Define
$f^{(s,0)}(\mathbf{x})=\sum_{j\in K_{s}}\mathbf{x}_{I_{j}^{L}}\pmod{2N_{1}},$ (18)
$f^{(s,1)}(\mathbf{x})=\sum_{j\in K_{s}}(j-1)\mathbf{x}_{I_{j}^{L}}\pmod{Q_{1}}.$
Lemma V.3
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f^{(0,0)}(\mathbf{x})$, $f^{(0,1)}(\mathbf{x})$, $f^{(1,0)}(\mathbf{x})$ and $f^{(1,1)}(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
Proof:
We first consider the case $j_{1}=1$, then $I_{j_{1}}^{L}=\emptyset$ and so we only need to determine $\mathbf{x}_{I_{j_{2}}^{L}}$. Let $s\in\{0,1\}$ be such that $j_{2}\equiv s\pmod{2}$. If $j_{1}\equiv s\pmod{2}$, then by Equation 17, $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{2}\}$. By the first line of Equation 18, we have $\mathbf{x}_{I_{j_{2}}^{L}}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ since $0\leq\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}$. If $j_{1}\not\equiv s\pmod{2}$, then $j_{1}+1\equiv s\pmod{2}$. Since $j_{2}-j_{1}\geq 2$, it holds that $j_{1}+1\neq j_{2}$. By Equation 17, $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{1}+1,j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}(j-1)\mathbf{x}_{I_{j}^{L}}\right)\pmod{Q_{1}}$. Then it follows from Equation 18 that
$\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{2N_{1}},$ (19)
$j_{1}\mathbf{x}_{I_{j_{1}+1}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$ (20)
Since $0\leq\mathbf{x}_{I_{j_{1}+1}^{L}},\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}$ and $0\leq\delta^{(0)}<2N_{1}$, we conclude from Equation 19 that $\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}=\delta^{(0)}$. Combining this with Equation 20, we get
$\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{Q_{1}},$ (21)
$j_{1}\mathbf{x}_{I_{j_{1}+1}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$
Since $0\leq j_{1}<j_{2}-1<Q_{1}$, we can get the unique solution of (21) in the field $\mathbb{F}_{Q_{1}}$: $\left(\mathbf{x}_{I_{j_{1}+1}^{L}}\pmod{Q_{1}},\mathbf{x}_{I_{j_{2}}^{L}}\pmod{Q_{1}}\right)$. Since $0\leq\mathbf{x}_{I_{j_{1}+1}^{L}},\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}\leq Q_{1}$, it holds that $\mathbf{x}_{I_{j_{1}+1}^{L}}=\mathbf{x}_{I_{j_{1}+1}^{L}}\pmod{Q_{1}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}=\mathbf{x}_{I_{j_{2}}^{L}}\pmod{Q_{1}}$. In particular, we obtained $\mathbf{x}_{I_{j_{2}}^{L}}$.
Next, we consider the case $j_{1}>1$. Let $s_{1},s_{2}\in\{0,1\}$ be such that $j_{1}\equiv s_{1}\pmod{2}$ and $j_{2}\equiv s_{2}\pmod{2}$. If $s_{1}=s_{2}=s$, then $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{1},j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}(j-1)\mathbf{x}_{I_{j}^{L}}\right)\pmod{Q_{1}}$. Then it follows from Equation 18 that
$\mathbf{x}_{I_{j_{1}}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{2N_{1}},$
$(j_{1}-1)\mathbf{x}_{I_{j_{1}}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$
Similar to the case $j_{1}=1$, we can uniquely solve $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$.
Now suppose $s_{1}\neq s_{2}$. In this case, we have $j_{1}+1\equiv s_{2}\pmod{2}$. Similar to the case when $j_{1}=1$ and $j_{1}\not\equiv j_{2}\pmod{2}$, we can obtain $\mathbf{x}_{I_{j_{2}}^{L}}$. It remains to show how to obtain $\mathbf{x}_{I_{j_{1}}^{L}}$. It is clear that $j_{2}+1\equiv s_{1}\pmod{2}$. Now similar to the case when $j_{1}=1$ and $j_{1}\not\equiv j_{2}\pmod{2}$, we can obtain $\mathbf{x}_{I_{j_{1}}^{L}}$. ∎
Once we know $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, it remains to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, where $I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b+1,j_{1}b\right]$ and $I_{j_{1}}^{\prime}=\left[p_{j_{2}-1}b+1,j_{2}b\right]$. Since $j_{2}-j_{1}\geq 2$, we have $(p_{j_{2}-1}-1)b+2>p_{j_{1}}b$. This implies that the $b$-burst-deletion in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ does not affect $\mathbf{x}_{I_{j_{1}}^{\prime}}$. Therefore, we conclude that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So to decode $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}$, we need the function $\phi(\cdot)$ defined in Lemma II.2.
Next, we will apply $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ to decode $\mathbf{x}_{I_{j}^{\prime}}$. Recall that $U(\mathbf{x})_{1}$ is $d$-regular. So by definition, we have $\left|I_{j}^{\prime}\right|=(p_{j}-p_{j-1})b\leq db\log n$. Then by Lemma II.2, $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{\log(db\log n)+O_{q,b}(1)}$. Let $Q_{2}\geq\max\{n/b,N_{2}\}$ be the smallest prime. Define
$g^{(0)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{2N_{2}},$ (22)
$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{2}}.$
Lemma V.4
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Furthermore, suppose that $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known. Then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}$.
Proof:
Since $j_{2}-j_{1}\geq 2$, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So by Lemma II.2, to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, it it sufficient to obtain $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known, it is easy to see that $\mathbf{x}_{I_{j}^{\prime}}$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{2}}$. Then it follows from Equation 22 that
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{2N_{2}},$ (23)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{2}}.$ (24)
Since $0\leq\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right),\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 23 that $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)=\delta^{(0)}$. Combining this with Equation 24, we get
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{2}},$ (25)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{2}}.$
Now similar to Lemma IV.6, we can uniquely solve $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$ from (25). ∎
V-B When Case (2) in Lemma V.2 happens
In this case, we can find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$. This implies that $\mathbf{x}_{k}$ is known to us for all $k\notin J$. Suppose $J=[c_{1},c_{2}]$. Then $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}\in\mathcal{B}_{b}^{2}\left(\mathbf{x}_{J}\right)$. To recover $\mathbf{x}_{J}$, we apply a trivial $2$-$b$-burst-deletion correcting code of length $2\left|J\right|$, which is given in Lemma II.5.
Let $P=3bd\log(n/b)+b-1$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$. Let the function $\psi(\cdot)$ be defined in Lemma II.5. Since $\left|J_{j}\right|\leq 2P$, we can view each $\psi\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7b\left\lceil\log q\right\rceil\log(2P/b)+O_{q,b}(1)}$ and define
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{0}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{1}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}}.\end{array}$ (26)
Lemma V.5
Suppose that we found an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$. Then given $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$, we can uniquely recover $\mathbf{x}_{J}$ (and thus $\mathbf{x}$) from $\mathbf{x}^{\prime}$.
Proof:
We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$. Without loss of generality, we assume $j_{0}\in L_{0}$. Then $\psi\left(\mathbf{x}_{J_{j}}\right)$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$. In addition, if $J_{j_{0}}=[c_{1},c_{2}]$, it is clear that $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}=\mathcal{B}_{2}^{b}\left(\mathbf{x}_{J_{0}}\right)$. It follows from Equation 26, that
$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, we have
$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
Now by Lemma II.5, one can uniquely recover $\mathbf{x}_{J_{0}}$ and thus $\mathbf{x}_{J}$. ∎
Theorem V.1
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1},Q_{2}$, $f^{(s,\ell)}(\mathbf{x})$, $g^{(s)}(\mathbf{x})$ (where $s,\ell\in\{0,1\}$), $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{s,0}<2N_{1}$ ($s=0,1$), $0\leq b_{s,1}<Q_{1}$ ($s=0,1$), $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{2}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f^{(s,\ell)}(\mathbf{x})=b_{s,\ell}\text{ for }s,\ell\in\{0,1\},\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{2}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $7\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
  • $q$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{2}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma V.2, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemmas V.3 and V.4 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma V.5 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{2}$ can correct two $b$-burst-deletions.
By Lemma V.1, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $U(\mathbf{x})_{1}$ is $d$-regular. Now the claim about redundancy follows from the pigeonhole principle. ∎
Remark V.2
Let $q=2$. Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $7\log n+(14b+11)\log\log n+O_{b}(1)$. Comparing this with Theorem IV.1, we can see that when $2\leq b\leq 16$, the code in Theorem IV.1 has lower redundancy, while when $b\geq 17$, the code in Theorem V.1 has lower redundancy.
VI Non-binary Two-deletion Correcting Codes
In this section, we always assume $q>2$. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.4 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23]
Lemma VI.1
[23] Suppose $q>2$ and $n>2$. There is a function $\mathsf{DVT}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+\log q}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathsf{DVT}(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma VI.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. Define
$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(I_{j}\right)\pmod{2N_{1}},$ (27)
$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(I_{j}\right)\pmod{Q}.$
The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ will help to deal with case (1) in Lemma IV.4. For case (2) in Lemma IV.4, let $P=3d\log n$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
Since $\left|J_{1}\right|\leq 2P=6d\log n$, we can view each $\xi_{1}\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{7\left\lceil\log q\right\rceil\log(6d\log n)+O_{q}(1)}$ (see Lemma II.4). For $s\in\{0,1\}$, let $K_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$ and define
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{0}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{1}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}.\end{array}$ (28)
Now we are ready to give our code.
Theorem VI.1
Suppose $q>2$. Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{1},c_{2}<N_{2}$, define the code $\mathcal{C}_{3}$ as
$\mathcal{C}_{3}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h_{e}(\mathbf{x})=c_{1},h_{o}(\mathbf{x})=c_{2}\end{array}\right\}$
Then $\mathcal{C}_{3}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{1}$ and $c_{2}$, such that the redundancy of $\mathcal{C}_{3}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
  • $q$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
The proof of Theorem VI.1 is similar to those of Theorems IV.1 and V.1, and thus omitted.
VII Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $7\log n+O(\log\log n)$ for all $b\geq 1$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. In particular, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$. Our technique fails in this setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence. We leave this problem for future research.
The following Lemma .1 is derived by slightly modifying the proof of [43, Proposition 3.3.7]. Before showing its proof, we list some facts.
Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}-\cdots$.
Fact .1
For all $10\leq x<$ 1, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
Proof:
When $10\leq x<$ 1, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$. Since $-\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. At last, it is easy to see that $-\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$. When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$. ∎
Lemma .1
Suppose that $q\geq 2$ is fixed and $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. Then it holds that
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Proof:
It is easy to verify that
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $=-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$ (29)
$=-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
$=1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
$=1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $10<\epsilon q<$ 1. Then it follows from .1 that $-\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $-\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$. Now by Equation 29, we have
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $\leq 1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
$=1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
$=1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$
$=1-\frac{\epsilon^{2}q^{2}}{2(q-1)\ln q}+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
$\overset{(a)}{\leq}1-\frac{\epsilon^{2}}{4},$
where (a) follows from the fact that $\epsilon<\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)$. ∎

### v3

@@
 Abstract
-In this paper, we explore constructions for codes that correct two bursts of deletions, with each burst having length exactly $b$.
-Previously, the best known construction, derived using the syndrome compression technique, achieved a redundancy of at most $7\log n+O\left(\log n/\log\log n\right)$ bits.
-In this work, we present new constructions for all $q\geq 2$ that achieve redundancy at most $7\log n+O(\log\log n)$ when $b>1$.
-Additionally, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
+In this paper, we study codes that correct two bursts of deletions, with each burst having length exactly $b$, where $b>1$.
+Previously, the best known construction, derived using the syndrome compression technique, achieved at most $7\log n+O\left(\log n/\log\log n\right)$ redundant bits.
+In this work, we develop new idea to construct $q$-ary codes that achieve redundancy at most $6\log n+O(\log\log n)$ for all $b>1$ and $q\geq 2$.
+At last, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
 Index Terms:
@@
 A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$.
-In this paper, we construct codes which outperform the one derived by the syndrome compression technique.
+This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes.
+On the other hand, [17, 19] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique.
+Specifically, [17] gave a binary code with redundancy at most $7\log n+O(1)$ and [19] gave a binary code with redundancy at most $4\log n+O(\log\log n)$.
+Note that in a burst of deletions, all deletions occur consecutively.
+Therefore, intuitively, it is reasonable to deem that there is no big difference between two-deletion correcting codes and codes for correcting two $b$-burst-deletions (where $b>1$).
+This raises a natural question: for $b>1$, is there a construction which leads to codes which are as good as or even better than the one given by the syndrome compression technique.
+Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$.
 Our contributions include:
-  • A binary code of length $n$ with redundancy at most $7\log n+(15b-5)\log\log n+O(1)$;
-  • A $q$-ary code of length $n$ with redundancy at most $7\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
-  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $7\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
+  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of $2$-$b$-burst-deletion correcting codes;
+  • A binary code of length $n$ with redundancy at most $6\log n+(15b-5)\log\log n+O(1)$;
+  • A $q$-ary code of length $n$ with redundancy at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
+  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
 Here, it is assumed that $q$ and $b$ are constants with respect to $n$.
@@
   1. (D1)
-    the two bursts happen sequentially: from $\mathbf{x}$, we get $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then we get $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$ from $\mathbf{z}$;
+    the two bursts are caused by two channels: $\mathbf{x}$ passes the first channel, resulting in $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then $\mathbf{z}$ passes the second channel, resulting in $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$;
   2. (D2)
-    the two bursts happen simultaneously: from $\mathbf{x}$, we get $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n].
-This definition is motivated by the nanopore sequencing technology, in which a DNA strand pass a nanoscale hole nucleotide by nucleotide [36, 37].
-So we can assume that the two bursts do not overlap.
-In fact, (D1) and (D2) are equivalent. Firstly, it is clear that see that (D1) covers (D2). Next, we show that (D2) also covers (D1).
+    the two bursts are caused by a single channel: symbols in $\mathbf{x}$ pass a channel sequentially and we receive $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n].
+In fact, (D1) and (D2) are equivalent. Firstly, it is clear that (D1) covers (D2). Next, we show that (D2) also covers (D1).
+Remark II.1
+There is another possibility: the two bursts might overlap and result in a single burst that is shorter than $2b$.
+We do not take this situation into account, since it is covered by a more comprehensive problem: correcting two bursts of deletions, where each burst has length at most $b$.
+Our idea in this paper fails in this situation.
+We left this problem for future research.
 Observation II.1
@@
 ∎
-Remark II.1
-There is another possibility: the two bursts might overlap and result in a single burst that is shorter than $2b$.
-We do not take this situation into account, since it is covered by a more comprehensive problem: correcting two bursts of deletions, where each burst has length at most $b$.
-Our idea in this paper fails in this situation.
-We left this problem for future research.
 For $t\in\{1,2\}$ and $n>tb$, define
@@
 Taking $f=\psi$ gives a function satisfying (P1) as needed.
-Before proceeding to subsequent sections, we introduce a well-known result from number theory, which will be frequently used.
+Before proceeding to subsequent sections, we introduce two useful lemmas.
+The first one is a well-known result from number theory, which will be frequently used.
+The second one will be used in Sections IV-A and V-A.
 Lemma II.6 (Bertrand’s Postulate/Bertrand–Chebyshev theorem)
 For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
+Let $N>2$ be an integer.
+Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$.
+Let ????
+denote an unknown symbol (not in $[0,N-1]$).
+If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two).
+For our purpose in Sections IV-A and V-A, assume $j\geq i+2$.
+For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
+Lemma II.7
+Let $Q\geq N$ be the smallest prime. For all $0\leq a_{1},a_{2},b_{1},b_{2}<Q$ and $0\leq c<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
+  • (C1)
+    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\equiv a_{1}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\equiv a_{2}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\equiv b_{1}\pmod{Q}$,
+    $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\equiv b_{2}\pmod{Q}$;
+  • (C2)
+    $W(\mathbf{x})\equiv c\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+2N\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
+Then the code $\mathcal{C}$ can correct two bursts of erasures.
+In particular, there is a function $\mathcal{E}_{1}:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+6\log N+4}$, computable in linear time, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\mathcal{E}_{1}(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
+Proof:
+Firstly, if the correctness of the code $\mathcal{C}$ is proved, we can define
+$\mathcal{E}_{1}(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\pmod{Q},\right.$
+$\left.\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\pmod{Q},W(\mathbf{x})\pmod{nN^{2}}\right).$
+Here, we view $\mathcal{E}(\mathbf{x})$ as a binary vector.
+By Lemma II.6, we have $N\leq Q<2N$.
+Now it is easy to see that the length of $\mathcal{E}(\mathbf{x})$ is at most $\log n+6\log N+4$.
+It remains to prove the correctness of $\mathcal{C}$.
+Suppose $\mathbf{y}$ is obtained from a codeword $\mathbf{x}\in\mathcal{C}$ by two bursts of erasures.
+Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$.
+Suppose the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$.
+Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$.
+This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$.
+By definition, the two bursts in $\mathbf{x}$ do not overlap.
+It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$.
+To decode $\mathbf{x}$ from $\mathbf{y}$. We first calculate the two erased symbols in each row. Notice that $N>2$ and hence $Q$ is an odd prime. Suppose $\sigma_{1},\sigma_{2}\in[0,N-1]$ satisfy
+$\sigma_{1}+\sigma_{2}\equiv t_{1}\pmod{Q},$ (1)
+$\sigma_{1}^{2}+\sigma_{2}^{2}\equiv t_{2}\pmod{Q}$
+where $t_{1}$ and $t_{2}$ are given.
+Then we can uniquely recover the (multi)set $\{\sigma_{1},\sigma_{2}\}$ by solving system (1).
+Now we can determine the erased symbols in $A(\mathbf{y})_{1}$ and $A(\mathbf{y})_{2}$ by (C1).
+Suppose that the two erased symbols in $A(\mathbf{y})_{1}$ are $\alpha_{1},\alpha_{2}$, and the two erased symbols in $A(\mathbf{y})_{2}$ are $\alpha_{3},\alpha_{3}$.
+It remains to determine positions of $\alpha_{1}$ and $\alpha_{2}$ in $A(\mathbf{x})_{1}$, and positions of $\alpha_{3}$ and $\alpha_{4}$ in $A(\mathbf{x})_{2}$.
+We can assume $\alpha_{1}\neq\alpha_{2}$ or $\alpha_{3}\neq\alpha_{4}$, since otherwise, we can uniquely recover $\mathbf{x}$.
+There are at most four choices for the positions of $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ and for each choice, we get a candidate for $\mathbf{x}$.
+We need to show that among these candidates, there is exactly one equals $\mathbf{x}$.
+Firstly, by the relationship of $\mathbf{x}$ and $\mathbf{y}$, it is clear that there is at least candidate equals $\mathbf{x}$.
+Next, we will show there is at most one such candidate.
+To that end, let $(\beta_{1},\beta_{2})$ be a permutation of $(\alpha_{1},\alpha_{2})$, and $(\beta_{3},\beta_{4})$ be a permutation of $(\alpha_{3},\alpha_{4})$, where $(\beta_{1},\beta_{2})\neq(\alpha_{1},\alpha_{2})$ or $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$.
+Let $\mathbf{u}$ be obtained from $\mathbf{y}$ by assigning $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively.
+Let $\mathbf{v}$ be obtained from $\mathbf{y}$ by assigning $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively.
+By assumption, we have $\mathbf{u}\neq\mathbf{v}$.
+We need to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ is not a codeword in $\mathcal{C}$.
+Let $\Delta=W(\mathbf{u})-W(\mathbf{v})$.
+Then we have
+$\Delta=i_{1}(\alpha_{1}-\beta_{1})+i_{2}(\alpha_{2}-\beta_{2})+2N\left(i_{3}(\alpha_{3}-\beta_{3})+i_{4}(\alpha_{4}-\beta_{4})\right).$
+To complete the proof, it is sufficient to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ does not satisfies (C2).
+Suppose on the contrary that $W(\mathbf{u})\equiv c\equiv W(\mathbf{v})\pmod{nN^{2}}$.
+Then we have $\Delta\equiv 0\pmod{nN^{2}}$.
+If $(\beta_{1},\beta_{2})=(\alpha_{1},\alpha_{2})$, it holds that $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$.
+Therefore, we conclude that $\alpha_{3}\neq\alpha_{4}$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$.
+Then it follow that $\left|\Delta\right|=2N(i_{4}-i_{3})\left|\alpha_{3}-\alpha_{4}\right|$ and hence $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)<nN^{2}$, which is a contradiction.
+Similarly, if $(\beta_{3},\beta_{4})=(\alpha_{3},\alpha_{4})$, we have $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)<nN^{2}$, which is a contradiction.
+Now suppose $\alpha_{1}\neq\alpha_{2}$, $\alpha_{3}\neq\alpha_{4}$, $(\beta_{1},\beta_{2})=(\alpha_{2},\alpha_{1})$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$.
+Then we have $\Delta=(i_{2}-i_{1})(\alpha_{2}-\alpha_{1})+2N(i_{4}-i_{3})(\alpha_{4}-\alpha_{3})$.
+Without loss of generality, suppose $\alpha_{1}<\alpha_{2}$.
+If $\alpha_{3}<\alpha_{4}$, then $0<\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)+\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)\leq n(N^{2}-1)$, which is a contradiction.
+Now suppose $\alpha_{3}>\alpha_{4}$.
+Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$.
+When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to verify that $0<-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(2N(N-1)-1)+2N(N-1)<nN^{2}$, which is a contradiction.
+When $i_{4}-i_{3}=i_{2}-i_{1}-1$, we have $i_{2}\geq i_{1}+2$.
+Then it follows that
+$0<2N(\alpha_{3}-\alpha_{4})-2(\alpha_{2}-\alpha_{1})\leq-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor\left(2N(N-1)-1\right)-2N<nN^{2},$
+which is a contradiction. ∎
 III Bounds
@@
 For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$.
-Then it follows from [39, page 100, Theorem 1] that
+Then it follows from [37, page 100, Theorem 1] that
 $M_{q,n,b}\geq\sum_{\mathbf{x}\in\Sigma_{q}^{n}}\frac{1}{d(\mathbf{x})+1}.$ (2)
@@
 ∎
-In [40, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes.
+In [38, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes.
 Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$.
@@
 $\cup_{1\leq\ell\neq k\leq\frac{n}{b}}\left\{A(\ell,k)_{i}\right\}=\mathcal{B}_{2}(A(\mathbf{x})_{i}),$
-for all $1\leq i\leq b$. By [41, eq. (11)], we know that
+for all $1\leq i\leq b$. By [39, eq. (11)], we know that
 $\binom{r-1}{2}\leq\left|\mathcal{B}_{2}(\mathbf{v})\right|\leq\binom{r+1}{2}$
@@
 =$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
-where (a) follows from the well-known result (see the proof of [42, Theorem 3.1]):
+where (a) follows from the well-known result (see the proof of [40, Theorem 3.1]):
 |$\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|=\binom{m}{j-1}q(q-1)^{j-1}.$
-Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [43, Proposition 3.3.3] that
+Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [41, Proposition 3.3.3] that
 $\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\leq q^{mH_{q}(t/m)}=q^{mH_{q}(1-1/q-\epsilon)},$ (6)
@@
 $H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
-Now it follows from Equations 4 and 5 that
+Now it follows from Equations 5 and 6 that
 $\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}\overset{(b)}{<}(1.121)^{3b}\frac{q^{n}}{n^{3}},$ (7)
-where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 3 and 6. ∎
+where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 4 and 7. ∎
 Then next corollary is a direct consequence of Theorems III.1 and III.2.
@@
 Our process of decoding $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$ is divided in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
-For the first step, for all $1\leq j<r\left(A(\mathbf{x})_{1}\right)$, define
+For the first step, for all $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
 $f_{j}(\mathbf{x})=\left(A(\mathbf{x})_{2,p_{j}},\ldots,A(\mathbf{x})_{b,p_{j}}\right),$
-which can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{b-1}$. For $s=0,1$, let
-$K_{s}=\left\{2\leq j\leq r\left(A(\mathbf{x})_{1}\right):j\equiv s\pmod{2}\right\}.$
-Let $Q_{1}\geq\max\{n/b,N_{1}\}$ be the smallest prime. For $s=0,1$, Define
-$f^{(s,0)}(\mathbf{x})=\sum_{j\in K_{s}}f_{j-1}(\mathbf{x})\pmod{2N_{1}},$ (8)
-$f^{(s,1)}(\mathbf{x})=\sum_{j\in K_{s}}(j-1)f_{j-1}(\mathbf{x})\pmod{Q_{1}}.$
-In the following, we may assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4.
+which can be viewed as an integer in $\left[0,N-1\right]$, where $N=2^{b-1}$.
+For uniform discussion, we also let $f_{0}(\mathbf{x})=0$.
+Now define $L(\mathbf{x})=f_{0}(\mathbf{x})f_{1}(\mathbf{x})\cdots f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})\in[0,N-1]^{r\left(A(\mathbf{x})_{1}\right)+1}$.
+In other word, sequence $L(\mathbf{x})$ is the concatenation of $f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\ldots,f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})$.
+To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$.
+By Equation 8, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
+In the following, we always assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4.
+Now we have $j_{2}-1\geq j_{1}-1+2$.
+Then it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures.
+To determine the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$, we apply Lemma II.7.
+Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7.
+Define
+$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
+Since $r(A(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+1)+6\log N+4}=2^{\log n+O_{b}(1)}$.
+Now the following lemma follows from above analysis and Lemma II.7.
 Lemma IV.5
@@
 Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known.
-Then given $f^{(s,0)}(\mathbf{x})$ and $f^{(s,1)}(\mathbf{x})$ ($s=0,1$), we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
-Proof:
-To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$.
-By Equation 7, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
-When $j_{1}=1$, we have $p_{j_{1}-1}=0$ and hence $A(\mathbf{x})_{p_{j_{1}-1}}$ is not defined.
-So we only need to find $A(\mathbf{x})_{p_{j_{2}-1}}$.
-Let $s\in\{0,1\}$ be such that $j_{2}\equiv s\pmod{2}$.
-If $s=1$, then $f_{j-1}(\mathbf{x})$ is known for all $j\in K_{s}\setminus\{j_{2}\}$.
-Let $\delta=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$.
-Then it follows from the first line of Equation 8 that
-$f_{j_{2}-1}(\mathbf{x})\equiv\delta\pmod{2N_{1}}.$
-Since $0\leq f_{j_{2}-1}(\mathbf{x})<N_{1}$ and $0\leq\delta<2N_{1}$, we conclude that $f_{j_{2}-1}(\mathbf{x})=\delta$.
-If $s=0$, then $j_{1}+1\equiv s\pmod{2}$.
-Since $j_{2}-j_{1}\geq 2$, we have $j_{1}+1\neq j_{2}$.
-Then $f_{j-1}(\mathbf{x})$ is known for all $j\in K_{s}\setminus\{j_{1}+1,j_{2}\}$.
-Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$.
-Then it follows from Equation 8 that
-$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$ (9)
-$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (10)
-Since $0\leq f_{j_{1}}(\mathbf{x}),f_{j_{2}-1}(\mathbf{x})<N_{1}$ and $0\leq\delta^{(0)}<2N_{1}$, we conclude from Equation 9 that $f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})=\delta^{(0)}$.
-Combining this with Equation 10, we get
-$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (11)
-$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
-Since $0\leq j_{1}<j_{2}-1<r\left(A(\mathbf{x})_{1}\right)\leq n/b\leq Q_{1}$, we have $j_{1}\not\equiv j_{2}-1\pmod{Q_{1}}$, which implies that system (11) can be uniquely solved and the unique solution is $\left(f_{j_{1}}(\mathbf{x})\pmod{Q_{1}},f_{j_{2}-1}(\mathbf{x})\pmod{Q_{1}}\right)$.
-Since $0\leq f_{j_{1}}(\mathbf{x}),f_{j_{2}-1}(\mathbf{x})<N_{1}\leq Q_{1}$, we have $f_{j_{1}}(\mathbf{x})=f_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $f_{j_{2}-1}(\mathbf{x})=f_{j_{2}-1}(\mathbf{x})\pmod{Q_{1}}$.
-Now assume $j_{1}>1$.
-First, suppose $j_{1}\equiv j_{2}\pmod{2}$.
-Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$, where $s\in\{0,1\}$ be such that $j_{1}\equiv j_{2}\equiv s\pmod{2}$.
-Then it follows from Equation 8 that
-$f_{j_{1}-1}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
-$(j_{1}-1)f_{j_{1}-1}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
-Similar to previous case, we can uniquely solve $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$ from this equation system.
-Now suppose $j_{1}\equiv s_{1}\pmod{2}$ and $j_{2}\equiv s_{2}\pmod{2}$, where $s_{1},s_{2}\in\{0,1\}$ and $s_{1}\neq s_{2}$.
-We first show how to calculate $f_{j_{2}-1}(\mathbf{x})$.
-Since $j_{1}\not\equiv s_{2}\pmod{2}$, it holds that $j_{1}+1\equiv s_{2}\pmod{2}$.
-Note that $j_{1}+1\neq j_{2}$ since $j_{2}-j_{1}\geq 2$ by assumption.
-It is clear that $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}$.
-Let $\delta^{(0)}=\left(f^{(s_{2},0)}(\mathbf{x})-\sum_{j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s_{2},1)}(\mathbf{x})-\sum_{j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$.
-Then it follows from Equation 8 that
-$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
-$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
-Similarly, we can obtain $f_{j_{2}-1}(\mathbf{x})$ by solving this equation system.
-It remains to show how to calculate $f_{j_{1}-1}(\mathbf{x})$.
-If $j_{2}=r\left(A(\mathbf{x})_{1}\right)$, then $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{1}}\setminus\{j_{1}\}$.
-Therefore,
-$f_{j_{1}-1}(\mathbf{x})=\left(f^{(s_{1},0)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}.$
-If $j_{2}<r\left(A(\mathbf{x})_{1}\right)$, then $j_{2}+1\equiv j_{1}\pmod{2}$.
-It is clear that $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}$.
-Let $\delta^{(0)}=\left(f^{(s_{1},0)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s_{1},1)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$.
-Then it follows from Equation 8 that
-$f_{j_{1}-1}(\mathbf{x})+f_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
-$(j_{1}-1)f_{j_{1}-1}(\mathbf{x})+j_{2}f_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
-We can obtain $f_{j_{1}-1}(\mathbf{x})$ by solving this equation system. ∎
+Then given $f(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
 Upon knowing the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$, it remains to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
@@
 So $g_{j}(\mathbf{x})$ can be viewed as an integer in $[0,N_{2}-1]$, where $N_{2}=2^{(b-1)\log\log(n)+O_{b}(1)}$. Here, $O_{b}(1)$ is a constant dependent only on $b$.
-Let $Q_{2}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
+Let $Q_{1}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
 $g^{(0)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}g_{j}(\mathbf{x})\pmod{2N_{2}},$ (9)
-$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{2}}$
+$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{1}}$
 Lemma IV.6
@@
 Since the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are given, the value of $g_{j}(\mathbf{x})$ is known for all $j\neq j_{1},j_{2}$.
-Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{2}}$.
+Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{1}}$.
 Then it follows from Equation 9 that
 $g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{2}},$ (10)
-$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{2}}.$ (14)
+$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (11)
 Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 10 that $g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})=\delta^{(0)}$.
 Combining this with Equation 11, we get
-$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{2}},$ (15)
-$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{2}}.$
-Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{2}$, system (15) can be uniquely solved and the unique solution is
-$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{2}},g_{j_{2}}(\mathbf{x})\pmod{Q_{2}}\right).$
-Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{2}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{2}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{2}}$.
+$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (12)
+$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
+Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{1}$, system (12) can be uniquely solved and the unique solution is
+$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{1}},g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}\right).$
+Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{1}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}$.
 ∎
@@
 Suppose $n>2b$ and $b>1$.
-Let $N_{1},N_{2},N_{3},Q_{1},Q_{2}$, $f^{(s,\ell)}(\mathbf{x})$, $g^{(s)}(\mathbf{x})$ (where $s,\ell\in\{0,1\}$), $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above.
-For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{s,0}<2N_{1}$ ($s=0,1$), $0\leq b_{s,1}<Q_{1}$ ($s=0,1$), $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{2}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
-$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f^{(s,\ell)}(\mathbf{x})=b_{s,\ell}\text{ for }s,\ell\in\{0,1\},\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
+Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above.
+For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
+$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
 Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code.
-Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $7\log n+(15b-5)\log\log n+O_{b}(1)$.
+Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $6\log n+(15b-5)\log\log n+O_{b}(1)$.
 Proof:
@@
 By Lemma IV.2, there are at least $2^{n-1}$ $7$-regular sequences when $n\geq 9$.
-It then follows from the pigeonhole principle that there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
-$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot(2N_{1})^{2}\cdot(Q_{1})^{2}\cdot(2N_{2})\cdot Q_{2}\cdot(N_{3})^{2}},$
+It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
+$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot N_{1}\cdot(2N_{2})\cdot Q_{1}\cdot(N_{3})^{2}},$
 which implies by definition that
-$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+2\log Q_{1}+\log Q_{2}+2\log(2N_{1})+\log(2N_{2})+2\log N_{3}+1$
-$=4\log n+2\log Q_{1}+\log Q_{2}+(15b-5)\log\log n+O_{b}(1)$
+$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+\log(2N_{2})+\log Q_{1}+2\log N_{3}+1$
+$=5\log n+\log Q_{1}+(15b-5)\log\log n+O_{b}(1)$
 $\leq 6\log n+(15b-5)\log\log n+O_{b}(1).$
-The last inequality follows from the fact $n/b\leq Q_{1},Q_{2}<2n/b$, which is guaranteed by Lemma II.6. ∎
+The last inequality follows from the fact $n/b\leq Q_{1}<2n/b$, which is guaranteed by Lemma II.6. ∎
+IV-C Another construction when $b=2$
+Recall that in Section IV-A, the decoding of $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ is conducted in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
+This is because the interval $\left[p_{j-1},p_{j}\right]$ intersects both $I_{j-1}$ and $I_{j}$.
+A natural question is whether we can handle deletions in the interval $\left[p_{j-1},p_{j}\right]$ in one step, instead of two steps.
+Next, we show that this is possible when $b=2$.
+Recall that in Section IV-A, once the first row $A(\mathbf{x})_{1}$ is recovered, the two deletions in the second row $A(\mathbf{x})_{2}$ are determined in two non-overlapping intervals $\left[p_{j_{1}-1},p_{j_{1}}\right]$ and $\left[p_{j_{2}-1},p_{j_{2}}\right]$ in $[n/2]$ if we assume $p_{j_{2}-1}>p_{j_{1}}$ (otherwise, the two deletions can be corrected using the method in Section IV-B).
+In Lemmas 21, 22 and 23 in [42], Lu et al proved the follows result, which is a straightforward corollary of Lemma 5 and Observation 2 in [19].
+Lemma IV.8
+[42, Lemma 23] Let $m$, $P_{1}$ and $P_{2}$ be positive integers, where $P_{1}+P_{2}<m$.
+Let $P=\min\{P_{1},P_{2}\}$.
+There is a function $\mathcal{E}_{2}:\Sigma_{2}^{m}\rightarrow\Sigma_{2}^{2\log m+\log P+O(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{m}$, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two bits at two known non-overlapping intervals of length at most $P_{1}$ and $P_{2}$ respectively, then given $\mathcal{E}_{2}(\mathbf{x})$ and $\mathbf{y}$, one can uniqely and efficently recover $\mathbf{x}$.
+Note that $p_{j_{1}}-p_{j_{1}-1},p_{j_{2}}-p_{j_{2}-1}\leq 7\log(n/2)+1$.
+So, let $P_{1}=P_{2}=7\log(n/2)+1$ and $m=n/2$.
+Then we can apply the function $\mathcal{E}_{2}$ in Lemma IV.8 to handle interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (and $\left[p_{j_{2}-1},p_{j_{2}}\right]$) in one step, introducing only $2\log n+\log\log n+O(1)$ additional redundant bits.
+Now replacing the six functions $f^{(s,\ell)}$, $g^{(}0)$ and $g^{(}1)$ in the definition of $\mathcal{C}_{1}$ by only one function $\mathcal{E}_{2}$, we have the following corollary.
+Corollary IV.1
+Let $b=2$ and $n\geq 9$. There is a binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $6\log n+25\log\log n+O(1)$.
 V Non-binary $2$-$b$-Burst-Deletion Correcting Codes
@@
 Our first goal is to recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$.
-Note that if $j=1$, then $p_{j-1}=0$.
-So we only need to focus on $\mathbf{x}_{I_{j}^{L}}$ for $j\geq 2$.
-For $s=0,1$, let
-$K_{s}=\left\{2\leq j\leq r\left(U(\mathbf{x})_{1}\right):j\equiv s\pmod{2}\right\}.$
 Recall that $\left|I_{j}^{L}\right|\leq b-1$.
-So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N_{1}-1]$, where $N_{1}=q^{b-1}$.
-Let $Q_{1}\geq\max\{n/b,N_{1}\}$ be the smallest prime.
-For $s=0,1$, Define
-$f^{(s,0)}(\mathbf{x})=\sum_{j\in K_{s}}\mathbf{x}_{I_{j}^{L}}\pmod{2N_{1}},$ (18)
-$f^{(s,1)}(\mathbf{x})=\sum_{j\in K_{s}}(j-1)\mathbf{x}_{I_{j}^{L}}\pmod{Q_{1}}.$
+So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N-1]$, where $N=q^{b-1}$.
+Note that $I_{1}^{L}=\emptyset$.
+For uniform discussion, we also let $\mathbf{x}_{I_{1}^{L}}=0$ and $\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}=\mathbf{x}_{\left[n-b+2,n\right]}$.
+Now define $L(\mathbf{x})=\mathbf{x}_{I_{1}^{L}}\cdots\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}}\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}\in[0,N-1]^{r\left(U(\mathbf{x})_{1}\right)+2}$.
+In other word, sequence $L(\mathbf{x})$ is the concatenation of $\mathbf{x}_{I_{1}^{L}},\ldots,\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}},\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}$.
+By Equation 14, the value of $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\notin\{j_{1},j_{1}+1,j_{2},j_{2}+1\}$.
+Since $j_{2}\geq j_{1}+2$, it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures.
+To determine the values of $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, we apply Lemma II.7.
+Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7.
+Define
+$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
+Since $r(U(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+2)+10\log N+4}=2^{\log n+O_{q,b}(1)}$.
+Now the following lemma follows from above analysis and Lemma II.7.
 Lemma V.3
 Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
-Given $f^{(0,0)}(\mathbf{x})$, $f^{(0,1)}(\mathbf{x})$, $f^{(1,0)}(\mathbf{x})$ and $f^{(1,1)}(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
-Proof:
-We first consider the case $j_{1}=1$, then $I_{j_{1}}^{L}=\emptyset$ and so we only need to determine $\mathbf{x}_{I_{j_{2}}^{L}}$.
-Let $s\in\{0,1\}$ be such that $j_{2}\equiv s\pmod{2}$.
-If $j_{1}\equiv s\pmod{2}$, then by Equation 17, $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{2}\}$.
-By the first line of Equation 18, we have $\mathbf{x}_{I_{j_{2}}^{L}}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ since $0\leq\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}$.
-If $j_{1}\not\equiv s\pmod{2}$, then $j_{1}+1\equiv s\pmod{2}$.
-Since $j_{2}-j_{1}\geq 2$, it holds that $j_{1}+1\neq j_{2}$.
-By Equation 17, $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{1}+1,j_{2}\}$.
-Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}(j-1)\mathbf{x}_{I_{j}^{L}}\right)\pmod{Q_{1}}$.
-Then it follows from Equation 18 that
-$\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{2N_{1}},$ (19)
-$j_{1}\mathbf{x}_{I_{j_{1}+1}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$ (20)
-Since $0\leq\mathbf{x}_{I_{j_{1}+1}^{L}},\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}$ and $0\leq\delta^{(0)}<2N_{1}$, we conclude from Equation 19 that $\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}=\delta^{(0)}$.
-Combining this with Equation 20, we get
-$\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{Q_{1}},$ (21)
-$j_{1}\mathbf{x}_{I_{j_{1}+1}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$
-Since $0\leq j_{1}<j_{2}-1<Q_{1}$, we can get the unique solution of (21) in the field $\mathbb{F}_{Q_{1}}$: $\left(\mathbf{x}_{I_{j_{1}+1}^{L}}\pmod{Q_{1}},\mathbf{x}_{I_{j_{2}}^{L}}\pmod{Q_{1}}\right)$.
-Since $0\leq\mathbf{x}_{I_{j_{1}+1}^{L}},\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}\leq Q_{1}$, it holds that $\mathbf{x}_{I_{j_{1}+1}^{L}}=\mathbf{x}_{I_{j_{1}+1}^{L}}\pmod{Q_{1}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}=\mathbf{x}_{I_{j_{2}}^{L}}\pmod{Q_{1}}$.
-In particular, we obtained $\mathbf{x}_{I_{j_{2}}^{L}}$.
-Next, we consider the case $j_{1}>1$.
-Let $s_{1},s_{2}\in\{0,1\}$ be such that $j_{1}\equiv s_{1}\pmod{2}$ and $j_{2}\equiv s_{2}\pmod{2}$.
-If $s_{1}=s_{2}=s$, then $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{1},j_{2}\}$.
-Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}(j-1)\mathbf{x}_{I_{j}^{L}}\right)\pmod{Q_{1}}$.
-Then it follows from Equation 18 that
-$\mathbf{x}_{I_{j_{1}}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{2N_{1}},$
-$(j_{1}-1)\mathbf{x}_{I_{j_{1}}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$
-Similar to the case $j_{1}=1$, we can uniquely solve $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$.
-Now suppose $s_{1}\neq s_{2}$.
-In this case, we have $j_{1}+1\equiv s_{2}\pmod{2}$.
-Similar to the case when $j_{1}=1$ and $j_{1}\not\equiv j_{2}\pmod{2}$, we can obtain $\mathbf{x}_{I_{j_{2}}^{L}}$.
-It remains to show how to obtain $\mathbf{x}_{I_{j_{1}}^{L}}$.
-It is clear that $j_{2}+1\equiv s_{1}\pmod{2}$.
-Now similar to the case when $j_{1}=1$ and $j_{1}\not\equiv j_{2}\pmod{2}$, we can obtain $\mathbf{x}_{I_{j_{1}}^{L}}$.
-∎
+Given $f(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
 Once we know $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, it remains to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, where $I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b+1,j_{1}b\right]$ and $I_{j_{1}}^{\prime}=\left[p_{j_{2}-1}b+1,j_{2}b\right]$.
@@
 Then by Lemma II.2, $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{\log(db\log n)+O_{q,b}(1)}$.
-Let $Q_{2}\geq\max\{n/b,N_{2}\}$ be the smallest prime.
+Let $Q_{1}\geq\max\{n/b,N_{2}\}$ be the smallest prime.
 Define
 $g^{(0)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{2N_{2}},$ (15)
-$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{2}}.$
+$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{1}}.$
 Lemma V.4
@@
 Since $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known, it is easy to see that $\mathbf{x}_{I_{j}^{\prime}}$ is known for all $j\neq j_{1},j_{2}$.
-Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{2}}$.
+Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{1}}$.
 Then it follows from Equation 15 that
 $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{2N_{2}},$ (16)
-$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{2}}.$ (24)
+$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$ (17)
 Since $0\leq\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right),\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 16 that $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)=\delta^{(0)}$.
 Combining this with Equation 17, we get
-$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{2}},$ (25)
-$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{2}}.$
+$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{1}},$ (18)
+$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$
 Now similar to Lemma IV.6, we can uniquely solve $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$ from (18). ∎
@@
 Suppose $q\geq 2$, $n>2b$ and $b>1$.
-Let $N_{1},N_{2},N_{3},Q_{1},Q_{2}$, $f^{(s,\ell)}(\mathbf{x})$, $g^{(s)}(\mathbf{x})$ (where $s,\ell\in\{0,1\}$), $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above.
-For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{s,0}<2N_{1}$ ($s=0,1$), $0\leq b_{s,1}<Q_{1}$ ($s=0,1$), $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{2}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
-$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f^{(s,\ell)}(\mathbf{x})=b_{s,\ell}\text{ for }s,\ell\in\{0,1\},\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
+Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above.
+For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
+$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
 Then $\mathcal{C}_{2}$ is a $2$-$b$-burst-deletion correcting code.
-Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $7\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
+Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
   • $q$ is even, $d=7$ and $n\geq 9$;
@@
 Let $q=2$.
-Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $7\log n+(14b+11)\log\log n+O_{b}(1)$.
+Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $6\log n+(14b+11)\log\log n+O_{b}(1)$.
 Comparing this with Theorem IV.1, we can see that when $2\leq b\leq 16$, the code in Theorem IV.1 has lower redundancy, while when $b\geq 17$, the code in Theorem V.1 has lower redundancy.
@@
 Define
-$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(I_{j}\right)\pmod{2N_{1}},$ (27)
-$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(I_{j}\right)\pmod{Q}.$
+$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{2N_{1}},$ (20)
+$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{Q}.$
 The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ will help to deal with case (1) in Lemma IV.4. For case (2) in Lemma IV.4, let $P=3d\log n$ and
@@
 In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $7\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique.
-In particular, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
+Furthermore, when $b=2$, we show that there is a binary code with redundancy at most $6\log n+O\left(\log\log n\right)$.
+Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
 A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$.
@@
 We leave this problem for future research.
-The following Lemma .1 is derived by slightly modifying the proof of [43, Proposition 3.3.7]. Before showing its proof, we list some facts.
+The following Lemma .1 is derived by slightly modifying the proof of [41, Proposition 3.3.7]. Before showing its proof, we list some facts.
 Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}-\cdots$.

### v4

@@
 Abstract
-In this paper, we study codes that correct two bursts of deletions, with each burst having length exactly $b$, where $b>1$.
-Previously, the best known construction, derived using the syndrome compression technique, achieved at most $7\log n+O\left(\log n/\log\log n\right)$ redundant bits.
-In this work, we develop new idea to construct $q$-ary codes that achieve redundancy at most $6\log n+O(\log\log n)$ for all $b>1$ and $q\geq 2$.
-At last, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
+In this paper, we investigate codes designed to correct two bursts of deletions, where each burst has a length of exactly $b$, where $b>1$.
+The previous best construction, achieved through the syndrome compression technique, had a redundancy of at most $7\log n+O\left(\log n/\log\log n\right)$ bits.
+In contrast, our work introduces a novel approach for constructing $q$-ary codes that attain a redundancy of at most $5\log n+O(\log\log n)$ bits for all $b>1$ and $q\geq 2$.
+Additionally, for the case where $b=1$, we present a new construction of $q$-ary two-deletion correcting codes with a redundancy of $5\log n+O(\log\log n)$ bits, for all $q>2$.
 Index Terms:
@@
 In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4].
-For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lowere bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6].
-This implies that codes in [3, 4] have redundancy optimal up to a constant.
+For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6].
+This implies that the VT code in [3] and the Tenengolts code [4] have redundancy optimal up to a constant.
 Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
-To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19].
+To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19, 24].
 For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20].
@@
 The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding.
-A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best known result in redundancy.
-When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log(n)+O(\log\log(n))$ [21].
-And a following work [22] gave a $q$-ary $2$-deletion correcting code with redundancy $5\log(n)+O(\log\log(n))$ for all $q>2$.
+A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy.
+When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21].
+And in a following work [22] the authors presented a $q$-ary $2$-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
 When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23].
@@
 If deletions occur at consecutive positions, we call them a burst of deletions.
-Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 24], wireless sensor networks and satellite communication devices [25].
+Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 25], wireless sensor networks, and satellite communication devices [26].
 A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions.
-In 1970, Levenshtein gave a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [26].
+In 1970, Levenshtein presented a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [27].
 For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [28], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [29].
@@
 It was proved in [29, 31] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$.
-Therefore, the codes in [26] and [31, Theorem 9] have redundancy optimal up to a constant.
+Therefore, the codes in [27] and [32, Theorem 9] have redundancy optimal up to a constant.
 There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 21, 35, 36, 32].
 In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$.
-To the best of our knowledge, there is no explicit results about such codes.
-A related result can be found in a work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively.
+We call such codes $2$-$b$-burst correcting codes.
+To the best of our knowledge, there are no explicit results about such codes.
+A related result can be found in a work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively.
 Let $t=2$ and $t_{L}=b$.
@@
 This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes.
-On the other hand, [17, 19] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique.
-Specifically, [17] gave a binary code with redundancy at most $7\log n+O(1)$ and [19] gave a binary code with redundancy at most $4\log n+O(\log\log n)$.
+On the other hand, [17, 19, 21, 22, 24] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique.
+Specifically, [17] and [19, 24] presented binary codes with redundancy at most $7\log n+O(1)$ and $4\log n+O(\log\log n)$, respectively.
+For non-binary codes, the best-known redundancy is $5\log n+O(\log\log n)$ [21, 22].
 Note that in a burst of deletions, all deletions occur consecutively.
-Therefore, intuitively, it is reasonable to deem that there is no big difference between two-deletion correcting codes and codes for correcting two $b$-burst-deletions (where $b>1$).
-This raises a natural question: for $b>1$, is there a construction which leads to codes which are as good as or even better than the one given by the syndrome compression technique.
+Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and $2$-$b$-burst-deletion correcting codes (where $b>1$).
+This raises a natural question: for $b>1$, is there a construction that leads to codes that are as good as, or even better than, the one given by the syndrome compression technique?
 Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$.
@@
   • We establish lower and upper bounds on the size (or equivalently, the redundancy) of $2$-$b$-burst-deletion correcting codes;
-  • A binary code of length $n$ with redundancy at most $6\log n+(15b-5)\log\log n+O(1)$;
-  • A $q$-ary code of length $n$ with redundancy at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
+  • A binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for all $b>1$;
+  • A $q$-ary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for all $q>2$ and $b>1$;
   • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
 Here, it is assumed that $q$ and $b$ are constants with respect to $n$.
+Therefore, our results show that for $2$-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
 The rest of this paper is organized as follows.
@@
 In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions.
-Section IV and Section V deal with binary and non-binary codes, respectively.
-In Section VI, we show that our idea in Section V can result in a new construction of non-binary two-deletion correcting code.
-Finally Section VII concludes this paper.
+Section IV deals with codes for correcting two $b$-burst-deletions.
+In Section V, we give a new construction of non-binary two-deletion correcting code.
+Finally Section VI concludes this paper.
 II Preliminaries
-In this section, we introduce some necessary terminologies and notations.
+In this section, we first introduce some necessary definitions, auxiliary conclusions and related results.
 For an integer $q\geq 2$ and a positive integer $n$, denote $\Sigma_{q}=\left\{0,1,\ldots,q-1\right\}$ and $\Sigma_{q}^{n}$ the set of all $q$-ary sequences with $n$ symbols.
@@
     the two bursts are caused by a single channel: symbols in $\mathbf{x}$ pass a channel sequentially and we receive $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n].
-In fact, (D1) and (D2) are equivalent. Firstly, it is clear that (D1) covers (D2). Next, we show that (D2) also covers (D1).
 Remark II.1
@@
 We left this problem for future research.
+In fact, (D1) and (D2) are equivalent. Firstly, it is clear that (D1) covers (D2). Next, we show that (D2) also covers (D1).
 Observation II.1
@@
 Suppose $t\in\{1,2\}$.
-We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$.
+We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{t}^{b}(\mathbf{x})\cap\mathcal{B}_{t}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$.
 In particular, if $b=1$, we call $\mathcal{C}$ a $t$-deletion correcting code.
@@
 $A\left(\mathbf{x},n,n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
-We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix representation of $\mathbf{x}$.
+We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$.
 If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
@@
 $A\left(\mathbf{x},8,3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}.$
-In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as an array $A(\mathbf{x},n,b)$.
+In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},n,b)$.
 To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$.
 All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$.
-The next observation is a straightforward corollary of II.1.
+Throughout this paper, for a matrix $A$, denote by $A_{i}$ and $A_{i,j}$ the $i$-th row of $A$ and the entry in the $i$-th row and $j$-th column, respectively.
+The next observation is a straightforward result of II.1.
 Observation II.2
@@
 In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$.
-In Section VI, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
+In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
 Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available.
-A related result can be found in the work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively.
+A related result can be found in the work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively.
 For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$.
@@
 Lemma II.2
-[31, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
+[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
 Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
@@
 Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$.
-So we let $N=q^{b}n^{3}$.
+In fact, each codeword in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$ can be obtained in the following three steps:
+  1. 1)
+    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are at most $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
+  2. 2)
+    For each $\mathbf{z}^{(1)}$, insert a sequence of length $b$ into $\mathbf{z}^{(1)}$ and get a sequence $\mathbf{z}^{(2)}\in\Sigma_{q}^{n-b}$.
+For each $\mathbf{z}^{(1)}$, there are at most $q^{b}n$ possibilities for $\mathbf{z}^{(2)}$.
+  3. 3)
+    Insert a sequence of length $b$ into $\mathbf{z}^{(2)}$ to get a sequence $\mathbf{y}\in\Sigma_{q}^{n}$.
+Since $\mathcal{E}$ is a $1$-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
+Therefore, we can let $N=q^{b}n^{3}$.
 Now the redundancy of $\mathcal{E}^{\prime}$ is at most $7\log n+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$.
@@
 [17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}^{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
+This result can be extended to arbitrary finite alphabets in the following way.
 Lemma II.4
@@
 Taking $f=\psi$ gives a function satisfying (P1) as needed.
-Before proceeding to subsequent sections, we introduce two useful lemmas.
-The first one is a well-known result from number theory, which will be frequently used.
-The second one will be used in Sections IV-A and V-A.
+Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory. The second one will be used in Section IV-B.
 Lemma II.6 (Bertrand’s Postulate/Bertrand–Chebyshev theorem)
 For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
-Let $N>2$ be an integer.
+Let $N\geq 2$ be an integer.
 Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$.
@@
 If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two).
-For our purpose in Sections IV-A and V-A, assume $j\geq i+2$.
+For our purpose in Section IV-B, assume $j\geq i+2$.
 For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
 Lemma II.7
-Let $Q\geq N$ be the smallest prime. For all $0\leq a_{1},a_{2},b_{1},b_{2}<Q$ and $0\leq c<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
+For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
   • (C1)
-    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\equiv a_{1}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\equiv a_{2}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\equiv b_{1}\pmod{Q}$,
-    $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\equiv b_{2}\pmod{Q}$;
+    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\equiv a_{1}\pmod{2N}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\equiv a_{2}\pmod{2N}$;
   • (C2)
-    $W(\mathbf{x})\equiv c\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+2N\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
-Then the code $\mathcal{C}$ can correct two bursts of erasures.
-In particular, there is a function $\mathcal{E}_{1}:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+6\log N+4}$, computable in linear time, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\mathcal{E}_{1}(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
+    $W(\mathbf{x})\equiv b\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+(2N-1)\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
+Then the code $\mathcal{C}$ can correct two bursts (of length two) of erasures.
+In particular, there is a function $\varphi:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+4\log N+2}$, efficiently computable, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\varphi(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
 Proof:
 Firstly, if the correctness of the code $\mathcal{C}$ is proved, we can define
-$\mathcal{E}_{1}(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\pmod{Q},\right.$
-$\left.\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\pmod{Q},W(\mathbf{x})\pmod{nN^{2}}\right).$
-Here, we view $\mathcal{E}(\mathbf{x})$ as a binary vector.
-By Lemma II.6, we have $N\leq Q<2N$.
-Now it is easy to see that the length of $\mathcal{E}(\mathbf{x})$ is at most $\log n+6\log N+4$.
+$\varphi(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\pmod{2N},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N},W(\mathbf{x})\pmod{nN^{2}}\right).$
+Here, we view $\varphi(\mathbf{x})$ as a binary vector.
+There are at most $(2N)^{2}\cdot(nN^{2})$ values of $\varphi(\mathbf{x})$.
+As a result, the length of $\varphi(\mathbf{x})$, when viewed as a binary vector, is at most $\log n+4\log N+2$.
 It remains to prove the correctness of $\mathcal{C}$.
@@
 Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$.
-Suppose the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$.
+Denote the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$.
+Note that $i_{1},i_{2},i_{3}$ and $i_{4}$ are known to us.
 Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$.
 This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$.
-By definition, the two bursts in $\mathbf{x}$ do not overlap.
+Next, we describe how to decode $\mathbf{x}$ from $\mathbf{y}$.
+Clearly, it is sufficient to recover the values of $A(\mathbf{x})_{1,i_{1}}$, $A(\mathbf{x})_{1,i_{2}}$, $A(\mathbf{x})_{2,i_{3}}$ and $A(\mathbf{x})_{2,i_{4}}$.
+Let
+$\delta_{1}=\left(a_{1}-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\right)\pmod{2N},$ (1)
+$\delta_{2}=\left(a_{2}-\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\right)\pmod{2N}.$
+Since $0\leq A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}},A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}<2N$ and $0\leq\delta_{1},\delta_{2}<2N$, it follows from Condition (C1) and Equation 1 that $\delta_{1}=A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}}$ and $\delta_{2}=A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}$.
+For simpler notations, denote $\alpha_{1}=A(\mathbf{x})_{1,i_{1}}$ and $\alpha_{2}=A(\mathbf{x})_{2,i_{3}}$.
+Then we have $A(\mathbf{x})_{1,i_{2}}=\delta_{1}-\alpha_{1}$ and $A(\mathbf{x})_{2,i_{4}}=\delta_{2}-\alpha_{2}$.
+Therefore, it remains to obtain the values of $\alpha_{1}$ and $\alpha_{2}$.
+To that end, let
+$\Delta=\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}.$ (2)
+Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$.
+Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, it holds that
+$\Delta$ $=\left(W(\mathbf{x})-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{2,j}\right)\pmod{nN^{2}}$ (3)
+$=\left(i_{1}\alpha_{1}+i_{2}(\delta_{1}-\alpha_{1})+(2N-1)(i_{3}\alpha_{2}+i_{4}(\delta_{2}-\alpha_{2}))\right)\pmod{nN^{2}}.$
+Let $\Delta^{\prime}=\left(i_{2}\delta_{1}+(2N-1)i_{4}\delta_{2}-\Delta\right)\pmod{nN^{2}}$, which can be calculated since values of $\delta_{1}$, $\delta_{2}$ and $\Delta$ are known to us.
+Then it follows from Equation 3 that
+$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}\equiv\Delta^{\prime}\pmod{nN^{2}}.$ (4)
+Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $21\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/$ 2.
+Combining this with the fact $10\leq\alpha_{1},\alpha_{2}\leq N-$ 1, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$.
+Now Equation 4 implies
+$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}=\Delta^{\prime}.$ (5)
+Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$.
+By definition, the two bursts of erasures in $\mathbf{x}$ do not overlap.
 It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$.
-To decode $\mathbf{x}$ from $\mathbf{y}$. We first calculate the two erased symbols in each row. Notice that $N>2$ and hence $Q$ is an odd prime. Suppose $\sigma_{1},\sigma_{2}\in[0,N-1]$ satisfy
-$\sigma_{1}+\sigma_{2}\equiv t_{1}\pmod{Q},$ (1)
-$\sigma_{1}^{2}+\sigma_{2}^{2}\equiv t_{2}\pmod{Q}$
-where $t_{1}$ and $t_{2}$ are given.
-Then we can uniquely recover the (multi)set $\{\sigma_{1},\sigma_{2}\}$ by solving system (1).
-Now we can determine the erased symbols in $A(\mathbf{y})_{1}$ and $A(\mathbf{y})_{2}$ by (C1).
-Suppose that the two erased symbols in $A(\mathbf{y})_{1}$ are $\alpha_{1},\alpha_{2}$, and the two erased symbols in $A(\mathbf{y})_{2}$ are $\alpha_{3},\alpha_{3}$.
-It remains to determine positions of $\alpha_{1}$ and $\alpha_{2}$ in $A(\mathbf{x})_{1}$, and positions of $\alpha_{3}$ and $\alpha_{4}$ in $A(\mathbf{x})_{2}$.
-We can assume $\alpha_{1}\neq\alpha_{2}$ or $\alpha_{3}\neq\alpha_{4}$, since otherwise, we can uniquely recover $\mathbf{x}$.
-There are at most four choices for the positions of $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ and for each choice, we get a candidate for $\mathbf{x}$.
-We need to show that among these candidates, there is exactly one equals $\mathbf{x}$.
-Firstly, by the relationship of $\mathbf{x}$ and $\mathbf{y}$, it is clear that there is at least candidate equals $\mathbf{x}$.
-Next, we will show there is at most one such candidate.
-To that end, let $(\beta_{1},\beta_{2})$ be a permutation of $(\alpha_{1},\alpha_{2})$, and $(\beta_{3},\beta_{4})$ be a permutation of $(\alpha_{3},\alpha_{4})$, where $(\beta_{1},\beta_{2})\neq(\alpha_{1},\alpha_{2})$ or $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$.
-Let $\mathbf{u}$ be obtained from $\mathbf{y}$ by assigning $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively.
-Let $\mathbf{v}$ be obtained from $\mathbf{y}$ by assigning $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively.
-By assumption, we have $\mathbf{u}\neq\mathbf{v}$.
-We need to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ is not a codeword in $\mathcal{C}$.
-Let $\Delta=W(\mathbf{u})-W(\mathbf{v})$.
-Then we have
-$\Delta=i_{1}(\alpha_{1}-\beta_{1})+i_{2}(\alpha_{2}-\beta_{2})+2N\left(i_{3}(\alpha_{3}-\beta_{3})+i_{4}(\alpha_{4}-\beta_{4})\right).$
-To complete the proof, it is sufficient to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ does not satisfies (C2).
-Suppose on the contrary that $W(\mathbf{u})\equiv c\equiv W(\mathbf{v})\pmod{nN^{2}}$.
-Then we have $\Delta\equiv 0\pmod{nN^{2}}$.
-If $(\beta_{1},\beta_{2})=(\alpha_{1},\alpha_{2})$, it holds that $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$.
-Therefore, we conclude that $\alpha_{3}\neq\alpha_{4}$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$.
-Then it follow that $\left|\Delta\right|=2N(i_{4}-i_{3})\left|\alpha_{3}-\alpha_{4}\right|$ and hence $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)<nN^{2}$, which is a contradiction.
-Similarly, if $(\beta_{3},\beta_{4})=(\alpha_{3},\alpha_{4})$, we have $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)<nN^{2}$, which is a contradiction.
-Now suppose $\alpha_{1}\neq\alpha_{2}$, $\alpha_{3}\neq\alpha_{4}$, $(\beta_{1},\beta_{2})=(\alpha_{2},\alpha_{1})$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$.
-Then we have $\Delta=(i_{2}-i_{1})(\alpha_{2}-\alpha_{1})+2N(i_{4}-i_{3})(\alpha_{4}-\alpha_{3})$.
-Without loss of generality, suppose $\alpha_{1}<\alpha_{2}$.
-If $\alpha_{3}<\alpha_{4}$, then $0<\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)+\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)\leq n(N^{2}-1)$, which is a contradiction.
-Now suppose $\alpha_{3}>\alpha_{4}$.
-Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$.
-When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to verify that $0<-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(2N(N-1)-1)+2N(N-1)<nN^{2}$, which is a contradiction.
-When $i_{4}-i_{3}=i_{2}-i_{1}-1$, we have $i_{2}\geq i_{1}+2$.
-Then it follows that
-$0<2N(\alpha_{3}-\alpha_{4})-2(\alpha_{2}-\alpha_{1})\leq-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor\left(2N(N-1)-1\right)-2N<nN^{2},$
-which is a contradiction. ∎
+Therefore, we always have $i_{4}-i_{3}\geq 1$.
+This implies $(2N-1)(i_{4}-i_{3})>1$.
+Then it follows from Equation 5 that
+$(i_{2}-i_{1})\alpha_{1}\equiv\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}.$ (6)
+Note that $10\leq\alpha_{1}\leq N-$ 1.
+When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it it easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$.
+When $i_{4}-i_{3}=i_{2}-i_{1}-1$, since $i_{2}-i_{1}\geq 2$, we have $(2N-1)(i_{4}-i_{3})-(i_{2}-i_{1})\alpha_{1}=(i_{2}-i_{1})(2N-1-\alpha_{1})-(2N-1)\geq 2N-1-2\alpha_{1}>0$ and hence $(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$.
+Now it follows from Equation 6 that $(i_{2}-i_{1})\alpha_{1}=\Delta^{\prime\prime}$, where $\Delta^{\prime\prime}=\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$.
+From this, we get $\alpha_{1}=\frac{\Delta^{\prime\prime}}{(i_{2}-i_{1})}$.
+Then by Equation 5, we have $\alpha_{2}=\frac{\Delta^{\prime}-(i_{2}-i_{1})\alpha_{1}}{(2N-1)(i_{4}-i_{3})}$.
+Now the proof is completed.
+∎
 III Bounds
@@
 Let $\alpha(G)$ be the maximum size of an independent set of $G$.
-It is clear that a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$.
+A subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$.
 Therefore, we have $M_{q,n,b}=\alpha(G)$.
 For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$.
-Then it follows from [37, page 100, Theorem 1] that
+Then it follows from [38, page 100, Theorem 1] that
 $M_{q,n,b}\geq\sum_{\mathbf{x}\in\Sigma_{q}^{n}}\frac{1}{d(\mathbf{x})+1}.$ (7)
@@
 ∎
-In [38, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes.
+In [39, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes.
 Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$.
@@
 $\cup_{1\leq\ell\neq k\leq\frac{n}{b}}\left\{A(\ell,k)_{i}\right\}=\mathcal{B}_{2}(A(\mathbf{x})_{i}),$
-for all $1\leq i\leq b$. By [39, eq. (11)], we know that
+for all $1\leq i\leq b$. By [40, eq. (11)], we know that
 $\binom{r-1}{2}\leq\left|\mathcal{B}_{2}(\mathbf{v})\right|\leq\binom{r+1}{2}$
@@
 =$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
-where (a) follows from the well-known result (see the proof of [40, Theorem 3.1]):
+where (a) follows from the well-known result (see the proof of [41, Theorem 3.1]):
 |$\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|=\binom{m}{j-1}q(q-1)^{j-1}.$
-Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [41, Proposition 3.3.3] that
+Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [42, Proposition 3.3.3] that
 $\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\leq q^{mH_{q}(t/m)}=q^{mH_{q}(1-1/q-\epsilon)},$ (11)
 where $H_{q}(x)\triangleq x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function. Here, $\log_{q}x=\frac{\log x}{\log q}$ for any positive real number $x$.
-If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. So by Lemma .1, we have.
+If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma .1, we have
 $H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
-Now it follows from Equations 5 and 6 that
+Now it follows from Equations 10 and 11 that
 $\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}\overset{(b)}{<}(1.121)^{3b}\frac{q^{n}}{n^{3}},$ (12)
-where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 4 and 7. ∎
+where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 9 and 12. ∎
 Then next corollary is a direct consequence of Theorems III.1 and III.2.
@@
 where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code with maximum size.
-IV Binary $2$-$b$-Burst-Deletion Correcting Codes
-In this section, we will construct a class of binary $2$-$b$-burst-deletion correcting codes with redundancy at most $7\log n+O(\log\log n)$.
-First, we introduce some $1$-deletion correcting codes and $2$-deletion correcting codes, which will be useful later.
-Lemma IV.1
-[21, Lemma 1] For any integer $n\geq 3$, there exists a function $\mathsf{VT}:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{\log(n+1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\mathsf{VT}\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
+IV Codes For Correcting Two $b$-Burst-Deletions
+In this section, we construct $q$-ary $2$-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$.
+Our idea is to first locate positions of deletions in short intervals (which is accomplished in Section IV-A, Lemma IV.5) and then correct errors in these intervals (which is accomplished in Sections IV-B and IV-C).
+IV-A Approximately determine positions of deletions
 Definition IV.1 (regularity)
@@
 For the number of regular sequences, we have the following lemma.
-Lemma IV.2
+Lemma IV.1
 [19, Lemma 11] The number of $d$-regular sequences of length $n$ is at least $2^{n-1}$, as long as $d\geq 7$ and $n$ be such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$.
 In particular, when $d=7$, it suffices to require $n\geq 9$.
-Lemma IV.3
+Lemma IV.2
 [19, Theorem 7] Suppose $d\geq 7$.
@@
 If the first row $A(\mathbf{x})_{1}$ is regular, then by correcting the two deletions in $A(\mathbf{x})_{1}$, we can approximate the location of the two deletions, as ensured by the following lemma.
-Lemma IV.4
-[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$.
+Lemma IV.3
+[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$.
 Then we can
@@
     or find an interval $J\subseteq[n]$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, of length at most $3d\log n$ such that $i_{1},i_{2}\in J$.
-From now on, we assume that $A(\mathbf{x})_{1}$ is $7$-regular.
-Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$.
-Then by II.2, we have $A\left(\mathbf{x}^{\prime}\right)_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$.
-Lemma IV.3 ensures that when given $\eta\left(A(\mathbf{x})_{1}\right)$, one can uniquely recover $A(\mathbf{x})_{1}$ from $A\left(\mathbf{x}^{\prime}\right)_{1}$.
-Then by Lemma IV.4, we can know the approximate positions of deletions in $A(\mathbf{x})_{1}$.
-This further enables us to approximately determine the positions of deletions in $A(\mathbf{x})_{i}$ for all $2\leq i\leq b$.
-Based on the two cases in Lemma IV.4, we split our discussion into Section IV-A and Section IV-B.
-Our codes for correcting two $b$-burst-deletions is presented in Theorem IV.1, whose proof is divided in into Lemmas IV.5, IV.6 and IV.7.
-IV-A When Case (1) in Lemma IV.4 happens
-Let $A(\mathbf{x})_{1,I_{j}}$ be the $j$th run of $A(\mathbf{x})_{1}$, where $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$ (recall that $r\left(A(\mathbf{x})_{1}\right)$ is the number of runs in $A(\mathbf{x})_{1}$) and $I_{j}\subseteq[n/b]$ is the index set of $A(\mathbf{x})_{1,I_{j}}$.
-For simplicity, denote $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$ and $p_{r\left(A(\mathbf{x})_{1}\right)}=n/b$.
-Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $p_{j}-p_{j-1}=\left|I_{j}\right|\leq 7\log(n/b)$.
-Note that $I_{j}$ depends on $\mathbf{x}$.
-We omit $\mathbf{x}$ for simpler notations.
-Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $j_{1}<j_{2}$.
-Then for all $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (if $j_{1}=1$, then this deletion occurred in the interval $[1,p_{1}]$) and the other occurred in the interval $\left[p_{j_{2}-1},p_{j_{2}}\right]$.
-So $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin\left[p_{j_{1}-1},p_{j_{1}}\right]\cup\left[p_{j_{2}-1},p_{j_{2}}\right]$.
-Specifically, we have
-$A(\mathbf{x})_{i,k}=\begin{cases}A(\mathbf{x}^{\prime})_{i,k},\text{ if }k<p_{j_{1}-1},\\ A(\mathbf{x}^{\prime})_{i,k-1},\text{ if }p_{j_{1}}<k<p_{j_{2}-1},\\ A(\mathbf{x}^{\prime})_{i,k-2},\text{ if }k>p_{j_{2}}.\end{cases}$ (8)
-for all $2\leq i\leq b$.
-Therefore, we need to decode $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$, for all $2\leq i\leq b$.
-Notice that when $j\geq 2$, the interval $[p_{j-1},p_{j}]$ intersects with both $I_{j-1}$ and $I_{j}$.
-Our process of decoding $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$ is divided in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
-For the first step, for all $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
-$f_{j}(\mathbf{x})=\left(A(\mathbf{x})_{2,p_{j}},\ldots,A(\mathbf{x})_{b,p_{j}}\right),$
-which can be viewed as an integer in $\left[0,N-1\right]$, where $N=2^{b-1}$.
-For uniform discussion, we also let $f_{0}(\mathbf{x})=0$.
-Now define $L(\mathbf{x})=f_{0}(\mathbf{x})f_{1}(\mathbf{x})\cdots f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})\in[0,N-1]^{r\left(A(\mathbf{x})_{1}\right)+1}$.
-In other word, sequence $L(\mathbf{x})$ is the concatenation of $f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\ldots,f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})$.
-To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$.
-By Equation 8, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
-In the following, we always assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4.
-Now we have $j_{2}-1\geq j_{1}-1+2$.
-Then it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures.
-To determine the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$, we apply Lemma II.7.
-Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7.
-Define
-$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
-Since $r(A(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+1)+6\log N+4}=2^{\log n+O_{b}(1)}$.
-Now the following lemma follows from above analysis and Lemma II.7.
-Lemma IV.5
-Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$.
-Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known.
-Then given $f(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
-Upon knowing the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$, it remains to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
-Recall that $j_{2}-j_{1}\geq 2$.
-This implies $p_{j_{2}-1}>p_{j_{1}}$ and thus for all $2\leq i\leq b$, the deletion in $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ does not affect $A(\mathbf{x})_{i,I_{j_{1}}}$.
-Therefore, we can conclude that $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$.
-So we can apply Lemma IV.1 to decode $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}$, respectively.
-For $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
-$g_{j}(\mathbf{x})=\left(\mathsf{VT}\left(A(\mathbf{x})_{2,I_{j}}\right),\cdots,\mathsf{VT}\left(A(\mathbf{x})_{b,I_{j}}\right)\right).$
-Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $\left|I_{j}\right|\leq 7\log(n/b)$. Then it follows from Lemma IV.1 that
-$\mathsf{VT}\left(A(\mathbf{x})_{i,I_{j}}\right)\in\{0,1\}^{\log\left(7\log(n/b)+1\right)}.$
-So $g_{j}(\mathbf{x})$ can be viewed as an integer in $[0,N_{2}-1]$, where $N_{2}=2^{(b-1)\log\log(n)+O_{b}(1)}$. Here, $O_{b}(1)$ is a constant dependent only on $b$.
-Let $Q_{1}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
-$g^{(0)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}g_{j}(\mathbf{x})\pmod{2N_{2}},$ (9)
-$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{1}}$
-Lemma IV.6
-Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$.
-Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known.
-If the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are known, then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
-Proof:
-As discussed above, we have $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$.
-So by Lemma IV.1, to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, it is sufficient to obtain the values of $g_{j_{1}}(\mathbf{x})$ and $g_{j_{2}}(\mathbf{x})$.
-Since the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are given, the value of $g_{j}(\mathbf{x})$ is known for all $j\neq j_{1},j_{2}$.
-Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{1}}$.
-Then it follows from Equation 9 that
-$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{2}},$ (10)
-$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (11)
-Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 10 that $g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})=\delta^{(0)}$.
-Combining this with Equation 11, we get
-$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (12)
-$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
-Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{1}$, system (12) can be uniquely solved and the unique solution is
-$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{1}},g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}\right).$
-Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{1}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}$.
-∎
-IV-B When Case (2) in Lemma IV.4 happens
-In this case, we can find an interval $J\subseteq[n/b]$ of length at most $21\log(n/b)$ such that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$.
-Suppose $J=[c,d]$.
-Then for all $2\leq i\leq b$, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c-1,d]$, which is of length at most $121\log n+$ 1.
-Let $P=21\log\left(n/b\right)+1$ and
-$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/bP\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/bP\right\rceil-1.\end{cases}$
-Let $\xi(\cdot)$ be the function defined in Lemma II.3. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
-$h_{j}\left(\mathbf{x}\right)=\left(\xi\left(A(\mathbf{x})_{2,J_{j}}\right),\xi\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi\left(A(\mathbf{x})_{b,J_{j}}\right)\right).$
-Since $\left|J_{j}\right|\leq 2P=42\log\left(n/b\right)+2$ for all $j$, each $h_{j}(\mathbf{x})$ is a binary vector of length at most $7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)$ and so we can view $h_{j}(\mathbf{x})$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)}$.
-Let
-$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ even}}h_{j}(\mathbf{x})\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ odd}}h_{j}(\mathbf{x})\pmod{N_{3}}.\end{array}$ (13)
-Lemma IV.7
-Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$, where $J\subseteq[n/b]$ is an interval of length $21\log(n/b)$ and is known to us.
-Given $h_{e}\left(\mathbf{x}\right)$ and $h_{o}\left(\mathbf{x}\right)$, one can uniquely recover $A(\mathbf{x})_{i}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
-Proof:
-We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$.
-Without loss of generality, we assume that $j_{0}$ is even.
-Then $A(\mathbf{x})_{i,J_{j}}$ and thus $h_{j}(\mathbf{x})$ is known to us for each even $j\neq j_{0}$.
-So it remains to recover $A(\mathbf{x})_{J_{j_{0}}}$.
-If $J_{j_{0}}=[c,d]$, it is clear that $A(\mathbf{x}^{\prime})_{i,[c,d-2]}=\mathcal{B}_{2}\left(A(\mathbf{x})_{i,[c,d]}\right)$, for all $2\leq i\leq b$.
-By Lemma II.3, it is suffices to obtain $h_{j_{0}}(\mathbf{x})$.
-By Equation 13, we have
-$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
-Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, it follows that
-$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
-Now the proof is completed. ∎
-Now we are ready to present our main result in this section.
-Theorem IV.1
-Suppose $n>2b$ and $b>1$.
-Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above.
-For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
-$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
-Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code.
-Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $6\log n+(15b-5)\log\log n+O_{b}(1)$.
-Proof:
-Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$.
-Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n,b)$.
-Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$.
-Since $A(\mathbf{x})_{1}$ is $7$-regular, by Lemma IV.3, we can recover $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ based on $\eta\left(A(\mathbf{x})_{1}\right)$.
-Now according to Lemmas IV.4, IV.5, IV.6 and IV.7, it is clear that $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code.
-Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$.
-By Lemma IV.2, there are at least $2^{n-1}$ $7$-regular sequences when $n\geq 9$.
-It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
-$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot N_{1}\cdot(2N_{2})\cdot Q_{1}\cdot(N_{3})^{2}},$
-which implies by definition that
-$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+\log(2N_{2})+\log Q_{1}+2\log N_{3}+1$
-$=5\log n+\log Q_{1}+(15b-5)\log\log n+O_{b}(1)$
-$\leq 6\log n+(15b-5)\log\log n+O_{b}(1).$
-The last inequality follows from the fact $n/b\leq Q_{1}<2n/b$, which is guaranteed by Lemma II.6. ∎
-IV-C Another construction when $b=2$
-Recall that in Section IV-A, the decoding of $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ is conducted in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
-This is because the interval $\left[p_{j-1},p_{j}\right]$ intersects both $I_{j-1}$ and $I_{j}$.
-A natural question is whether we can handle deletions in the interval $\left[p_{j-1},p_{j}\right]$ in one step, instead of two steps.
-Next, we show that this is possible when $b=2$.
-Recall that in Section IV-A, once the first row $A(\mathbf{x})_{1}$ is recovered, the two deletions in the second row $A(\mathbf{x})_{2}$ are determined in two non-overlapping intervals $\left[p_{j_{1}-1},p_{j_{1}}\right]$ and $\left[p_{j_{2}-1},p_{j_{2}}\right]$ in $[n/2]$ if we assume $p_{j_{2}-1}>p_{j_{1}}$ (otherwise, the two deletions can be corrected using the method in Section IV-B).
-In Lemmas 21, 22 and 23 in [42], Lu et al proved the follows result, which is a straightforward corollary of Lemma 5 and Observation 2 in [19].
-Lemma IV.8
-[42, Lemma 23] Let $m$, $P_{1}$ and $P_{2}$ be positive integers, where $P_{1}+P_{2}<m$.
-Let $P=\min\{P_{1},P_{2}\}$.
-There is a function $\mathcal{E}_{2}:\Sigma_{2}^{m}\rightarrow\Sigma_{2}^{2\log m+\log P+O(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{m}$, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two bits at two known non-overlapping intervals of length at most $P_{1}$ and $P_{2}$ respectively, then given $\mathcal{E}_{2}(\mathbf{x})$ and $\mathbf{y}$, one can uniqely and efficently recover $\mathbf{x}$.
-Note that $p_{j_{1}}-p_{j_{1}-1},p_{j_{2}}-p_{j_{2}-1}\leq 7\log(n/2)+1$.
-So, let $P_{1}=P_{2}=7\log(n/2)+1$ and $m=n/2$.
-Then we can apply the function $\mathcal{E}_{2}$ in Lemma IV.8 to handle interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (and $\left[p_{j_{2}-1},p_{j_{2}}\right]$) in one step, introducing only $2\log n+\log\log n+O(1)$ additional redundant bits.
-Now replacing the six functions $f^{(s,\ell)}$, $g^{(}0)$ and $g^{(}1)$ in the definition of $\mathcal{C}_{1}$ by only one function $\mathcal{E}_{2}$, we have the following corollary.
-Corollary IV.1
-Let $b=2$ and $n\geq 9$. There is a binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $6\log n+25\log\log n+O(1)$.
-V Non-binary $2$-$b$-Burst-Deletion Correcting Codes
+We briefly explain why Lemma IV.3 is helpful.
+Suppose $\mathbf{x}\in\Sigma_{2}^{n}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$.
+Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n-2b,b)$.
+Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}(A(\mathbf{x})_{i})$ for all $1\leq i\leq 2$ If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals.
+Then by II.2, we can approximate error positions in $A(\mathbf{x})_{i}$ within one or two short intervals, for all $2\leq i\leq b$.
+However, the above analysis is valid only when $\mathbf{x}$ is a binary sequence.
+Since we deal with binary and non-binary cases in a unified way, it is convenient to generalize the notion of regularity to general alphabets.
+To that end, we associate with any $q$-ary sequence a binary sequence.
 Let $q\geq 2$.
-In this section, we will extend our idea in Section IV to construct $q$-ary codes capable of correcting two $b$-burst-deletions.
-Surprisingly, our construction in this section gives rise to a binary code which is better than that in Theorem IV.1 when $b\geq 17$.
-The first step is to establish a result similar to Lemma IV.4 to approximately determine the error positions. To that end, we associate with any $q$-ary sequence a binary sequence.
 For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$.
 Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$.
-Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$.
+Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$ and when $q=2$, we have $u(\mathbf{x})=\mathbf{x}$.
 It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
-Remark V.1
+Remark IV.1
 Although any $x\in\Sigma_{q}$ can be decomposed as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, it might exist some $u\in\{0,1\}$ and $0\leq v<\left\lceil q/2\right\rceil$ such that $\left\lceil q/2\right\rceil u+v\notin\Sigma_{q}$.
@@
 But when $q$ is odd, we have $\left\lceil q/2\right\rceil+\left\lceil q/2\right\rceil-1=q\notin\Sigma_{q}$.
-We first generalize the concept of $d$-regular sequence to sequences over general alphabets.
-Definition V.1
+Definition IV.2
 A sequence $\mathbf{x}\in\Sigma_{q}^{n}$ is said to be $d$-regular if $u(\mathbf{x})$ is $d$-regular.
@@
 […]
-Lemma V.1
+Lemma IV.4
 Let $q\geq 2$.
@@
 On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2.
-Then it follows from Lemma IV.2 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
+Then it follows from Lemma IV.1 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
 Now suppose that $q\geq 3$ is odd.
-In this case, previous argument does not hold.
+In this case, the previous argument does not hold.
 This is because when $q$ is odd, $u_{x}=1$ and $v_{x}=(q-1)/2$, we have $\left\lceil q/2\right\rceil u_{x}+(q-1)/2=q\notin\Sigma_{q}$.
@@
 We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$.
-By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}^{3}\right)=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$.
+By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}\right)^{3}=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$.
 Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$.
 For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$.
-So follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
+So it follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
 For $\mathbf{v}\in\Sigma_{q}^{m}$, let $\overline{\mathbf{v}}\triangleq(q-v_{1})\cdots(q-v_{m})$.
@@
 =$\mathcal{R}^{\prime}=\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-k\left\lfloor n/m\right\rfloor}.\end{array}\right\}$
-It is easy to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus
+It is straightforward to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus
 $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}^{\prime}\right|$ $=\left|Q\right|^{k}q^{n-km}$
@@
 ∎
-In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$.
+Notations: In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$.
+Denote the $i$-th row of $A(\mathbf{x})$ and $U(\mathbf{x})$ by $A(\mathbf{x})_{i}$ and $U(\mathbf{x})_{i}$, respectively.
 It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$.
 Let
-$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\}.$
-Then $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
-Like in Section IV-A, let $U(\mathbf{x})_{1,I_{j}}$ ($1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$) be all runs in $U(\mathbf{x})_{1}$.
+$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\},$
+which is a subset of $[n]$. Then we have $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
+Recall that $r\left(U(\mathbf{x})_{1}\right)$ is the number of runs in $U(\mathbf{x})_{1}$.
+For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run.
 Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$.
-Recall that $I_{j}$ is an interval in $[n/b]$.
-For all $j\geq 1$ We associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$.
+By definition, $I_{j}$ is an interval in $[n/b]$.
+For all $j\geq 1$, we associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$.
 In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$.
@@
 Figure 1: Illustration of relationship between $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$.
-Lemma V.2
+Lemma IV.5
 Let $q\geq 2$, $b>1$ and $n>2b$.
@@
 Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular.
-Let $\eta(\cdot)$ be the function defined in Lemma IV.3.
+Let $\eta(\cdot)$ be the function defined in Lemma IV.2.
 Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can
@@
 So $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
-Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.3, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$.
-In addition, by Lemma IV.4, we can
+Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.2, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$.
+In addition, by Lemma IV.3, we can
   • either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $k_{1}\in I_{j_{1}}$ and $k_{2}\in I_{j_{2}}$;
@@
 By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
-Note that for the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$.
+For the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$.
 Now the proof is completed.
 ∎
-In the following, we always assume $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$.
-Furthermore, suppose that $A(\mathbf{x})_{1}$ is $d$-regular.
-We need to find conditions on $\mathbf{x}$ which enable us to recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
-As in Section IV, we split our discussion into Sections V-A and V-B.
-Our main result in this section is presented in Theorem V.1.
-V-A When Case (1) in Lemma V.2 happens
+From now on, we assume that $\mathbf{x}\in\Sigma_{q}^{n}$ and $A(\mathbf{x})_{1}$ is $d$-regular, where $d$ will be specified later.
+Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$.
+Then by II.2, we have $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$.
+Since $U(\mathbf{x})_{1}$ is regular, Lemma IV.2 ensures that when given $\eta\left(U(\mathbf{x})_{1}\right)$, one can uniquely recover $U(\mathbf{x})_{1}$ from $U\left(\mathbf{x}^{\prime}\right)_{1}$.
+Then by Lemma IV.3, we can know the approximate positions of the two deletions in $U(\mathbf{x})_{1}$.
+According to Lemma IV.5, we can approximately determine the positions of the two bursts in $\mathbf{x}$.
+Based on the two cases in Lemma IV.5, we split our discussion into Section IV-B and Section IV-C.
+Our codes for correcting two $b$-burst-deletions are presented in Theorem IV.1, whose proof is divided into Lemmas IV.6 and IV.7.
+Before proceeding, we need the following observation.
+Observation IV.1
+Let $m>b$.
+Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}\in\mathcal{B}_{1}^{b}(\mathbf{u})$.
+Then for any $b\leq i\leq m-b$, we have $\mathbf{v}_{[1,i-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[1,i]}\right)$ and $\mathbf{v}_{[i+1,m-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[i+1,m]}\right)$.
+IV-B When Case (1) in Lemma IV.5 happens
 In this case, we already found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
@@
 $x_{k}=\begin{cases}x^{\prime}_{k},\text{ if }k<p_{j_{1}-1}b-b+2,\\ x^{\prime}_{k-b},\text{ if }p_{j_{1}}b<k<p_{j_{2}-1}b-b+2,\\ x^{\prime}_{k-2b},\text{ if }p_{j_{2}}b<k\leq n.\end{cases}$ (13)
-We can assume $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma V.2.
-Our first goal is to recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$.
-Recall that $\left|I_{j}^{L}\right|\leq b-1$.
-So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N-1]$, where $N=q^{b-1}$.
-Note that $I_{1}^{L}=\emptyset$.
-For uniform discussion, we also let $\mathbf{x}_{I_{1}^{L}}=0$ and $\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}=\mathbf{x}_{\left[n-b+2,n\right]}$.
-Now define $L(\mathbf{x})=\mathbf{x}_{I_{1}^{L}}\cdots\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}}\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}\in[0,N-1]^{r\left(U(\mathbf{x})_{1}\right)+2}$.
-In other word, sequence $L(\mathbf{x})$ is the concatenation of $\mathbf{x}_{I_{1}^{L}},\ldots,\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}},\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}$.
-By Equation 14, the value of $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\notin\{j_{1},j_{1}+1,j_{2},j_{2}+1\}$.
-Since $j_{2}\geq j_{1}+2$, it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures.
-To determine the values of $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, we apply Lemma II.7.
-Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7.
-Define
-$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
-Since $r(U(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+2)+10\log N+4}=2^{\log n+O_{q,b}(1)}$.
-Now the following lemma follows from above analysis and Lemma II.7.
-Lemma V.3
+In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is know to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
+Therefore, to decode $\mathbf{x}$ from $\mathbf{x}^{\prime}$, it is sufficient to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
+In the rest of this subsection, it is always assumed that $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma IV.5 (or equivalently, case (2) in Lemma IV.3).
+Since $j_{2}-j_{1}\geq 2$, deletions in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ do not affect $\mathbf{x}_{I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}}$.
+As a result, (13) implies $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$.
+Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}\cup I_{j_{1}}}$, $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
+$\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}-1}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right),$ (14)
+$\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right).$
+Similarly, we have
+$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}-1}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right),$ (15)
+$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-1}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right).$
+Recall that the range of $j_{1}$ is $\left[1,r(U(\mathbf{x})_{1})\right]$.
+To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 00) and $\mathbf{x}^{\prime}_{\left[p_{-2}b+1,-b\right]}$ be an empty sequence.
+By (14) and (15), to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$, we need to apply to each $\mathbf{x}_{I_{j}^{\prime}}$ a code for correcting single $b$-burst-deletion.
+Let $\phi(\cdot)$ be the function defined in Lemma II.2.
+Then by Lemma II.2, our goal boils down to recovering the values of $\phi\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$.
+Since $A(\mathbf{x})_{1}$ is $d$-regular, we have $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$.
+Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $[0,N-1]$, where $N=2^{\log\log(n)+O_{q,d,b}(1)}$.
+Here, $O_{q,d,b}(1)$ denotes a constant dependent only on $q,d$ and $b$.
+Define $\overline{\phi}(\mathbf{x})=\phi\left(\mathbf{x}_{I_{0}^{\prime}}\right)\phi\left(\mathbf{x}_{I_{1}^{\prime}}\right)\cdots\phi\left(\mathbf{x}_{I_{r(U(\mathbf{x})_{1})}^{\prime}}\right)\in[0,N-1]^{r(U(\mathbf{x})_{1})+1}$.
+According to (13), when given $\mathbf{x}^{\prime}$, the value of $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
+Therefore, sequence $\overline{\phi}(\mathbf{x})$ suffered two bursts (of length two) of erasures.
+To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.7 and define
+$f(\mathbf{x})=\varphi\left(\overline{\phi}(\mathbf{x})\right).$
+Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.7, $f(\mathbf{x})$ can be viewed as an integer in $[0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
+The following lemma follows from the above analysis, Lemmas II.2 and II.7.
+Lemma IV.6
 Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
-Given $f(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
-Once we know $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, it remains to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, where $I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b+1,j_{1}b\right]$ and $I_{j_{1}}^{\prime}=\left[p_{j_{2}-1}b+1,j_{2}b\right]$.
-Since $j_{2}-j_{1}\geq 2$, we have $(p_{j_{2}-1}-1)b+2>p_{j_{1}}b$.
-This implies that the $b$-burst-deletion in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ does not affect $\mathbf{x}_{I_{j_{1}}^{\prime}}$.
-Therefore, we conclude that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$.
-So to decode $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}$, we need the function $\phi(\cdot)$ defined in Lemma II.2.
-Next, we will apply $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ to decode $\mathbf{x}_{I_{j}^{\prime}}$.
-Recall that $U(\mathbf{x})_{1}$ is $d$-regular.
-So by definition, we have $\left|I_{j}^{\prime}\right|=(p_{j}-p_{j-1})b\leq db\log n$.
-Then by Lemma II.2, $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{\log(db\log n)+O_{q,b}(1)}$.
-Let $Q_{1}\geq\max\{n/b,N_{2}\}$ be the smallest prime.
-Define
-$g^{(0)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{2N_{2}},$ (15)
-$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{1}}.$
-Lemma V.4
-Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
-Furthermore, suppose that $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known.
-Then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}$.
+Given $f(\mathbf{x})$, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
+IV-C When Case (2) in Lemma IV.5 happens
+Recall that case (2) in Lemma IV.5 corresponds to case (2) in Lemma IV.3.
+Therefore, for case (2), instead of looking at $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we look at $A(\mathbf{x})$ and $A(\mathbf{x}^{\prime})$.
+In this case, we have already decoded $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ by using $\eta\left(U(\mathbf{x})_{1}\right)$.
+Furthermore, by case (2) in Lemma IV.3, we found an interval $[c_{1},c_{2}]\subseteq[n/b]$ of length at most $3db\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$).
+By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c_{1}-1,c_{2}]$, for all $2\leq i\leq b$.
+Therefore, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$.
+Note that $\left|[c_{1}-1,c_{2}]\right|$=$\left|J\right|+1\leq 3db\log(n/b)+1$. Let $P=3db\log(n/b)+1$ and define
+$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/bP\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/bP\right\rceil-1.\end{cases}$
+Since $c_{1}$ and $c_{2}$ are known, we can find some $j_{0}$ such that $[c_{1}-1,c_{2}]\subseteq J_{j_{0}}$.
+Suppose $J_{j_{0}}=[d_{1},d_{2}]$.
+According to above discussion, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin J_{j_{0}}$.
+To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$.It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i,J_{j_{0}}}\right)$.
+Therefore, we can apply Lemma II.4 to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$.
+Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
+$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$
+Note that in the above definition of $g_{j}(\mathbf{x})$, we distinguish the two cases where $q>2$ and $q=2$.
+This is because when $q=2$, we have $A(\mathbf{x})=U(\mathbf{x})$, and hence the first row of $A(\mathbf{x})$ was already recovered.
+Since $\left|J_{j}\right|\leq 2P=6db\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $+O_{q,d,b}(1)$.
+Therefore, each $g_{j}(\mathbf{x})$ is a binary vector of length at most $7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)$ (when $q>2$) or $7(b-1)\log\log n+O_{d,b}(1)$ (when $q=2$).
+Therefore, we can view $g_{j}(\mathbf{x})$ as an integer in $\left[0,N_{2}-1\right]$, where
+$N_{2}=\begin{cases}2^{7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)},\mbox{ if }q>2,\\ 2^{7(b-1)\log\log n+O_{d,b}(1)},\mbox{ if }q=2.\\ \end{cases}$ (16)
+For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/bP\right\rceil-1:j\equiv s\pmod{2}\right\}$. For $s\in\{0,1\}$, define
+$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{s}}g_{j}(\mathbf{x})\pmod{N_{2}}.$ (17)
+Lemma IV.7
+For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for each $2\leq i\leq b$.
+In particular, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
 Proof:
-Since $j_{2}-j_{1}\geq 2$, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$.
-So by Lemma II.2, to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, it it sufficient to obtain $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$.
-Since $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known, it is easy to see that $\mathbf{x}_{I_{j}^{\prime}}$ is known for all $j\neq j_{1},j_{2}$.
-Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{1}}$.
-Then it follows from Equation 15 that
-$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{2N_{2}},$ (16)
-$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$ (17)
-Since $0\leq\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right),\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 16 that $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)=\delta^{(0)}$.
-Combining this with Equation 17, we get
-$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{1}},$ (18)
-$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$
-Now similar to Lemma IV.6, we can uniquely solve $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$ from (18). ∎
-V-B When Case (2) in Lemma V.2 happens
-In this case, we can find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$.
-This implies that $\mathbf{x}_{k}$ is known to us for all $k\notin J$.
-Suppose $J=[c_{1},c_{2}]$.
-Then $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}\in\mathcal{B}_{b}^{2}\left(\mathbf{x}_{J}\right)$.
-To recover $\mathbf{x}_{J}$, we apply a trivial $2$-$b$-burst-deletion correcting code of length $2\left|J\right|$, which is given in Lemma II.5.
-Let $P=3bd\log(n/b)+b-1$ and
-$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
-For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$.
-Let the function $\psi(\cdot)$ be defined in Lemma II.5.
-Since $\left|J_{j}\right|\leq 2P$, we can view each $\psi\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7b\left\lceil\log q\right\rceil\log(2P/b)+O_{q,b}(1)}$ and define
-$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{0}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{1}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}}.\end{array}$ (19)
-Lemma V.5
-Suppose that we found an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$.
-Then given $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$, we can uniquely recover $\mathbf{x}_{J}$ (and thus $\mathbf{x}$) from $\mathbf{x}^{\prime}$.
+Without loss of generality, assume $j_{0}\in L_{0}$.
+Then $g_{j}(\mathbf{x})$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$.
+By above discussion and Lemma II.4, to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ for all $2\leq i\leq b$, it is sufficient to obtain the value of $g_{j_{0}}(\mathbf{x})$.
+It follows from Equation 17, that
+$g_{j_{0}}(\mathbf{x})\equiv\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}.$
+Since $0\leq g_{j_{0}}(\mathbf{x})<N_{2}$, we have
+$g_{j_{0}}(\mathbf{x})=\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}.$
+Now the proof is completed. ∎
+We are ready to present our main result in this section.
+Theorem IV.1
+Suppose $q\geq 2$, $n>2b$ and $b>1$.
+Let $N_{1}$, $N_{2}$, $f(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above.
+Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$.
+For any $0\leq a<N_{0}$, $0\leq b<N_{1}$ and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{1}$ as
+$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
+Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code.
+Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O_{q,b}(1)$:
+  • $q>2$ is even, $d=7$ and $n\geq 9$;
+  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
+If $q=2$, $d=7$ and $n\geq 9$, the redundancy is upper bounded by $5\log n+14b\log\log n+O_{b}(1)$.
 Proof:
-We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$.
-Without loss of generality, we assume $j_{0}\in L_{0}$.
-Then $\psi\left(\mathbf{x}_{J_{j}}\right)$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$.
-In addition, if $J_{j_{0}}=[c_{1},c_{2}]$, it is clear that $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}=\mathcal{B}_{2}^{b}\left(\mathbf{x}_{J_{0}}\right)$.
-It follows from Equation 19, that
-$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
-Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, we have
-$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
-Now by Lemma II.5, one can uniquely recover $\mathbf{x}_{J_{0}}$ and thus $\mathbf{x}_{J}$. ∎
-Theorem V.1
-Suppose $q\geq 2$, $n>2b$ and $b>1$.
-Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above.
-For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
-$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
-Then $\mathcal{C}_{2}$ is a $2$-$b$-burst-deletion correcting code.
-Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
-  • $q$ is even, $d=7$ and $n\geq 9$;
-  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
-Proof:
-Suppose $\mathbf{x}\in\mathcal{C}_{2}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$.
-Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma V.2, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
-For the former case, Lemmas V.3 and V.4 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$.
-For the latter case, Lemma V.5 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$.
-Therefore, $\mathcal{C}_{2}$ can correct two $b$-burst-deletions.
-By Lemma V.1, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $U(\mathbf{x})_{1}$ is $d$-regular. Now the claim about redundancy follows from the pigeonhole principle. ∎
-Remark V.2
-Let $q=2$.
-Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $6\log n+(14b+11)\log\log n+O_{b}(1)$.
-Comparing this with Theorem IV.1, we can see that when $2\leq b\leq 16$, the code in Theorem IV.1 has lower redundancy, while when $b\geq 17$, the code in Theorem V.1 has lower redundancy.
-VI Non-binary Two-deletion Correcting Codes
-In this section, we always assume $q>2$.
+Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$.
+Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
+For the former case, Lemma IV.6 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$.
+For the latter case, Lemma IV.7 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$.
+Therefore, $\mathcal{C}_{1}$ can correct two $b$-burst-deletions.
+By Lemma IV.4, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $A(\mathbf{x})_{1}$ is $d$-regular.
+It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, and $c_{1}$, such that
+$\left|\mathcal{C}_{1}\right|\geq\frac{q^{n-1}}{N_{0}\cdot N_{1}\cdot(N_{2})^{2}},$
+which implies by definition that
+$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+2\log N_{2}+1$
+$=5\log n+14\log\log n+2\log N_{2}+O_{q,b}(1).$
+Now the claim for redundancy follows from (16). ∎
+V Non-binary Two-deletion Correcting Codes
+In this section, it is always assumed that $q>2$.
 We aim to give a new construction of $q$-ary two-deletion correcting codes.
@@
 We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular.
-Then Lemma IV.4 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions.
+Then Lemma IV.3 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions.
 This enables us to approximately determine error positions in $\mathbf{x}$.
@@
 We use the code given in [23]
-Lemma VI.1
+Lemma V.1
 [23] Suppose $q>2$ and $n>2$.
@@
 Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$.
-Therefore, by Lemma VI.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$.
+Therefore, by Lemma V.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$.
 Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime.
+By Lemma II.6, we have $n\leq Q<2n$.
 Define
@@
 $f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{Q}.$
-The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ will help to deal with case (1) in Lemma IV.4. For case (2) in Lemma IV.4, let $P=3d\log n$ and
+The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ can help to deal with case (1) in Lemma IV.3, as shown in the next lemma.
+Lemma V.2
+Suppose that the two deletions in $\mathbf{x}$ occurred in two known intervals $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$, respectively.
+Then given $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
+Proof:
+Notice that $x_{k}$ is known to us for all $k\notin I_{j_{1}}\cup I_{j_{2}}$.
+It remains to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$.
+It is easy to see that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}+1,p_{j_{1}}-1\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2},p_{j_{2}}-2\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{2}}}\right)$.
+By Lemma V.1, to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$, it is sufficient to know the values of $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)$.
+Note that $\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)$ is known to us for all $j\neq j_{1},j_{2}$.
+Let $\delta_{0}=\left(f_{0}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{2N_{1}}$ and $\delta_{1}=\left(f_{1}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{Q}$.
+Then it follows from Equation 18 that
+$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{2N_{1}},$ (19)
+$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}.$ (20)
+Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<2N_{1}$ and $0\leq\delta_{0}<2N_{1}$, we can conclude from Equation 19 that $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\delta_{0}$.
+Combining this with Equation 20, we get
+$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{Q},$ (21)
+$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}.$
+Since $1\leq j_{1}<j_{2}\leq n\leq Q$, we have $j_{1}\not\equiv j_{2}\pmod{Q}$.
+Therefore, system (21) has a unique solution in the field $\mathbb{F}_{Q}$: $\left(\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q},\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}\right)$.
+Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right),\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<N_{1}\leq Q$, we have $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q}$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}$.
+Now the proof is completed.
+∎
+For case (2) in Lemma IV.3, we follow similar idea in Section IV-C. Let $P=3d\log n$ and
 $J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
@@
 For $s\in\{0,1\}$, let $K_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$ and define
-$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{0}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{1}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}.\end{array}$ (21)
-Now we are ready to give our code.
-Theorem VI.1
+$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{s}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}.\$ (22)
+Lemma V.3
+Suppose we are in case (2) in Lemma IV.3. Given $h^{(0)}$ and $h^{(1)}$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
+The proof is similar to that of Lemma IV.7, and thus omitted.
+Our $q$-ary two-deletion correcting code is given in the next theorem.
+Theorem V.1
 Suppose $q>2$.
-Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above.
-For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{1},c_{2}<N_{2}$, define the code $\mathcal{C}_{3}$ as
-$\mathcal{C}_{3}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h_{e}(\mathbf{x})=c_{1},h_{o}(\mathbf{x})=c_{2}\end{array}\right\}$
-Then $\mathcal{C}_{3}$ is a two-deletion correcting code.
-Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{1}$ and $c_{2}$, such that the redundancy of $\mathcal{C}_{3}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
+Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above.
+For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{2}$ as
+$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
+Then $\mathcal{C}_{2}$ is a two-deletion correcting code.
+Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
   • $q$ is even, $d=7$ and $n\geq 9$;
   • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
-The proof of Theorem VI.1 is similar to those of Theorems IV.1 and V.1, and thus omitted.
-VII Conclusion
-In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $7\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique.
-Furthermore, when $b=2$, we show that there is a binary code with redundancy at most $6\log n+O\left(\log\log n\right)$.
+The proof of Theorem V.1 follows directly from Lemmas V.2 and V.3. The claim for redundancy follows from Lemma IV.4 and Lemma II.6.
+VI Conclusion
+In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $5\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique.
 Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
@@
 We leave this problem for future research.
-The following Lemma .1 is derived by slightly modifying the proof of [41, Proposition 3.3.7]. Before showing its proof, we list some facts.
+The following Lemma .1 is derived by slightly modifying the proof of [42, Proposition 3.3.7]. Before showing its proof, we list some facts.
 Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}-\cdots$.

### v5

@@
 In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4].
-For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6].
+For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|\mathcal{C}\right|\right)$ [^1] All logarithms in this paper are to the base $2$., is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6].
 This implies that the VT code in [3] and the Tenengolts code [4] have redundancy optimal up to a constant.
@@
 The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding.
-A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy.
-When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21].
-And in a following work [22] the authors presented a $q$-ary $2$-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
+A straightforward application of this method will give a $q$-ary $t$-deletion correcting code with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy.
+When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary 22-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21].
+And in a following work [22] the authors presented a $q$-ary 22-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
 When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23].
@@
 Therefore, the codes in [27] and [32, Theorem 9] have redundancy optimal up to a constant.
-There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 21, 35, 36, 32].
+There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 35, 21, 36, 37, 32].
+For readers’ convenience, we summarize previous results on codes correcting bursts of deletions in Table I.
 In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$.
-We call such codes $2$-$b$-burst correcting codes.
+We call such codes 22-$b$-burst correcting codes.
 To the best of our knowledge, there are no explicit results about such codes.
-A related result can be found in a work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively.
+A related result can be found in a work of Sima et al [34, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively ($(t,t_{L})$ burst deletions, for short).
 Let $t=2$ and $t_{L}=b$.
-Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$.
+Then their result gives a binary 22-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$.
 A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$.
@@
 Note that in a burst of deletions, all deletions occur consecutively.
-Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and $2$-$b$-burst-deletion correcting codes (where $b>1$).
+Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and 22-$b$-burst-deletion correcting codes (where $b>1$).
 This raises a natural question: for $b>1$, is there a construction that leads to codes that are as good as, or even better than, the one given by the syndrome compression technique?
@@
 Our contributions include:
-  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of $2$-$b$-burst-deletion correcting codes;
-  • A binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for all $b>1$;
-  • A $q$-ary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for all $q>2$ and $b>1$;
-  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
+  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of 22-$b$-burst-deletion correcting codes;
+  • A binary 22-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for any $b>1$;
+  • A $q$-ary 22-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for any $q>2$ and $b>1$;
+  • A new construction of $q$-ary two-deletion correcting codes of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for any $q>2$.
 Here, it is assumed that $q$ and $b$ are constants with respect to $n$.
-Therefore, our results show that for $2$-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
+Therefore, our results show that for 22-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
+TABLE I: Previous codes correcting bursts of deletions and corresponding methods
+Redundancies of $q$-ary Codes References      Core Methods
+single burst
+of size
+exactly $b$
+$b\log(n/b+1)$
+($q=2$)
+[28]
+1. representing each codeword as an array with $b$ rows
+2. imposing a VT constraint on each row
+$\log n+(b-1)\log\log n+O(1)$
+($q=2$, $q>2$)
+[29]
+[30]
+1. representing each codeword as an array with $b$ rows
+2. encoding the first row with a VT code with run-length limited constraint
+3. encoding the rest rows with shifted VT codes
+$\log n+O(1)$ ($q\geq 2$) [32]
+1. representing each codeword as an array with $b$ rows
+2. representing this array as a $q^{b}$-ary vector
+3. imposing two types of sum constraints on this vector
+4. consider the signature of this vector
+5. imposing a VT-type constraint together with three types of sum constraints
+   on the signature
+single burst
+of size
+at most $b$
+$\log n+1$ ($b=2,q=2$)
+[27]
+[36]
+1. imposing a sum constraint on rank sequences of codewords
+       or
+2. imposing a VT-type constraint on differential sequences of codewords
+$(b-1)\log n+\left(\binom{b}{2}-1\right)\log\log n$
+            +O(1)+O(1)
+($q=2$)
+[29]
+   extension of the construction of $b$-burst-deletion correcting codes
+   in the same work
+$\left\lceil\log b\right\rceil\left(\log n+\binom{b+1}{2}\log\log n\right)$
+           +O(1)+O(1)
+($q=2$)
+[33]
+1. denoting each $t\in[1,b]$ as $t=2^{i}\cdot j$, where $i\geq 0$ and $j$ is odd
+2. representing each code word as an array with $2^{i}$ rows
+3. imposing a VT-type constraint alongside a “balanced” constraint on the
+   first row to approximate the error positions
+4. representing each codeword as an array with $t$ rows and encoding
+   each row with a shifted VT code
+$\log n+\binom{b+1}{2}\log\log n+O(1)$
+($q=2$)
+[35]
+1. each codeword $\mathbf{x}$ is required to be $(\bm{p},\delta)$-dense
+2. associating with $\mathbf{x}$ a vector of integers $a_{\bm{p}}(\mathbf{x})$
+3. imposing a constraint on the number of $\bm{p}$ in $\mathbf{x}$ and a VT-type constraint
+   on $a_{\bm{p}}(\mathbf{x})$ to approximate positions of errors
+5. for each $1\leq t\leq b$, representing $\mathbf{x}$ as an array with $t$ rows and
+   encoding each row with shifted VT codes
+$4\log n+o\left(\log n\right)$ [34] syndrome compression technique
+$\log n+8\log\log n+o\left(\log\log n\right)$
+($q=2$)
+[21]
+1. applying the same method in [35] to approximate positions of errors
+2. applying a code with $4\log n+o(\log n)$ redundant bits to correct burst
+   deletions in short intervals
+$\log n+(8\log q+8)\log\log n$
+      $o\left(\log q\log\log n\right)$
+($q>2$ is even)
+1. representing each codeword as a binary array with $\left\lceil\log q\right\rceil$ rows
+2. encoding the first row with the binary code in the same work
+3. applying a code with $4\log n+o(\log n)$ redundant bits to correct
+   errors in remaining rows
+$\log n+8\log\log n+o\left(\log\log n\right)$
+($q\geq 2$)
+[37]    generalization of binary codes in [21]
+$\log n+\log q\log\log n+O(1)$
+($b=2$, $q>2$ is even)
+[36]
+1. representing each codeword as a binary array with $\left\lceil\log q\right\rceil$ rows
+2. encoding the first row with the binary code in the same work
+   with additional pattern-limited constraint
+3. applying a $P$-bounded version of the binary code to remaining rows
+$\log n+b\log\log n+O(1)$
+($q\geq 2$)
+[32]
+1. associating each codeword with a binary sequence
+2. applying the same method in [35] to the binary sequence to approximate
+   error positions
+3. for each $1\leq b^{\prime}\leq b$, applying a bounded $b^{\prime}$-burst-deletion correcting code
+   to correct errors in short intervals
+$(t,t_{L})$
+burst
+deletions
+$4t\log n+o(\log\log n)$
+($q=2$)
+[34]
+1. $t$-mixed sequences
+2. syndrome compression technique
 The rest of this paper is organized as follows.
@@
 Section IV deals with codes for correcting two $b$-burst-deletions.
-In Section V, we give a new construction of non-binary two-deletion correcting code.
+In Section V, we give a new construction of non-binary two-deletion correcting codes.
 Finally Section VI concludes this paper.
 II Preliminaries
-In this section, we first introduce some necessary definitions, auxiliary conclusions and related results.
+In this section, we introduce some necessary definitions, auxiliary conclusions and related results.
 For an integer $q\geq 2$ and a positive integer $n$, denote $\Sigma_{q}=\left\{0,1,\ldots,q-1\right\}$ and $\Sigma_{q}^{n}$ the set of all $q$-ary sequences with $n$ symbols.
@@
 The number of runs of $\mathbf{x}$ is denoted by $r(\mathbf{x})$.
-For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, $1$, 00 and $1$.
+For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, 11, 0 and $1$.
 So $r\left(\mathbf{x}\right)=5$.
@@
 Let $b$ and $n$ be two positive integers satisfying $b<n$.
-When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $11\leq i\leq n-b+$ 1.
+When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $1\leq i\leq n-b+1$.
 In this paper, we focus on codes correcting two $b$-burst-deletions. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$, where $n>2b$. There are two ways to define two $b$-burst-deletions:
@@
 Then there exist two intervals $I_{1}=[j_{1},j_{1}+b-1]$, $I_{2}=[j_{2},j_{2}+b-1]\subseteq[n]$, where $j_{2}-j_{1}\geq b$, such that $\mathbf{y}=\mathbf{x}_{[n]\setminus(I_{1}\cup I_{2})}$.
-In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlap substrings of length $b$ from $\mathbf{x}$.
+In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlapping substrings of length $b$ from $\mathbf{x}$.
 Proof:
-By assumption, there is $11\leq i_{1}\leq n-b+$ 1 and $11\leq i_{2}\leq n-2b+$ 1 such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$.
+By assumption, there is $1\leq i_{1}\leq n-b+1$ and $1\leq i_{2}\leq n-2b+1$ such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$.
 If $i_{2}\geq i_{1}$, let $j_{1}=i_{1}$ and $j_{2}=i_{2}+b$.
 Then the conclusion follows.
-Now suppose $11\leq i_{2}\leq i_{1}-$ 1.
+Now suppose $1\leq i_{2}\leq i_{1}-1$.
 If $i_{2}\leq i_{1}-b$, let $j_{1}=i_{2}$ and $j_{2}=i_{1}$.
@@
 For $t\in\{1,2\}$ and $n>tb$, define
-[…]
+$\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\Sigma_{q}^{n-tb}:\begin{array}[]{c}\mathbf{y}\text{ is obtained from }\mathbf{x}\\ \text{by }t\;b\text{-burst-deletion(s)}\end{array}\right\}.$
 When $b=1$, we use notation $\mathcal{B}_{t}(\mathbf{x})$ instead of $\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)$.
@@
 Here, “efficiently” means that the time complexity of decoding $\mathbf{x}$ from $\mathbf{y}$ is polynomial in $n$.
-In this paper, we construct $t$-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
+In this paper, we construct 22-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
 The redundancy of a code $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is defined to be $\rho(\mathcal{C})=\log\left(q^{n}/\left|\mathcal{C}\right|\right)$.
@@
 In addition, we always assume that $q$ and $b$ are fixed with respect to the code-length $n$.
-Let $n^{\prime}$ and $n$ be two positive integers satisfying $n^{\prime}<n$.
-For each sequence $\mathbf{x}\in\Sigma_{q}^{n}$, let $\widetilde{\mathbf{x}}$ be the zero padding of $\mathbf{x}$ to the shortest length that is greater than $n$ and is divisible by $n^{\prime}$, that is, $\widetilde{\mathbf{x}}=\mathbf{x}0^{\left\lceil n/n^{\prime}\right\rceil n^{\prime}-n}$, and then $n^{\prime}\mid\left|\widetilde{\mathbf{x}}\right|$.
-We can represent $\mathbf{x}$ as an $n^{\prime}\times\left\lceil n/n^{\prime}\right\rceil$ array $A\left(\mathbf{x},n,n^{\prime}\right)=\left[a_{i,j}\right]$, where $a_{i,j}=\tilde{x}_{i+n^{\prime}j}$ for all $1\leq i\leq n^{\prime}$ and $10\leq j\leq\left\lceil n/n^{\prime}\right\rceil-$ 1.
-In other words, the $i$-th row of $A\left(\mathbf{x},n,n^{\prime}\right)$ is
-$A\left(\mathbf{x},n,n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
-We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$.
-If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
+[…]
+$A\left(\mathbf{x},n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
+We call $A\left(\mathbf{x},n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$.
+If $n^{\prime}$ is clear from the context, we will denote $A\left(\mathbf{x},n^{\prime}\right)$ by $A(\mathbf{x})$.
 For example, let $n=7$, $\mathbf{x}=1011010\in\Sigma_{2}^{7}$ and $n^{\prime}=2$. Then $\widetilde{\mathbf{x}}=10110100\in\Sigma_{2}^{8}$ and
-$A\left(\mathbf{x},7,2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}.$
+$A\left(\mathbf{x},2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}$.
 If $n^{\prime}=3$, then $\widetilde{\mathbf{x}}=\mathbf{x}00$ and so
-$A\left(\mathbf{x},8,3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}.$
-In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},n,b)$.
+$A\left(\mathbf{x},3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}$.
+In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},b)$.
 To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$.
@@
 Observation II.2
-For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{2}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$.
+For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$.
 Any two $b$-burst-deletions in $\mathbf{x}$ will induce two deletions in each row of $A(\mathbf{x})$.
@@
 Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy.
-In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for all even $q>2$.
-In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$.
-In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
+In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for any even $q>2$.
+In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for any $q>2$.
+In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for any $q>2$.
 Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available.
-A related result can be found in the work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively.
-For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$.
+A related result can be found in the work of Sima et al [34, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively.
+For $t=2$ and $t_{L}=b$, their result provides a binary 22-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$.
 Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20].
@@
 $\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\mathcal{E}~:~\mathbf{y}\neq\mathbf{x}\text{ and }\mathcal{B}_{2}^{b}\left(\mathbf{y}\right)\cap\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)\neq\emptyset\right\}.$
+In other words, $\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$ is the set of all sequences (except $\mathbf{x}$) in $\mathcal{E}$ whose error-ball intersects with that of $\mathbf{x}$.
 Lemma II.1
@@
     if $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}\in\mathcal{N}_{\Sigma_{q}^{n}}\left(\mathbf{x}\right)$, then $f\left(\mathbf{x}\right)\neq f\left(\mathbf{y}\right)$.
-Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time, such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
-Let $\mathcal{E}$ be a $1$-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1.
+Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time [^2] In this paper, when saying that a function is computable in polynomial/linear time, we mean that this function is computable in time polynomial/linear in the code-length $n$.
+such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
+Let $\mathcal{E}$ be a 11-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1.
 Then Lemma II.1 asserts that if $\mathbf{x},\mathbf{y}\in\mathcal{E}$ are distinct codewords and $\bar{f}(\mathbf{x})=\bar{f}(\mathbf{y})$, we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$.
-Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code =$\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a $2$-$b$-burst-deletion correcting code.
-Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$ [^2] We select $\mathcal{E}$ to be a $1$-$b$-burst correcting code to obtain a better redundancy.
-If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$..
+Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code $\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a 22-$b$-burst-deletion correcting code.
+Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$.
+[^3] We select $\mathcal{E}$ to be a 11-$b$-burst correcting code to obtain a better redundancy.
+If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$.
 For the choice of $\mathcal{E}$, we have the following result.
 Lemma II.2
-[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
+[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n\geq b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
 Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
-This lemma gives a $1$-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$.
-Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$.
+This lemma gives a 11-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$.
+Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|<q^{b}n^{3}$.
 In fact, each codeword in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$ can be obtained in the following three steps:
   1. 1)
-    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are at most $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
+    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are less than $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
   2. 2)
@@
     Insert a sequence of length $b$ into $\mathbf{z}^{(2)}$ to get a sequence $\mathbf{y}\in\Sigma_{q}^{n}$.
-Since $\mathcal{E}$ is a $1$-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
+Since $\mathcal{E}$ is a 11-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
 Therefore, we can let $N=q^{b}n^{3}$.
@@
 Lemma II.3
-[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}^{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
+[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ (i.e., $\mathbf{y}$ is obtained from $\mathbf{x}$ by two deletions), one can uniquely and efficiently recover $\mathbf{x}$.
 This result can be extended to arbitrary finite alphabets in the following way.
@@
 $M(\mathbf{x})\triangleq\begin{pmatrix}x_{1,1}&\cdots&x_{1,n}\\ \vdots&\cdots&\vdots\\ x_{\left\lceil\log q\right\rceil,1}&\cdots&x_{\left\lceil\log q\right\rceil,n}\end{pmatrix}$,
-where $x_{k,i}\in\{0,1\}$ be such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
+where $x_{k,i}\in\{0,1\}$ such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
 Denote the $k$-th row of $M(\mathbf{x})$ by $M(\mathbf{x})_{k}$.
@@
 Proof:
-Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{y})=A(\mathbf{y},n-b,b)$.
+Let $A(\mathbf{x})=A(\mathbf{x},b)$ and $A(\mathbf{y})=A(\mathbf{y},b)$.
 Since $\mathbf{y}\in\mathcal{B}_{2}^{b}(\mathbf{x})$, we have $A(\mathbf{y})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$.
@@
 Taking $f=\psi$ gives a function satisfying (P1) as needed.
-Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory. The second one will be used in Section IV-B.
-Lemma II.6 (Bertrand’s Postulate/Bertrand–Chebyshev theorem)
-For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
+Before proceeding to subsequent sections, we introduce a useful lemma, which will be used in Section IV-B.
 Let $N\geq 2$ be an integer.
 Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$.
-Let ????
-denote an unknown symbol (not in $[0,N-1]$).
-If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two).
+Let ??
+denote an unknown symbol (not in [0,N1][0,N-1]).
+If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for any $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two).
 For our purpose in Section IV-B, assume $j\geq i+2$.
-For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
-Lemma II.7
-For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
+For a sequence $\mathbf{x}$ over the alphabet [0,N1][0,N-1], denote $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
+In the following lemma, denote $A(\mathbf{x})=A(\mathbf{x},2)$, i.e., the matrix representation of $\mathbf{x}$ with two rows.
+Lemma II.6
+For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be the set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
   • (C1)
@@
 Firstly, if the correctness of the code $\mathcal{C}$ is proved, we can define
-$\varphi(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\pmod{2N},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N},W(\mathbf{x})\pmod{nN^{2}}\right).$
+[…]
+$\quad\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N}$,
+$\quad W(\mathbf{x})\pmod{nN^{2}}\Bigg{)}$.
 Here, we view $\varphi(\mathbf{x})$ as a binary vector.
@@
 To that end, let
-$\Delta=\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}.$ (2)
-Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$.
-Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, it holds that
+$\Delta=$ $\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}\right$. (2)
+$\quad\quad\left.-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}$.
 $\Delta$ $=\left(W(\mathbf{x})-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{2,j}\right)\pmod{nN^{2}}$ (3)
 $=\left(i_{1}\alpha_{1}+i_{2}(\delta_{1}-\alpha_{1})+(2N-1)(i_{3}\alpha_{2}+i_{4}(\delta_{2}-\alpha_{2}))\right)\pmod{nN^{2}}$.
+Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$.
+Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, we have (3).
 Let $\Delta^{\prime}=\left(i_{2}\delta_{1}+(2N-1)i_{4}\delta_{2}-\Delta\right)\pmod{nN^{2}}$, which can be calculated since values of $\delta_{1}$, $\delta_{2}$ and $\Delta$ are known to us.
@@
 $(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}\equiv\Delta^{\prime}\pmod{nN^{2}}$. (4)
-Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $21\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/$ 2.
-Combining this with the fact $10\leq\alpha_{1},\alpha_{2}\leq N-$ 1, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$.
+Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $1\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/2$.
+Combining this with the fact $0\leq\alpha_{1},\alpha_{2}\leq N-1$, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$.
 Now Equation 4 implies
@@
 $(i_{2}-i_{1})\alpha_{1}\equiv\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$. (6)
-Note that $10\leq\alpha_{1}\leq N-$ 1.
-When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it it easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$.
+Note that $0\leq\alpha_{1}\leq N-1$.
+When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$.
 When $i_{4}-i_{3}=i_{2}-i_{1}-1$, since $i_{2}-i_{1}\geq 2$, we have $(2N-1)(i_{4}-i_{3})-(i_{2}-i_{1})\alpha_{1}=(i_{2}-i_{1})(2N-1-\alpha_{1})-(2N-1)\geq 2N-1-2\alpha_{1}>0$ and hence $(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$.
@@
 III Bounds
-We could not find any existing upper or lower bounds on the maximum size of a $2$-$b$-burst-deletion code. In this section, we will derive these bounds.
-Let $M_{q,n,b}$ be the maximum size of a $2$-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
+We could not find any existing upper or lower bounds on the maximum size of a 22-$b$-burst-deletion code. In this section, we will derive these bounds.
+Let $M_{q,n,b}$ be the maximum size of a 22-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
 Theorem III.1
-$M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
+The maximum size of a 22-$b$-burst-deletion correcting code satisfies $M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
 Proof:
@@
 Let $\alpha(G)$ be the maximum size of an independent set of $G$.
-A subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$.
+By definition, a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a 22-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$.
 Therefore, we have $M_{q,n,b}=\alpha(G)$.
 For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$.
-Then it follows from [38, page 100, Theorem 1] that
+Then it follows from [39, page 100, Theorem 1] that
 $M_{q,n,b}\geq\sum_{\mathbf{x}\in\Sigma_{q}^{n}}\frac{1}{d(\mathbf{x})+1}.$ (7)
@@
 ∎
-In [39, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes.
+In [40, Section IV-B], the authors proved an upper bound of 11-$b$-burst-deletion correcting codes.
 Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$.
@@
 For $q\geq 2$, let $f(q)=\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$.
-When $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
+If $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, the maximum size of a 22-$b$-burst-deletion correcting code satisfies
 $M_{q,n,b}\leq\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}.$
 Proof:
-Define $m=n/b-1$.
-Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a $2$-$b$-busrt-deletion correcting code.
+Define m=n/b1m=n/b-1.
+Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a 22-$b$-burst-deletion correcting code.
 Set $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and $t=\left(1-\frac{1}{q}-\epsilon\right)m$.
@@
 $\cup_{1\leq\ell\neq k\leq\frac{n}{b}}\left\{A(\ell,k)_{i}\right\}=\mathcal{B}_{2}(A(\mathbf{x})_{i}),$
-for all $1\leq i\leq b$. By [40, eq. (11)], we know that
+for all $1\leq i\leq b$. By [41, eq. (11)], we know that
 $\binom{r-1}{2}\leq\left|\mathcal{B}_{2}(\mathbf{v})\right|\leq\binom{r+1}{2}$
@@
 Since $\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ and $\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|=\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|$ for all $\mathbf{x}$, it follows that
-$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2},$ (8)
+$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|$ $\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ (8)
+$\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2}$,
 for all $\mathbf{x}\in\mathcal{C}_{1}$.
-Since $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a $2$-$b$-burst-deletion correcting code.
+Since $\mathcal{C}$ is a 22-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a 22-$b$-burst-deletion correcting code.
 So we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for all distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}_{1}$.
 Then it follows from Equation 8 that $\left|\mathcal{C}_{1}\right|\binom{t+1}{2}\leq\sum_{\mathbf{x}\in\mathcal{C}_{1}}\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\leq q^{n-2b}$ and hence
-$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ $=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$ (9)
+$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}$ $\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ (9)
+$=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$
 $=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2b^{2}}{(1-1/q-\epsilon)^{2}(1-b/n)^{2}}$
@@
 as long as $\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, which is possible when $n$ is sufficiently large.
-Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define a set
-$\mathcal{C}^{\prime}\triangleq\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$
+Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define $\mathcal{C}^{\prime}$ to be the following set
+$\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$
 Since $r\left(A(\mathbf{x})_{i}\right)\leq t+1$ for any $\mathbf{x}\in\mathcal{C}_{2}$ and $1\leq i\leq b$, we conclude that $\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|$.
 So it suffices to estimate an upper bound of $\left|\mathcal{C}^{\prime}\right|$.
-Since $m=n/b-1$, by definition of $\mathcal{C}^{\prime}$, we have
-$\left|\mathcal{C}^{\prime}\right|$ =|$=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (10)
+Since m=n/b1m=n/b-1, by definition of $\mathcal{C}^{\prime}$, we have
+$\left|\mathcal{C}^{\prime}\right|$ $=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (10)
 $=\left|\cup_{j=1}^{t+1}\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|^{b}$
-$=\left(\sum_{j=1}^{m+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
-=$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
-where (a) follows from the well-known result (see the proof of [41, Theorem 3.1]):
+$=\left(\sum_{j=1}^{t+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
+$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}$
+$=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
+where (a) follows from the well-known result (see the proof of [42, Theorem 3.1]):
 $\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|=\binom{m}{j-1}q(q-1)^{j-1}.$
-Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [42, Proposition 3.3.3] that
+Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [43, Proposition 3.3.3] that
 $\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\leq q^{mH_{q}(t/m)}=q^{mH_{q}(1-1/q-\epsilon)},$ (11)
 where $H_{q}(x)\triangleq x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function. Here, $\log_{q}x=\frac{\log x}{\log q}$ for any positive real number $x$.
-If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma .1, we have
+If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma A.1, we have
 $H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
 Now it follows from Equations 10 and 11 that
-$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}\overset{(b)}{<}(1.121)^{3b}\frac{q^{n}}{n^{3}},$ (12)
-where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 9 and 12. ∎
+$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}$ $\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}$ (12)
+$=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}$
+[…]
+where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $\sqrt[n]{n}<1.121$ when $n\geq 30$.
+We conclude the proof by combining Equations 9 and 12.
+∎
 Then next corollary is a direct consequence of Theorems III.1 and III.2.
@@
 When $n\geq\max\left\{30,\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}\right\}$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
-$2\log n+(2b-2)\log q+2\log(q-1)-2\log b-2\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
-where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code with maximum size.
+$2\log n+\log\left(\frac{q^{2b-2}(q-1)^{2}}{4b^{2}}\right)\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
+where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a 22-$b$-burst-deletion correcting code with maximum size.
+Proof:
+Firstly, the upper bound follows straightforward from Theorem III.1 just by noticing that $\binom{n}{2}^{2}\leq n^{4}$.
+Since $n\geq\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}$, we have $\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}\leq\frac{4b^{2}}{q^{2b-2}(q-1)^{2}}\frac{q^{n}}{n^{2}}$.
+Now the lower bound follows from Theorem III.2.
+∎
 IV Codes For Correcting Two $b$-Burst-Deletions
-In this section, we construct $q$-ary $2$-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$.
+In this section, we construct $q$-ary 22-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$.
 Our idea is to first locate positions of deletions in short intervals (which is accomplished in Section IV-A, Lemma IV.5) and then correct errors in these intervals (which is accomplished in Sections IV-B and IV-C).
@@
 In particular, when $d=7$, it suffices to require $n\geq 9$.
+The next lemma ensures that we can efficiently correct two deletions in $d$-regular sequences.
 Lemma IV.2
 [19, Theorem 7] Suppose $d\geq 7$.
-There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
+There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{x}^{\prime}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
 Here, $O_{d}(1)$ denotes a constant depending only on $d$.
-If the first row $A(\mathbf{x})_{1}$ is regular, then by correcting the two deletions in $A(\mathbf{x})_{1}$, we can approximate the location of the two deletions, as ensured by the following lemma.
+Furthermore, once $\mathbf{x}$ is recovered, locations of the two deletions in $\mathbf{x}$ can be approximated, as ensured by the following lemma.
 Lemma IV.3
 [21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$.
-Then we can
+When given $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we can
   1. $(1)$
@@
     or find an interval $J\subseteq[n]$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, of length at most $3d\log n$ such that $i_{1},i_{2}\in J$.
-We briefly explain why Lemma IV.3 is helpful.
+A proof of Lemma IV.3 is given in [21]. Here we show the intuition behind this lemma.
+Example IV.1
+  1. $(1)$
+    Let $\mathbf{x}=000111$ and $\mathbf{x}^{\prime}=0011$.
+Then $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by deleting one bit in the run $000$ and one bit in the run $111$.
+This corresponds to case (1) in Lemma IV.3.
+  2. $(2)$
+    $\mathbf{x}=10010101110$ and $\mathbf{x}^{\prime}=100101110$.
+Then $\mathbf{x}^{\prime}$ can be obtained from $\mathbf{x}$ by deleting $x_{2}$ and $x_{4}$, or by deleting $x_{3}$ and $x_{4}$, or by deleting $x_{4}$ and $x_{5}$, or by deleting $x_{7}$ and $x_{8}$, or by deleting $x_{7}$ and $x_{9}$.
+Therefore, we can only locate error positions in the substring $001010111$, which is the concatenation of a run $00$, an alternating substring $1010$ and a run $111$.
+If $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular, then each run has length at most $d\log n$ and each alternating substring has length at most $d\log n$.
+This is why in case (2) of Lemma IV.3 we can locate error positions in an interval of length at most $3d\log n$.
+Now we briefly explain why Lemma IV.3 is helpful.
 Suppose $\mathbf{x}\in\Sigma_{2}^{n}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$.
-Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n-2b,b)$.
-Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}(A(\mathbf{x})_{i})$ for all $1\leq i\leq 2$ If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals.
-Then by II.2, we can approximate error positions in $A(\mathbf{x})_{i}$ within one or two short intervals, for all $2\leq i\leq b$.
-However, the above analysis is valid only when $\mathbf{x}$ is a binary sequence.
-Since we deal with binary and non-binary cases in a unified way, it is convenient to generalize the notion of regularity to general alphabets.
+Let $A(\mathbf{x})=A(\mathbf{x},b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},b)$ be the matrix representations of $\mathbf{x}$ and $\mathbf{x}^{\prime}$, respectively.
+If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals.
+By II.2, this further reveals information of error positions in remaining rows of $A(\mathbf{x})$, or equivalently, error positions in $\mathbf{x}$.
+Detailed analysis will be given Lemma IV.5.
+Before that, it is convenient to generalize the notion of regularity to general alphabets.
 To that end, we associate with any $q$-ary sequence a binary sequence.
@@
 Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$.
+We call $u(\mathbf{x})$ the binary sequence associated with $\mathbf{x}$.
 Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$ and when $q=2$, we have $u(\mathbf{x})=\mathbf{x}$.
 It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
+Example IV.2
+Let $q=3$. We have that $\left\lceil q/2\right\rceil=2$. If $x=1$, then $u_{x}=0$ and $v_{x}=1$. If $x=2$, then $u_{1}=1$ and $v_{x}=0$.
 Remark IV.1
@@
 In other words, $\mathbf{x}$ is $d$-regular if and only if any substring of $\mathbf{x}$ of length at least $d\log n$ contains two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$.
-[…]
+Let $\mathcal{R}_{q,n,d}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\mathbf{x}\text{ is }d\text{-regular}\right\}$.
+To estimate a lower bound on our code size in Theorem IV.1, we need a lower bound on $\left|\mathcal{R}_{q,n,d}\right|$.
 Lemma IV.4
@@
   • If $q$ is odd, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n-1}$ for all $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$.
-Note that when $d\geq 10$, we have $-1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large.
+Note that when $d\geq 10$, we have $1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large.
 In particular, when $d=10$, it suffices to require $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
@@
 Suppose first that $q$ is even.
-Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2.
-On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2.
+Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<q/2$.
+On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $0\leq v_{x}<q/2$.
 Then it follows from Lemma IV.1 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
@@
 For $m\geq 2$, define
-$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }0\leq v_{i},v_{i+1}\leq(q-1)/2\right\},$
-$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }(q+1)/2\leq v_{i},v_{i+1}<q\right\}.$
+$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\nexists 1\leq i<m\text{ such that}\\ 0\leq v_{i},v_{i+1}\leq(q-1)/2\end{array}\right\},$
+$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\nexists 1\leq i<m\text{ such that}\\ (q+1)/2\leq v_{i},v_{i+1}<q\end{array}\right\}.$
 We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$.
@@
 Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$.
-For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$.
+For $\mathbf{v}\in S_{m}^{L}$, it must hold that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$.
 So it follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
@@
 Let $m=\left\lfloor\frac{d}{2}\log n\right\rfloor$ and
-$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\text{ and }v_{j}.v_{j+1}\geq(q+1)/2\right\}.$
-Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$. Now let $k=\left\lfloor n/m\right\rfloor$ and define
-=$\mathcal{R}^{\prime}=\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-k\left\lfloor n/m\right\rfloor}.\end{array}\right\}$
-It is straightforward to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus
+$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\\ \text{and }v_{j}.v_{j+1}\geq(q+1)/2\end{array}\right\}.$
+In other words, $Q$ is the set of length-$m$ sequences which contain two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$.
+Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$.
+Now let $k=\left\lfloor n/m\right\rfloor$ and define $\mathcal{R}^{\prime}$ to be the following set
+$\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-km}\end{array}\right\}.$
+It is easy to see that for any $\mathbf{x}\in\mathcal{R}^{\prime}$, any substring of $\mathbf{x}$ of length at least $d\log n$ must contain some $\mathbf{x}^{(i)}$, where $1\leq i\leq k$.
+Therefore, $\mathbf{x}$ is $d$-regular and hence $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$.
+This implies that
 $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}^{\prime}\right|$ $=\left|Q\right|^{k}q^{n-km}$
@@
 ∎
-Notations: In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$.
-Denote the $i$-th row of $A(\mathbf{x})$ and $U(\mathbf{x})$ by $A(\mathbf{x})_{i}$ and $U(\mathbf{x})_{i}$, respectively.
-It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$.
-Let
-$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\},$
-which is a subset of $[n]$. Then we have $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
+In Appendix B, we will discuss how to encode a sequence into a $d$-regular sequence.
+Now we return to the aim of this subsection: approximately determine error positions.
+Let $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}$ be obtained from $\mathbf{x}$ by a 22-$b$-burst-deletion.
+Let $u(\mathbf{x})$ and $u(\mathbf{x}^{\prime})$ be the binary sequences associated with $\mathbf{x}$ and $\mathbf{x}^{\prime}$, respectively (see the paragraph prior to Example IV.2).
+Then we have $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}^{b}\left(u(\mathbf{x})\right)$.
+By the relationship between $\mathbf{x}$ and $u(\mathbf{x})$, it suffices to locate the 22-$b$-burst-deletion in $u(\mathbf{x})$.
+In the rest of this section, let $U(\mathbf{x})=A(u(\mathbf{x}),b)$ be the matrix representation of $u(\mathbf{x})$.
+Denote the $i$-th row of $U(\mathbf{x})$ by $U(\mathbf{x})_{i}$.
+Similarly, we can define $U(\mathbf{x}^{\prime})$ and $U(\mathbf{x}^{\prime})_{i}$.
+Since $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}^{b}\left(u(\mathbf{x})\right)$, it follows from II.2 that $U(\mathbf{x}^{\prime})_{i}$ is obtained from $U(\mathbf{x})_{i}$ by two deletions.
+Suppose that $U(\mathbf{x})_{1}$ is $d$-regular and $\eta\left(U(\mathbf{x})_{1}\right)$ is given.
+According to Lemma IV.2, we can efficiently recover $U(\mathbf{x})_{1}$.
+Then by Lemma IV.3, the two deletions in $U(\mathbf{x})_{1}$ can be approximately located.
+In Lemma IV.5 below, we will show that this helps to locate errors in $u(\mathbf{x})$.
+To explain the idea, more notations are needed.
+For $1\leq i\leq b$ and $1\leq j\leq n/b$, let $U(\mathbf{x})_{i,j}$ be the $j$-th coordinate of $U(\mathbf{x})_{i}$.
+By the relationship between $u(\mathbf{x})$ and $U(\mathbf{x})$, it is easy to see that $U(\mathbf{x})_{1,j}=u(\mathbf{x})_{1+(j-1)b}$.
+This justifies the following definition
+$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\}.$
+Note that $I_{n,b}$ is a subset of $[n]$. Clearly, we have $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
+Example IV.3
+Let $n=10$, $b=2$ and $\mathbf{x}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}2\in\Sigma_{3}^{10}$.
+Then $u(\mathbf{x})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1$.
+By definition, we have $I_{n,b}=\{1,3,5,7,9\}$ and $u(\mathbf{x})_{I_{n,b}}=01100$.
+On the other hand, the matrix representations of $u(\mathbf{x})$ is
+$U(\mathbf{x})=\begin{pmatrix}0&1&1&0&0\\ 0&1&0&0&1\end{pmatrix}$.
+It is easy to see that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
 Recall that $r\left(U(\mathbf{x})_{1}\right)$ is the number of runs in $U(\mathbf{x})_{1}$.
-For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run.
+For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run of $U(\mathbf{x})_{1}$.
 Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$.
@@
 In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$.
-In other words, $I_{j}^{L}$ is the interval of length at most $b-1$ in $[n]$ to the left of $I_{j}^{\prime}$.
-See Figure 1 for an illustration.
+In other words, $I_{j}^{L}$ is the interval of length at most b1b-1 in $[n]$ to the left of $I_{j}^{\prime}$.
 Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$.
 We omit $\mathbf{x}$ for simpler notations.
+Example IV.4
+Let $\mathbf{x}$ be the sequence in Example IV.3.
+There are three runs in $U(\mathbf{x})_{1}=01100$: 0, $11$ and $00$.
+Recall that $p_{j}$ indicates where the $j$-th run ends.
+Then we have $p_{1}=1$, $p_{2}=3$, $p_{3}=5$, $I_{1}=[1,1]$, $I_{2}=[2,3]$ and $I_{3}=[4,5]$.
+All intervals $I_{j}^{\prime}$ and $I_{j}^{L}$ are listed in the following:
+$\begin{array}[]{cc}I_{1}^{L}=\emptyset,I_{1}^{\prime}=[1,2],\\ I_{2}^{L}=[2,2],I_{2}^{\prime}=[3,6],\\ I_{3}^{L}=[6,6],I_{3}^{\prime}=[7,10].\end{array}$
+We briefly explain why we define intervals $I_{j}^{\prime}$ and $I_{j}^{L}$.
+A $b$-burst-deletion in $u(\mathbf{x})$ induces one deletion in each row of $U(\mathbf{x})$.
+If it is known that the deletion in the first row $U(\mathbf{x})_{1}$ occurs in the interval $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, we can only locate the deletion in the $i$-th row in the interval $\left[p_{j-1},p_{j}\right]$ when $i\geq 2$ (see II.2).
+Note that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
+When looking at $u(\mathbf{x})$, interval $I_{j}^{L}$ (if not empty) corresponds to the last b1b-1 coordinates in the $p_{j-1}$-th column of $U(\mathbf{x})$ and interval $I_{j}^{\prime}$ corresponds to columns of $U(\mathbf{x})$ indexed by $I_{j}$.
+As a result, the $b$-burst-deletion in $u(\mathbf{x})$ can be located in the interval $I_{j}^{L}\cup I_{j}^{\prime}$.
+See Figure 1 for an illustration.
 […]
@@
 Figure 1: Illustration of relationship between $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$.
+In the rest of this section, let $A(\mathbf{x})=A(\mathbf{x},b)$ be the matrix representation of $\mathbf{x}$ and $A(\mathbf{x})_{1}$ be the first row of $A(\mathbf{x})$.
 Lemma IV.5
@@
 Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$.
+That is to say, $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by a 22-$b$-burst-deletion.
 Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular.
@@
   2. $(2)$
-    or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
+    or find an interval $J\subseteq[n]$ of length at most $3bd\log(n/b)+b-1$, such that $D_{1},D_{2}\subseteq J$.
 Proof:
@@
 Recall that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
-So $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
+This implies that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
 Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.2, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$.
+Then we can know the number of runs in $U(\mathbf{x})_{1}$ and $I_{j}$ for all $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$.
 In addition, by Lemma IV.3, we can
   • either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $k_{1}\in I_{j_{1}}$ and $k_{2}\in I_{j_{2}}$;
-  • or find an interval $[c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
-By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
+  • or find an interval $c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
+Similar to the discussion in the paragraph above Lemma IV.5, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
 For the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$.
@@
 From now on, we assume that $\mathbf{x}\in\Sigma_{q}^{n}$ and $A(\mathbf{x})_{1}$ is $d$-regular, where $d$ will be specified later.
-Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$.
-Then by II.2, we have $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$.
-Since $U(\mathbf{x})_{1}$ is regular, Lemma IV.2 ensures that when given $\eta\left(U(\mathbf{x})_{1}\right)$, one can uniquely recover $U(\mathbf{x})_{1}$ from $U\left(\mathbf{x}^{\prime}\right)_{1}$.
-Then by Lemma IV.3, we can know the approximate positions of the two deletions in $U(\mathbf{x})_{1}$.
-According to Lemma IV.5, we can approximately determine the positions of the two bursts in $\mathbf{x}$.
+Let $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$, where $D_{1},D_{2}\subseteq[n]$ are two unknown disjoint intervals of length $b$.
+Suppose that $\eta\left(U(\mathbf{x})_{1}\right)$ is given.
+The first step of decoding $\mathbf{x}$ from $\mathbf{x}^{\prime}$ is to decode $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$.
+After this step, we can know the number of runs in $U(\mathbf{x})_{1}$ and $I_{j}$ for all $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$.
+Then by Lemma IV.5, one can approximately locate the two bursts in $\mathbf{x}$ if $\eta\left(U(\mathbf{x})_{1}\right)$ is given.
 Based on the two cases in Lemma IV.5, we split our discussion into Section IV-B and Section IV-C.
 Our codes for correcting two $b$-burst-deletions are presented in Theorem IV.1, whose proof is divided into Lemmas IV.6 and IV.7.
+In Section IV-B and Section IV-C, we follow notations defined in this subsection.
 Before proceeding, we need the following observation.
@@
 Let $m>b$.
-Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}\in\mathcal{B}_{1}^{b}(\mathbf{u})$.
+Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}$ is obtained from $\mathbf{u}$ by single $b$-burst-deletion.
 Then for any $b\leq i\leq m-b$, we have $\mathbf{v}_{[1,i-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[1,i]}\right)$ and $\mathbf{v}_{[i+1,m-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[i+1,m]}\right)$.
@@
 $x_{k}=\begin{cases}x^{\prime}_{k},\text{ if }k<p_{j_{1}-1}b-b+2,\\ x^{\prime}_{k-b},\text{ if }p_{j_{1}}b<k<p_{j_{2}-1}b-b+2,\\ x^{\prime}_{k-2b},\text{ if }p_{j_{2}}b<k\leq n.\end{cases}$ (13)
-In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is know to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
+In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is known to us for any $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
 Therefore, to decode $\mathbf{x}$ from $\mathbf{x}^{\prime}$, it is sufficient to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
@@
 Since $j_{2}-j_{1}\geq 2$, deletions in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ do not affect $\mathbf{x}_{I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}}$.
-As a result, (13) implies $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$.
-Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}\cup I_{j_{1}}}$, $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
+As a result, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ (i.e., $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$ is obtained from $\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}$ by single $b$-burst-deletion) and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$ due to the fact that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\subseteq I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\subseteq I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}$.
+Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}$ (i.e., the concatenation of $\mathbf{x}_{I_{j_{1}-1}}$ and $\mathbf{x}_{I_{j_{1}}}$), $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
 $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}-1}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right),$ (14)
@@
 Recall that the range of $j_{1}$ is $\left[1,r(U(\mathbf{x})_{1})\right]$.
-To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 00) and $\mathbf{x}^{\prime}_{\left[p_{-2}b+1,-b\right]}$ be an empty sequence.
+To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 0) and $\mathbf{x}^{\prime}_{\left[p_{-1}b+1,-b\right]}$ be an empty sequence.
 By (14) and (15), to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$, we need to apply to each $\mathbf{x}_{I_{j}^{\prime}}$ a code for correcting single $b$-burst-deletion.
@@
 Then by Lemma II.2, our goal boils down to recovering the values of $\phi\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$.
-Since $A(\mathbf{x})_{1}$ is $d$-regular, we have $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$.
-Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $[0,N-1]$, where $N=2^{\log\log(n)+O_{q,d,b}(1)}$.
+Since $U(\mathbf{x})_{1}$ is $d$-regular and each $U(\mathbf{x})_{1,I_{j}}$ is a run of $U(\mathbf{x})_{1}$, we have $\left|I_{j}\right|\leq d\log(n/b)$ and hence $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$.
+Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in [0,N1][0,N-1], where $N=2^{\log\log(n)+O_{q,d,b}(1)}$.
 Here, $O_{q,d,b}(1)$ denotes a constant dependent only on $q,d$ and $b$.
@@
 Therefore, sequence $\overline{\phi}(\mathbf{x})$ suffered two bursts (of length two) of erasures.
-To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.7 and define
+To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.6 and define
 $f(\mathbf{x})=\varphi\left(\overline{\phi}(\mathbf{x})\right).$
-Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.7, $f(\mathbf{x})$ can be viewed as an integer in $[0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
-The following lemma follows from the above analysis, Lemmas II.2 and II.7.
+Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.6, $f(\mathbf{x})$ can be viewed as an integer in $0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
+The following lemma follows from the above analysis, Lemmas II.2 and II.6.
 Lemma IV.6
 Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$.
-Given $f(\mathbf{x})$, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
+Given $f(\mathbf{x})$, we can efficiently recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
 IV-C When Case (2) in Lemma IV.5 happens
@@
 In this case, we have already decoded $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ by using $\eta\left(U(\mathbf{x})_{1}\right)$.
-Furthermore, by case (2) in Lemma IV.3, we found an interval $[c_{1},c_{2}]\subseteq[n/b]$ of length at most $3db\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$).
-By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c_{1}-1,c_{2}]$, for all $2\leq i\leq b$.
-Therefore, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$.
-Note that $\left|[c_{1}-1,c_{2}]\right|$=$\left|J\right|+1\leq 3db\log(n/b)+1$. Let $P=3db\log(n/b)+1$ and define
+Furthermore, by case (2) in Lemma IV.3, we found an interval $c_{1},c_{2}]\subseteq[n/b]$ of length at most $3d\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$).
+By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $c_{1}-1,c_{2}]$, for all $2\leq i\leq b$.
+Let $A(\mathbf{x})_{i,k}$ be the $k$-th coordinate in the $A(\mathbf{x})_{i}$.
+Then $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$.
+In fact, it holds that $A(\mathbf{x})_{i,k}=A(\mathbf{x}^{\prime})_{i,k}$ when $1\leq k<c_{1}-1$ and $A(\mathbf{x})_{i,k}=A(\mathbf{x}^{\prime})_{i,k-2}$ when $c_{2}<k\leq n/b$.
+Let $P=3d\log(n/b)+1$. Define
 $J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/bP\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/bP\right\rceil-1.\end{cases}$
-Since $c_{1}$ and $c_{2}$ are known, we can find some $j_{0}$ such that $[c_{1}-1,c_{2}]\subseteq J_{j_{0}}$.
+Since $c_{1}$ and $c_{2}$ are known and $\left|[c_{1}-1,c_{2}]\right|\leq 3d\log(n/b)+1$, we can find some $j_{0}$ such that $c_{1}-1,c_{2}]\subseteq J_{j_{0}}$.
 Suppose $J_{j_{0}}=[d_{1},d_{2}]$.
 According to above discussion, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin J_{j_{0}}$.
-To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$.It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i,J_{j_{0}}}\right)$.
+To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$.
+It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$ is obtained from $A(\mathbf{x})_{i,J_{j_{0}}}$ by two deletions.
 Therefore, we can apply Lemma II.4 to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$.
-Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
-$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$
+Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $1\leq j\leq\left\lceil n/bP\right\rceil-1$, let function $g_{j}(\mathbf{x})$ be defined in (16).
+$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$ (16)
 Note that in the above definition of $g_{j}(\mathbf{x})$, we distinguish the two cases where $q>2$ and $q=2$.
 This is because when $q=2$, we have $A(\mathbf{x})=U(\mathbf{x})$, and hence the first row of $A(\mathbf{x})$ was already recovered.
-Since $\left|J_{j}\right|\leq 2P=6db\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $+O_{q,d,b}(1)$.
+Since $\left|J_{j}\right|\leq 2P=6d\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $O_{q,d,b}(1)$.
 Therefore, each $g_{j}(\mathbf{x})$ is a binary vector of length at most $7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)$ (when $q>2$) or $7(b-1)\log\log n+O_{d,b}(1)$ (when $q=2$).
 Therefore, we can view $g_{j}(\mathbf{x})$ as an integer in $\left[0,N_{2}-1\right]$, where
-$N_{2}=\begin{cases}2^{7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)},\mbox{ if }q>2,\\ 2^{7(b-1)\log\log n+O_{d,b}(1)},\mbox{ if }q=2.\\ \end{cases}$ (16)
+[…]
 For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/bP\right\rceil-1:j\equiv s\pmod{2}\right\}$. For $s\in\{0,1\}$, define
@@
 Lemma IV.7
-For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for each $2\leq i\leq b$.
+For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for all $i$.
 In particular, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
@@
 $\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
-Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code.
+Then $\mathcal{C}_{1}$ is a 22-$b$-burst-deletion correcting code.
 Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O_{q,b}(1)$:
@@
 Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$.
-Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
+Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $3bd\log(n/b)+b-1$, such that $D_{1},D_{2}\subseteq J$.
 For the former case, Lemma IV.6 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$.
@@
 By Lemma IV.4, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $A(\mathbf{x})_{1}$ is $d$-regular.
-It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, and $c_{1}$, such that
+It then follows from the pigeonhole principle that there are some $a$, $b$, $c_{0}$, and $c_{1}$, such that
 $\left|\mathcal{C}_{1}\right|\geq\frac{q^{n-1}}{N_{0}\cdot N_{1}\cdot(N_{2})^{2}},$
@@
 In this section, it is always assumed that $q>2$.
+Recall that $\mathcal{B}_{t}(\mathbf{x})$ is the set of sequences obtained from $\mathbf{x}$ by $t$ deletions.
 We aim to give a new construction of $q$-ary two-deletion correcting codes.
@@
 For the former case, we need a $q$-ary single-deletion correcting code.
-We use the code given in [23]
+We use the code given in [23].
 Lemma V.1
@@
 Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$.
-Therefore, by Lemma V.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$.
+Therefore, by Lemma V.1, $\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$.
 Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime.
-By Lemma II.6, we have $n\leq Q<2n$.
+By the following lemma, we have $n\leq Q<2n$.
+Lemma V.2 (Bertrand–Chebyshev theorem)
+For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
 Define
@@
 $J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
+Let $\xi_{1}(\cdot)$ be the function in Lemma II.4, which can help to correct two deletions in $q$-ary sequences.
 Since $\left|J_{1}\right|\leq 2P=6d\log n$, we can view each $\xi_{1}\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{7\left\lceil\log q\right\rceil\log(6d\log n)+O_{q}(1)}$ (see Lemma II.4).
@@
 Then $\mathcal{C}_{2}$ is a two-deletion correcting code.
-Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
+Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there are some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
   • $q$ is even, $d=7$ and $n\geq 9$;
   • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
-The proof of Theorem V.1 follows directly from Lemmas V.2 and V.3. The claim for redundancy follows from Lemma IV.4 and Lemma II.6.
+The proof of Theorem V.1 follows directly from Lemmas V.3 and V.4. The claim for redundancy follows from Lemma IV.4 and Lemma V.2.
 VI Conclusion
@@
 Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
+In this work, it is required that the two bursts have the same length.
+Allowing the two deletion-bursts to have different sizes would make the error model more general and applicable to a wider range of practical scenarios.
+Indeed, such a generalization would be more realistic, as burst deletions in real-world applications (e.g., communication systems or storage systems) may not always have equal lengths.
+The case where the two bursts have different lengths seems more involved.
+Extending our work to handle bursts of different sizes is a promising direction for future research.
 A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$.
-Our technique fails in this setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence.
-We leave this problem for future research.
-The following Lemma .1 is derived by slightly modifying the proof of [42, Proposition 3.3.7]. Before showing its proof, we list some facts.
-Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}-\cdots$.
-Fact .1
-For all $10\leq x<$ 1, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
+Our technique fails in these setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence.
+This problem is also left for future research.
+Acknowledgement
+The authors express their gratitude to the anonymous reviewers for their detailed and constructive comments which are very helpful to the improvement of the presentation of this paper.
+The authors would also like to thank Prof.
+Alberto Ravagnani, the associate editor, for his excellent editorial job.
+Appendix A A Lemma
+The following Lemma A.1 is derived by slightly modifying the proof of [43, Proposition 3.3.7].
+Before showing its proof, we list some facts.
+Let $\ln(\cdot)$ be the natural logarithm.
+Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}$-$\cdots$.
+Fact A.1
+For all $0\leq x<1$, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
 Proof:
-When $10\leq x<$ 1, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$.
-Since $-\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$.
-At last, it is easy to see that $-\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$.
-When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
+When $0\leq x<1$, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$.
+Since $\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $\ln(1+x)\leq-x+\frac{x^{2}}{2}$.
+At last, it is easy to see that $\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$.
+When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
 ∎
-Lemma .1
+Lemma A.1
 Suppose that $q\geq 2$ is fixed and $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. Then it holds that
@@
 It is easy to verify that
-$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $=-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$ (23)
-$=-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
-$=1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
-$=1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
-Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $10<\epsilon q<$ 1.
-Then it follows from .1 that $-\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $-\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$.
+$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ (24)
+$=$ $-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)$
+$\quad-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$
+$=$ $-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)$
+$\quad+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
+$=$ $1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)\right$.
+$\quad\left.-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
+$=$ $1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)\right$.
+$\quad\quad\quad\quad\quad\left.-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
+Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $0<\epsilon q<1$.
+Then it follows from A.1 that $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$.
 Now by Equation 24, we have
-$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $\leq 1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}\right.\right.$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
-$=1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
-$=1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$
+$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$
+$\leq$ $1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}\right$.
+$\quad\quad\left.+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
+$=$ $1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}\right$.
+$\quad\quad\quad\quad\quad\quad\quad\left.+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
+$=$ $1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
+$\quad\quad\quad\quad\quad+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$
 $=$ $1-\frac{\epsilon^{2}q^{2}}{2(q-1)\ln q}+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
@@
 where (a) follows from the fact that $\epsilon<\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)$. ∎
+Appendix B Encoding sequences into regular sequences
+In this section, we give two methods to encode a sequence into a $d$-regular sequence using only one redundant symbol.
+The first method is an instantiation of the encoding idea in the proof of [19, Lemma 11].
+The second method relies on the sequence replacement technique.
+The first method works for more flexible parameters $n$ and $d$.
+The time complexity of the second method is lower.
+For any $q\geq 2$, let $Q\subseteq\Sigma_{q}^{m}$ be the set of sequences containing two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$.
+According to [19, Lemma 10], it holds that $\left|Q\right|\geq 2^{m}-2\times(1.62)^{m+2}$ when $q=2$.
+By the definition of $q$-ary $d$-regular sequences, it is clear that $\left|Q\right|\geq q^{m}-2\times(1.62)^{2}\times(0.81q)^{m}$ for any even $q$.
+In Lemma IV.4, it is shown that $\left|Q\right|\geq q^{m}-2\times(0.87q)^{m}$.
+B-A The first method
+Suppose that $d$ and $n$ satisfy conditions in Lemma IV.4. In this subsection, let $m=\left\lfloor\frac{d}{2}\log n\right\rfloor$, $k=\left\lfloor n/m\right\rfloor$ and
+$\mathcal{R}^{\prime}=\bigg{\{}\mathbf{x}=\mathbf{x}^{(0)}\cdots\mathbf{x}^{(k-1)}$ $\mathbf{x}^{(k)}\in\Sigma_{q}^{n}:$
+$\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 0\leq i\leq k-1\\ \mathbf{x}^{(k)}\in\Sigma_{q}^{n-km}\end{array}\bigg{\}}.$
+Then in the proof of [19, Lemma 11] and Lemma IV.4, it is shown that each sequence in $d$-regular and $\left|\mathcal{R}^{\prime}\right|\geq q^{n-1}$.
+We aim to encode a sequence of length n1n-1 into $\mathcal{R}^{\prime}$.
+Before showing the algorithm, we need to define some functions.
+For $\mathbf{x}\in\Sigma_{q}^{n-1}$, define $f_{\textup{dec}}(\mathbf{x})=\sum_{i=1}^{n-1}x_{i}q^{i-1}$.
+It is clear that $f_{\textup{dec}}(\mathbf{x})\in\left[0,q^{n-1}-1\right]$ and $f_{\textup{dec}}(\mathbf{x})$ can be computed in $O(n)$ time.
+We further assume that $n$ is sufficiently large such that $m\geq 8$ when $q$ is odd, or such that $m\geq 12$ when $q$ is even.
+Under this assumption, it is easy to verify that $\left|Q\right|\geq q^{m-1}\geq q^{n-km}$.
+Since $\left|Q\right|^{k}-1+\left|Q\right|^{k}\left(q^{n-km}-1\right)=\left|\mathcal{R}^{\prime}\right|-1\geq q^{n-1}-1$, each integer $a$ in $\left[0,q^{n-1}-1\right]$ can be uniquely written as $a=\sum_{i=0}^{k}a_{i}\left|Q\right|^{i}$, where $0\leq a_{0},a_{1},\ldots,a_{k-1}<\left|Q\right|$ and $0\leq a_{k}<q^{n-km}\leq\left|Q\right|$.
+Denote $g_{Q}(a)=\left(a_{0},\ldots,a_{k-1},a_{k}\right)$.
+The time complexity of computing $g_{Q}(a)$ is $O\left(\log_{\left|Q\right|}\left(q^{n-1}-1\right)\right)=O\left(\log\left(q^{n-1}-1\right)/\log\left|Q\right|\right)=O\left(n/m\right)$.
+Since $\left|Q\right|\leq q^{m}=O\left(n^{d\log q/2}\right)$, we can efficiently construct the set $Q$ by brute-force.
+There is a bijection between the set $Q$ and the integer set $\left[0,\left|Q\right|-1\right]$.
+We can build a lookup table which gives such a bijection.
+Then for each sequence in $Q$ we can find its corresponding integer in $\left[0,\left|Q\right|-1\right]$ in $O\left(n^{d\log q/2}\right)$ time.
+Similarly, there is a bijection between $\Sigma_{q}^{n-km}$ and the set $\left[0,q^{n-km}-1\right]$ and for each sequence in $\Sigma_{q}^{n-km}$, we can find its corresponding integer in $\left[0,q^{n-km}-1\right]$ in $O\left(n^{d\log q/2}\right)$ time by a lookup table.
+Our encoding algorithm is described as follows.
+For a sequence $\mathbf{x}\in\Sigma_{q}^{n-1}$, compute $f_{\textup{dec}}(\mathbf{x})$ and $g_{Q}\left(f_{\textup{dec}}(\mathbf{x})\right)$.
+Suppose that $g_{Q}\left(f_{\textup{dec}}(\mathbf{x})\right)=\left(a_{0},\ldots,a_{k-1},a_{k}\right)$, where $0\leq a_{0},\ldots,a_{k-1}<\left|Q\right|$ and $0\leq a_{k}<q^{n-km}$.
+As analyzed above, we can map $a_{i}$ to a sequence $\mathbf{y}^{(i)}$ in $Q$ for each $0\leq i\leq k-1$, and map $a_{k}$ to a sequence $\mathbf{y}^{(k+1)}$ in $\Sigma_{q}^{n-km}$.
+At last, concatenate these sequences and we get the output $\mathbf{y}=\mathbf{y}^{(0)}\cdots\mathbf{y}^{(k-1)}\mathbf{y}^{(k)}$, which is $d$-regular.
+Clearly, the time complexity of the encoding process is $O\left((k+1)n^{d\log q/2}\right)=O\left(n^{d\log q/2+1}/m\right)$.
+Decoding $\mathbf{y}$ to the original sequence is straightforward. And the time complexity is also $O\left(n^{d\log q/2+1}/m\right)$.
+B-B The second method
+In this subsection, we give a new encoding algorithm based on the sequence replacement technique.
+This technique has been widely used in the literature to encode a sequence to a sequence without forbidden substrings [29, 44, 36, 38].
+Let $m=\left\lfloor\frac{d}{3}\log n\right\rfloor$ (where $d$ will be specified later) and $Q$ be the set defined at the beginning of this section.
+Then we have that $\left|\Sigma_{q}^{m}\setminus Q\right|\leq(c\cdot q)^{m}$ for some $0<c<1$.
+Therefore, there is an injective mapping $f_{Q}$ from $\Sigma_{q}^{m}\setminus Q$ to the set $\Sigma_{q}^{\left\lceil m\left(1+\log_{q}c\right)\right\rceil}$.
+Then the inverse mapping $f_{Q}^{-1}:\left\{f_{Q}:\mathbf{z}\in\Sigma_{q}^{m}\setminus Q\right\}\rightarrow\Sigma_{q}^{m}\setminus Q$ is well-defined.
+In addition, by building a lookup table, both $f_{Q}$ and $f_{Q}^{-1}$ can be computed in $O\left((cq)^{m}\right)=O\left(n^{\frac{d}{3}\log(cq)}\right)$ time.
+In the rest of this subsection, denote $k=\left\lfloor n/m\right\rfloor$.
+Fix a suitable $d$ such that when $n$ is sufficiently large, we have $m\geq\left\lceil m\left(1+\log_{q}c\right)\right\rceil+\left\lceil\log_{q}k\right\rceil+4$.
+For a sequence $\mathbf{x}\in\Sigma_{q}^{n-1}$, we first append a symbol 0 to $\mathbf{x}$ and get the sequence $\tilde{\mathbf{x}}=\mathbf{x}0$.
+This appended symbol 0 will help to indicate when our decoding algorithm ends.
+Next, we partition sequence $\tilde{\mathbf{x}}$ as $\tilde{\mathbf{x}}=\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(k)}\tilde{\mathbf{x}}^{(k+1)}$, where $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}$ for each $1\leq i\leq k$ and $\tilde{\mathbf{x}}^{(k+1)}\in\Sigma_{q}^{n-km}$.
+For any symbol $\alpha\in\Sigma_{q}$ and integer $\ell\geq 1$, let $\alpha^{\ell}$ denote the sequence consisting of $\ell$ symbols $\alpha$.
+For simpler notations, denote $m_{1}=\left\lfloor\frac{m-\left\lceil m\left(1+\log_{q}c\right)\right\rceil-\left\lceil\log_{q}k\right\rceil}{2}\right\rfloor$ and $m_{2}=\left\lceil\frac{m-\left\lceil m\left(1+\log_{q}c\right)\right\rceil-\left\lceil\log_{q}k\right\rceil}{2}\right\rceil$.
+The idea behind the encoding algorithm is as follows.
+  1. Step 1
+    For each $1\leq i\leq k$, check if $\tilde{\mathbf{x}}^{(i)}\in Q$.
+If not, delete $\tilde{\mathbf{x}}^{(i)}$ from $\tilde{\mathbf{x}}$ and still denote the resulted sequence by $\tilde{\mathbf{x}}$.
+  2. Step 2
+    Then we append the sequence $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ to the end of $\tilde{\mathbf{x}}$.
+Here, we use $b_{q}(i)$ to denote the $q$-ary representation of integer $i$.
+Clearly, $b_{q}(i)\in\Sigma_{q}^{\left\lceil\log_{q}k\right\rceil}$.
+  3. Step 3
+    Note that the substring $\tilde{\mathbf{x}}^{(k+1)}$ will never be deleted after Step 1 and Step 2.
+Continue Step 1 and Step 2 until the left of $\tilde{\mathbf{x}}^{(k+1)}$ is empty or all $\tilde{\mathbf{x}}^{(i)}$ to the left of $\tilde{\mathbf{x}}^{(k+1)}$ is in $Q$.
+Since $m\geq\left\lceil m\left(1+\log_{q}c\right)\right\rceil+\left\lceil\log_{q}k\right\rceil+4$, we have $m_{2}\geq m_{1}\geq 2$.
+Therefore, once Step 2 is executed, a sequence in $Q$ is appended to the end.
+Note that the length of $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ is $m$.
+This implies that the length of $\tilde{\mathbf{x}}$ does not change after Step 2.
+Denote the sequence at the end of this algorithm by $\mathbf{y}$.
+The fact that $m=\left\lfloor\frac{d}{3}\log n\right\rfloor$ implies $d\log n>3m-3$.
+Then it is easy to see that any substring of $\mathbf{y}$ of length at least $d\log n$ either contains a $\tilde{\mathbf{x}}^{(i)}$ (which is on the left of $\tilde{\mathbf{x}}^{(k+1)}$) or contains a substring of the form $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ (which is on the right of $\tilde{\mathbf{x}}^{(k+1)}$).
+Therefore, $\mathbf{y}$ is $d$-regular.
+The formal algorithm is given in Algorithm 1.
+Note that in Algorithm 1, it always holds that $\tilde{\mathbf{x}}^{(j+1)}=\tilde{\mathbf{x}}^{(k+1)}$.
+In each loop, if $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}\setminus Q$, a block of length $m$ to the left of $\tilde{\mathbf{x}}^{(j+1)}$ is deleted.
+Therefore, in Line 6, the substring between $\tilde{\mathbf{x}}^{(j+1)}$ and $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ is $\tilde{\mathbf{x}}_{\left[n-(k-j)m+1,n\right]}$, which the suffix of $\tilde{\mathbf{x}}$ (in previous loop) of length (kj)m(k-j)m.
+There are at most $k$ loops.
+In each loop, if $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}\setminus Q$, we have to compute $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)$ and $b_{q}(i)$.
+Therefore, the overall time complexity is $O\left(kn^{\frac{d}{3}\log(cq)}\right)=O\left(n^{\frac{d}{3}\log(cq)+1}/m\right)$.
+Denote the encoder based on Algorithm 1 by $\textup{Enc}(\cdot)$.
+We present the decoding algorithm in Algorithm 2, where the function $\textup{Dec}(\cdot)$ computes the decimal representation of an input sequence.
+In the Initialization step of Algorithm 1, a symbol 0 is appended to $\mathbf{x}$.
+In Line 6 of Algorithm 1, the appended block ends with symbol q1q-1.
+Therefore, in Algorithm 2, the fact that $x_{n}=q-1$ implies that there is an appended block which has not been deleted and the decoder has to go into another while loop.
+Suppose that $\mathbf{x}=\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(i-1)}\tilde{\mathbf{x}}^{(i+1)}\cdots$ $\tilde{\mathbf{x}}^{(j)}\tilde{\mathbf{x}}^{(j+1)}\tilde{\mathbf{x}}_{\left[n-(k-j)m+1,n\right]}f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}.$ Then it is clear that $\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil+1,n-m_{1}-m_{2}\right]}=b_{q}(i)$ and $\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil-\left\lceil m\left(1+\log_{q}c\right)\right\rceil+1,n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil\right]}$.
+Therefore, we get the index $i$ in Line 4 and $f_{Q}^{-1}$ is the deleted $\tilde{\mathbf{x}}^{(i)}$.
+When the while loop ends, we get the sequence $\mathbf{x}0$.
+In Line 8, the last symbol 0 is deleted.
+Now the correctness of Algorithm 2 is clear.
+There are at most $k$ while loops and in each loop, we have to compute $\textup{Dec}(\cdot)$ and $f_{Q}^{-1}(\mathbf{u})$.
+Therefore, the overall time complexity is $O\left(kn^{\frac{d}{3}\log(cq)}\right)=O\left(n^{\frac{d}{3}\log(cq)+1}/m\right)$.
+Input: $\mathbf{x}\in\Sigma_{q}^{n-1}$
+Output: $\tilde{\mathbf{x}}\in\mathcal{R}_{q,n,d}$
+1 Initialization
+2 $\tilde{\mathbf{x}}\leftarrow\mathbf{x}0$, write $\tilde{\mathbf{x}}=\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(k)}\tilde{\mathbf{x}}^{(k+1)}$, where $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}$ for each $1\leq i\leq k$ and $\tilde{\mathbf{x}}^{(k+1)}\in\Sigma_{q}^{n-km}$
+3 $i\leftarrow 1$, $j\leftarrow k$
+4 while $i\leq j$ do
+5  if $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}\setminus Q$ then
+6     $\tilde{\mathbf{x}}\leftarrow\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(i-1)}\tilde{\mathbf{x}}^{(i+1)}\cdots\tilde{\mathbf{x}}^{(j)}\tilde{\mathbf{x}}^{(j+1)}\tilde{\mathbf{x}}_{\left[n-(k-j)m+1,n\right]}f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$
+7     $j\leftarrow j-1$
+8 else
+9     $i\leftarrow i+1$
+10  end if
+11
+12 end while
+return $\tilde{\mathbf{x}}$
+Algorithm 1 Encoding a sequence into a $d$-regular sequence
+Input: $\tilde{\mathbf{x}}=\textup{Enc}(\mathbf{x})\in\Sigma_{q}^{n-1}$
+Output: $\mathbf{x}\in\Sigma_{q}^{n-1}$
+1 Initialization
+2 $\mathbf{x}\leftarrow\tilde{\mathbf{x}}$
+3 while $x_{n}=q-1$ do
+4  $i_{0}\leftarrow\textup{Dec}\left(\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil+1,n-m_{1}-m_{2}\right]}\right)$
+5  $\mathbf{u}\leftarrow\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil-\left\lceil m\left(1+\log_{q}c\right)\right\rceil+1,n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil\right]}$
+6  $\mathbf{x}\leftarrow\mathbf{x}_{\left[1,(i_{0}-1)m\right]}f_{Q}^{-1}\left(\mathbf{u}\right)\mathbf{x}_{\left[(i_{0}-1)m+1,n-m\right]}$
+7 end while
+[…]
+return $\mathbf{x}$
+Algorithm 2 Decoding a $d$-regular sequence into the original sequence